A Complete Bibliography of Publications in the ICES Journal of Marine Science (1990–1999)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

05 October 2023 Version 1.02

Title word cross-reference

a [DJ98]. G_{loss} [O’B99].

/DNA [BB94].

0- [NN91b]. 0-group [Arr96, BWH97, FG96, GG98, HH93, MJ97a, PLPN90].

1-group [GG98, NN91b]. 11-year [HW97]. 1800s [RM96a]. 1989 [Øst90]. 1990 [Bev91a, Dan97, HA92b, HA92a, LP91b]. 1992 [Mad94]. 1993 [Ano94c, Dic94]. 1995 [Ano96e, BFL97]. 1996 [GR98a, GRDCW98]. 1997 [Øst99].

20th [Thu97]. 25 [Ano94c].

4VsW [SP96].

75^{o} [Ric95].

âge [AL95].

abalone [HS99]. abnormalities [PM96a]. Absence [LGKM97]. Abundance [SW96, AR90, BS92, Box97, BI92, CHØ92, CNF90, CvdK96, ES92, FL98, FS93, GW93, GKT98, GMA98, GG98, HS91, HJR96, HJ99, HF94, Kap99, KZ97, Kre94, Lar92, MJ97a, MK99b, MP91, MGF96, MW96, Mis96, MSG96, NTWW95b, OF95, PLvdM^{+}96, PBSN98, PCM96, RvLDH96, RM96b, Sam91, SPL98, She99, SSJK98, Ste96, Tay95, VGHO97, VR98, HN90]. Acadian [MBSM90]. Acartia [KB95, THJ^{+}95]. accordance [MI96]. Accounting [DW99]. Accuracy [CM91, Ham96b]. acetate [RW91, RCE^{+}93]. acid [AGH^{+}93]. Acoustic [Ham96b, ML96, MSWL96, OM96, Orl99, Ros98, SWCG94, SCW96, YF96, Agl96, AY99, BMS98, BJM99, Foo90b, FS93, Gle96, GW93, GKT98, Gre97a, GD96, HM95, HGOT99, IMH96, Klo96, KWK97, LR99, MKP96, MAC96, MA92a, Mur96, Mut96, Nak90, Pet93, PCM96, Rey98, SDW^{+}96, SF96, SWB96, Spa90c, Sta94, TF96, TSAS96, WT96, YCL96, GD96, RGHC96]. Acoustical [SSD96, MJ96]. Acoustics [Sim96, ET96, MH96]. across [MBT95, WMO98]. activities [CR97, DM96, DBH98]. activity [Ber95, DLMS97, Gar97, HLAM95, PB97, UC92]. acutorostrata [HGLN95, MRL92, MEH97, MP99, NTWW95b]. acutus [SSH95]. Adams [SG95]. ADAPT [PK95a]. adaptations [Mau95]. Adaptive [PMF^{+}96]. Address [Mau96]. adequacy [KO97]. adjacent [CHØ92, JG93, SSFB97]. Admiralty [MRSF91]. ADN [BBG91, BB94]. Adolf [Hem95]. Adriatic [MUCM98]. adult [Ims99, hLHM^{+}92, MJ97a, TT96, ZM93]. advantage [KP97]. adverse [Rym97]. advice [Coo99]. Aegean [Ste99]. aeglefinus [Alb94b, BWH97, JJ96a, PV93]. Aequipecten [BB92]. aerial [EMN^{+}94]. Aeromonas [WTB99]. affect [AEH^{+}99, RR99]. affecting [Daw91, RM96b, WW96]. affinity [Ims99]. Africa [GDJC99, HS99, LHSA98, PB97, PH94]. African [Ham96b, Pay99, Wal94]. after [Kai98, SK90b]. Age [GCG96, KCB99, Ned90, AL95, BS92, BMM^{+}98, CM91, CR93, DKLM99, GMG98, HKN^{+}91, JM99, Jør90, KHØ95, KP97, LL97, LDU98, MC90, MS90, Mil92, MB97, MC99, RW91, RCE^{+}93, SN91, She99, Swa99]. age- [MS90]. age-0 [BS92, Swa99]. age-2 [BS92]. age-at-length [MC90]. age-composition [KHØ95]. age/length [DKLM99]. aggregate [SN99, dG96a]. aggregated [BHS96]. aggregation [Pet98]. aids [TSAS96]. Akita [SKS97]. Alan [Bev91a]. Alaska [BS92, GKT98, Swa99, TT96]. Alaskan [PCH91, WT96]. albacares [BJM99]. albirostris [NTWW95b]. alga [FPW91]. Algal [NBBS95]. algorithm [Cla94, MSG96, SBSH97]. algorithms [Coo99]. allocation [CC99]. Allozyme [Bea91, BCJ^{+}97]. allozymes [KP97]. along [JBC97, KN94, Køi99, McD98, NN91a, Nie99]. alternative [CR99, Dem99, Hor92a, SC96]. amberjack [RG93]. Ambient [OMN98a, OMN98b, MON98]. American [FRK93, MC99, WHW99]. americanus [CZM98]. amino [AGH^{+}93]. Ammodytes [Gau90]. among [Mis93, OB97, Ver97]. analyse [PL97]. analyses [GGF94, Mur96]. analysing [MUV90]. Analysis [JM99, PL97, PV93, PCM96, SN99, SFG96, Ste96, Swa97, BE90, BCCP96, CFUCY95, Cla94, Coo97, EMN^{+}94, GCG96, GMG98, Gud95, HL97, Jon90, JPG93, LS98a, LP91a, MC90, Moh99, OB97, PBSN98, She97, She99, SAC96, SWB96, SS99, SK95b, TZD^{+}95, VM98, VKA98, dVDP^{+}98, vLRvR^{+}96]. Anarhichas [FPHM90, PN93, PM96a]. Anchoa [HGDC94]. anchoita [BP91]. anchovies [BCCP96]. anchovy [BP91, BV96, Ham96b, HGDC94, JPG93, Nie94, Wal94]. Andersen [Mad94]. angle [HO96]. angles [Rey98]. Anglian [TIR97]. Angola [MLCB99]. Anguilla [Køi91, MA99]. Anguillicola [Køi91]. animals [AB96, SE96]. Annual [MFL98, BMS98, MH95, MUCM98, PCM96, RM96b, SFCP99]. annuli [Wal94]. Anoplopoma [KZ97]. Antarctic [HD96b, MGMS94, Mur96]. Antarctica [Atk95, MRSF91, SSH95]. anuli [dBH95]. appears [Jen95]. applicable [TA94]. Application [DM99, ET96, WT96, Bar98, CC99, GDJC99, GMG98, Har97, HLB91, MK98, MUV90, Pat98, Pet93, SN91]. applied [Chr97, HN91a, HN91b, Jen91]. Applying [MK99a]. approach [Chr97, DM99, Eve93, Hay98, HGOT99, Hor91, HL97, KOS^{+}99, MSPG92, Oel99, RVD90, SSS99, SMGB97, SRAL99]. approaches [Ste96]. April [Mad94, Øst99]. aquaculture [GB97]. Arabian [Smi95]. Archipelago [KRKS97, Raj93]. Arctic [BFL97, BFKR97, Jak92, Jør90, Jør92, MB96a, MRF97, MON98, OMN98a, OMN98b, Ynd99a]. Arctica [WK94, Wit96]. Arcto [SS93b, SS96]. Arctoscopus [SKS97]. arcuatus [RR98]. area [Agl96, CNF90, FS93, Foo95, HH93, MFPT98, Pat98, PR98, SOBP99]. areal [TL96]. areas [Ber95, DRR94, OR97, SSFB97]. Arendal [ADS98]. Argentina [Ber93, Gay98]. Argentinidae [Ber93]. argentinus [BI92]. aristotelis [Wan97]. ARN [BB94]. ARN/ADN [BB94]. aromatic [BSMK99, SK97]. array [Foo90a]. Arthur [Dic94]. Artificial [HG96, RR99, WM99]. artisanal [EGB^{+}96]. Aspartate [Ber95, HLAM95]. Aspects [BB92, SSS98, Ber95, Mal90]. aspera [PPS90b]. assemblage [PR98, RG96]. assemblages [GR98b, GK97, KBC^{+}92, RA97]. assess [BRK97a, BR93, McG97]. assessed [EMN^{+}94]. Assessing [ML95, Pat98, Ye98, vdML95, CR99, Gle96, KP97, MS90, RGH93]. Assessment [ADS96, CR93, MSC98, TT96, BB94, Ber95, Gre97a, HM99, Hay98, LDU98, MK98, MGMS94, SRJ98, SS93a, Som99, Ull96, WK99, dG96b]. assessments [FN99, Pet93]. Associated [Ano93b, LDU98, MGN91, SW96]. Association [Möl90]. Associations [CZM98, SP96]. assumption [HLB91]. Atlantic [AS95, AS97, AGH^{+}93, AH92, BP91, BME96, BW95, BDC^{+}94, BG93, BRK97b, BTC93, BCJ^{+}97, Bra95, BR93, BI92, CLAD97, CAWD97, CG95, Cas99, DMAG97, DPHM97, EHJN91, FE97, FØK94, FRK93, HTG97, HRL97, HJL99, HGLN95, IOEJ97, JG97, JB91, Jen99, Jon97, JG93, KP97, KWN95, LGKM97, Mac90, MRL92, MEH97, MP99, MBSM90, McD98, McG97, MT97a, MK99b, MRTB95, McV97, MM97, MT96, MB97, MFL98, MBT95, Nak90, Ned90, Nil97, NLP97, ODR92, OMF^{+}97, PT98, RVD90, Rit97, RP96, SRJ98, SHKL97, SE98, SC99, SP96, SL97, SNSC94, Ver97, WWS93, WW96, YWMW97, ZM93, dSJ95]. atlantica [Gre94, KLF96]. atlanticus [KWK97]. atmospheric [CH98]. attempt [CB93]. atypical [WTB99]. August [Ano96e, Mad94, GRDCW98]. auklets [Hun97]. Aurelia [BMM94, Mut94, Mut96, OIF95, SB94]. aureolum [Rai93]. aurita [SB94, BMM94, BP90, Mut94, Mut96, OIF95]. austral [HRP90]. Australia [AGKB91]. Australian [SSS99]. australis [MMHC98]. autocorrelation [vdML95]. automatic [GMG98]. autumn [AGH^{+}93, BMS98, MF91]. autumn-spawned [MF91]. auxiliary [Spa90c]. availability [GW93, HGLN95]. Avoidance [HD90, MLCB99, Nøt98]. Azores [WB95]. Azov [VVM95].

B [Bla93]. Back [FAN99, CB93, HJB90]. Back-calculated [FAN99]. back-calculation [CB93]. background [WB96]. backscatter [GD96, SAC96]. backscattered [Jaf96]. backscattering [CSW92, IMH96]. Bahía [Gay98]. Bait [GHM97, Løk98]. Balaenoptera [HGLN95, MRL92, MEH97, MP99, NTWW95b]. Balearic [OR97]. baleen [Bes93, CHØ92]. Ballantrae [RCM90]. ballast [Hay98]. Balsfjord [BE90, HN90]. Baltic [LM99, SE98, Arr96, AH96, Arr98, BSMK99, BLM99, DKLM99, Gis99, Gry99, HIL98, KRKS97, KNI92, Køi99, KP97, KSP96, LMW^{+}99, MK99b, ML99, Nil97, OLO98, Orl99, ROH^{+}98, SB94, SMGB97, Spa90c, Spa95, Spa96, Thu97, VHV^{+}98, WTB99, WH97]. Bank [SCM94, Alb95, HBA98, MBSM90, PV93, PLPN90, RCM90, WMO98]. Barents [KN99, BK96, BG94, FAN99, Fos92, FS97, GA94, GKT98, GG98, Har97, HP99, JJ96a, KHØ95, LOÅ97, NHP^{+}95, Sak97, SK97, Ynd99b, dBT98]. based [BW95, CC99, CHR96, DSO96, GCG96, Has96, Kap99, PPS92, RWJF90, RP95, SWB96, SS95a, Ye98, dBH95, vLRvR^{+}96]. baseline [Bal93]. Basil [Bla93]. Basin [RSN96, TSP^{+}96, WH97]. basins [ADS96]. bass [SHM95, SH98]. Bathymetric [Mau95]. Bay [LHSA98, MRSF91, OHN99, HGDC94, AL95, BS98, Gui95, JPG93, LS98b, OIF95, PCH91, SHM95, SH98, Sma98, VM98, VPT99]. Bayesian [MK98]. bays [CCC^{+}98]. be [BCJ^{+}97, LL97, MSRK99, NR96, Ull96]. beaked [NTWW95b]. beam [BJM99, ET96, HM95, Ham96a, MMHC98, MK96, MC98, Mut96, Ped96b, Rey98, RvL96, RBSV98, SFG96, SBH95, SHB96, SBSH97, SW96, SJ96, TF96]. beam-trawl [RvL96]. beamtrawl [WK94]. beamtrawling [BH92]. bed [KRKS97, RBSV98]. behavior [MJ96]. behavioral [BTS^{+}97]. behaviour [GHM97, HM95, HHA95, ML96, MA92a, MAI96, PMF^{+}96, SFG96, THJ^{+}95, WRT95]. behavioural [Jon97, Mac99]. Being [Hou90, OBD^{+}97]. Belgian [JBC97, MUV90]. Bellingshausen [Atk95]. beluga [HJR96]. beneath [WM99]. Benguela [HVMI^{+}95, VR98]. Benthic [HA92b, HA92a, AGKB91, Kai98, KBC^{+}92, Phi98, RPW^{+}99, RBSV98, SN99]. benthopelagic [Klo96]. benthos [BED^{+}93, BME96, FBG96, HSFAM99, Kin92, RRL^{+}92]. Bergen [Knu90]. Bering [Swa99, Hun97, RTK95, SSKP94]. better [Hor91, SF96]. between [AS97, Agl96, BTS^{+}97, BSMK99, CBG96, CZM98, CR97, CH98, Cus95a, DB92, FG96, Fry91, GG98, Hou94, Hul95, IMH96, Jen95, Jon97, KS90, LP91a, MAC96, MS96, MABD92, Möl90, OR97, Rij91, RvLDH96, RBSV98, SKS97, SN99, SP96, SBSH97, Spa96, SK95b, THJ^{+}95, WH98, vdML95]. between-haul [Fry91]. Beverton [Ano96e]. bias [MK99a, Rey96, SBH95]. bib [HH93]. bifurcations [Gre95]. bigeye [BJM99]. Bight [HCR94, Gre94, GRN96, Hic98, ML99, Rad98, SB94]. Bioacoustical [HP95]. biochemical [UC92]. Biocoenotic [GRN96]. Biodiversity [SSJK98, RPW^{+}99]. bioeconomic [Chr95, Chr97]. bioenergetics [Han96]. Biological [Jak92, Kim97, Pep90, BRK97b, DBH98, Fra97, Mac90, MH95, Neh98, RGHC96, Ull96]. Biology [Alb94a, HN90, TCA97, VGF97]. Biomagnification [Eve93]. Biomass [LLLL99, Pet98, Sak97, AH92, AG95, CH98, Ham96b, HBC^{+}92, Hor93, HHHS92, hLHM^{+}92, MSC98, MABD92, PL96, PW95, PM96b, RTK95, She91, SFG96, Spa90a, SV91, SS95a, Sta94, SS95b, Thu97, VHV^{+}98, ZF98, Ban95]. Biomass-dependent [Pet98]. biosynthesis [KH95]. biota [FN99, Ped96a]. bipunctata [KLF96]. Biscay [VM98, AL95, Gui95, JPG93, VPT99]. bivalve [Cla94, WK94]. bivalvia [WK94]. black [DAP97, Mut94, Nie94]. bladder [Rob92]. Blanca [Gay98]. bloom [MGN91, Rai93]. blue [BMM^{+}98, HHB96, KL97, MMHC98, OVP98]. blue-mouth [HHB96]. bluefin [HLB91, PKMP99]. bluemouth [KCB99]. BMB [LM99]. BMB/ICES [LM99]. board [CLB92, ROH^{+}98]. body [Mok92a, PPS90b]. Bohuslän [Cor99a]. Boreal [SSS98]. borealis [HN90]. Boreogadus [GA94]. Börje [WO93]. Bornholm [WH97]. boscii [SPL98]. Bothidae [SV98]. Bottom [Hee96, ES92, GWE99, KN99, Phi98, PB97, VR98]. bottom-trawl [GWE99, KN99]. bottom-up [VR98]. Boundary [Swa97, Mal90]. box [PR98]. Boyle [MI96]. Brachyura [GGF94]. brachyuran [McD98]. branchialis [VGHO97]. Brazilian [BP90]. Breeding [HW97, BK96, Gré97b, HTG97, OR97, WHG98]. British [Bra94, NTWW95b, RCE^{+}93]. broadband [TL96]. brood [HW97]. Brosme [BMM^{+}98]. brown [Bod96, Cus95b, FPW91]. bubble [Nøt98]. Bückmann [Hem95]. budget [APD92]. Buen [HCA93]. buoyancy [CNF90]. Burry [CWC^{+}97]. by-catch [Løk98, Phi98, TBHL97]. bycatch [EHEW94]. byssus [BB92].

cachinnans [Mun97]. Cage [Dem99]. caged [BFNP97]. Calanoid [FVC94, Ric95]. Calanoida [Mau95, SSH95]. Calanoides [SSH95]. Calanus [BHKH94, GA98, PTO95, SSH95]. calculated [FAN99]. calculation [CB93]. calibrating [Foo90b]. calibration [She97]. caligid [JDLH97]. Caligidae [MLBM98, MH98b, SKHM98]. Caligus [SKHM98]. Callionymidae [SV98]. Cameroon [NLCS99]. Can [NR96, THJ^{+}95, Ull96]. Canada [CR97, MBHQ96, ODR92, SOBP99]. Canadian [CLAD97, CAWD97, Ham96a, LGKM97, MB97, SL97]. Cancer [AL91]. canicula [OVP98]. cannibalism [PH94]. canonical [SK95b]. Cantabrian [OVP98]. canyon [SVS99]. Cap [CP96, SRJ98]. capacity [Ser90]. Cape [BH94, PB97, PH94]. capelin [AG98, BG94, Fos92, GG98, MP91, Nak90, Ros98, SCM94]. capensis [BH94, Wal94]. capillata [BMM94]. captively [BTS^{+}97]. capture [FA97, PLPN90, WHW99]. captured [AGH^{+}93, TIR97]. Carangidae [VGF97]. carbo [CC93, Gré97b, Gré97b]. carbon [AS97, BRK97a]. Carotenoids [NLP97]. Case [Sin91, HS99, JCP91, SRAL99, SG98]. Catch [Gré97b, CR93, ES92, EHEW94, GCM99, GHM97, Gud95, HHA95, Jen90, KN99, LL97, Løk98, PL97, Phi98, RSN96, SKS97, Sam91, SN91, She91, She99, Ste99, TBHL97]. catch-at-age [CR93, SN91, She99]. catch-at-length [Gud95]. catchability [GWE99, MGF96, SNSC94, TJ97]. catches [AGKB91, HJL99, Kap99, KP97, MSPG92, Pat98, PB97, RKCM92]. caught [Hee96]. Causal [Spa96]. caused [Kim97]. causes [BP96, Gre95, Laa96, Ped96a]. Cd [SKL98]. cell [NBBS95]. Celtic [HF94, Hor93, RPW^{+}99, TBHL97]. central [HHA95, Spa96, WH98, MP99]. Centurial [MM97]. century [Thu97]. Cephalopoda [CDL^{+}94, CBR95, HM99, JM99, LN93]. Cessation [Jen99]. Cestode [Mac90]. chaetognaths [KLF96]. chain [SH95]. chains [Man93]. chalcogramma [BS92, PPS90a, Swa99, Tra96]. chambered [LR96]. change [KHK94, MM97, RG96]. Changes [BMS98, Bod96, His96, Kai98, Laa96, MFGP96, PR98, Rij91, RvLDH96, RvL96, Rob92, SNSC94, vdWRvdLF90, Alb95, AH96, Arr98, AG95, BK96, BP96, CP96, Cas99, DLMS97, FH96, GR98b, GW93, GRN96, Hic98, Jen99, Jør90, Jør92, KWN95, LGKM97, MW96, MM97, Neh98, NR96, ODR92, Ped96a, PT98, Rad98, RR99, Ric95, Ste99, SG98, SVJ99, Tor90, WH96]. changing [Mil95]. Channel [GK97, Laa96, Mar93, RPW^{+}99, WHM97]. characterised [Pet97]. characteristics [GS96, SG95, SCW96]. characterization [SWCG94]. characterized [Pet98]. characterizing [Sma98]. characters [JPG93]. chemical [CZM98, Col98, LBDH98, ZM93]. Chesapeake [SHM95, SH98]. chez [AL95, BB94]. Chile [CBG96, HHA95]. China [LLLL99]. Chinook [CR99]. Chlamys [BB92]. chlorophyll [BC96, DJ98, LPL95, LS98b]. choice [MAYC96]. chokka [LHSA98, LDU98]. choosing [Jon90]. circular [Foo95]. circulation [Tur92]. clam [GCM99, RW91, RR98, SG95]. clarity [BS98]. class [Ser90, She97]. classes [GG98]. classification [MSWL96]. climate [GRN96, MM97, PT98, Smi95]. climatic [BEM96, LBDH98, Neh98, Tay95]. closed [Ye98]. Clupea [Arr96, AH96, BMS98, BH98, Fos96, Gry99, Hog94, HS98, JB91, KRKS97, Mac99, MF91, MF92, Mok92a, Mok92b, Nøt98, Raj93, RCM90, SPCLT96, Tor91, ZM93, dBH95, dBTT98, dBT98]. clupeoid [MBR95]. clusters [MKP96]. Clyde [TCA97, RCM90]. CMP [MUV90]. Cnidaria [OIF95]. co [SV98, SVJ99]. co-occurring [SV98]. co-reared [SVJ99]. coast [AL91, AGKB91, Bod96, CCC^{+}98, DJ98, FHA^{+}98, FPW91, HGD92, JBC97, KN94, McD98, ML99, MUV90, OF95, PU93, PB97, Rai93, RRL^{+}92, TIR97, Tor91, UJA^{+}93, dPVV^{+}96, Nie99, WK99]. Coastal [Ano94c, Alb95, ADS96, ADS98, BP90, BP91, BC96, FH96, GK97, HH93, HKN^{+}91, HBA98, KLF96, Kre94, KS90, Laa96, Lor91, Md95, MP91, SKS97, SE98, SN99, SJK90, SK90a, SK90b, dVDP^{+}98]. cockles [CWC^{+}97]. Cod [RdC95, AW92, AWEH94, BDS96, BH98, BG94, Bra94, Bra95, BWH97, CP96, Cas99, CR97, CKF91, Coo98, DJ92, Ehr91, ES92, FG96, FØK94, GA94, God95, Han96, HP99, HGD92, Jak92, Jør90, Jør92, KS90, KWN95, ML99, MON98, Mil92, MUV90, Moh99, MT96, MBHQ96, NL99, NP98, NK99, OMN98a, OMN98b, Pat98, PU93, PM96b, PLPN90, RAFC92, RJ90, RDvBH91, RP96, SRJ98, SC99, SSFB97, SP96, SMGB97, Spa95, Spa96, SS93b, SS96, SVJ99, SJK90, SK90a, SK90b, SNSC94, TA94, TJ97, WH97, Ynd99a, dBTT98, dSJ95]. cod-end [RAFC92]. codend [Chr95, Has96, KSP96, SPCP97, Ste99, SPCLT96]. codends [PS97]. coexistence [McV97]. Coho [BTS^{+}97]. cohort [Jon90, dBTT98]. cohorts [Col99, FØK94]. coindetii [GCG96, GMG98]. collapse [MBHQ96]. collection [NLCS99]. Collett [KWK97]. colony [Gar97]. colour [FL98]. Columbia [RSN96]. column [Mau95]. Combined [MC90, TZD^{+}95]. Combining [RKCM92, Ste96]. commercial [Lar92, PS97, RKCM92, WRT95]. commercially [Spa90b]. Common [CH98, FPHM90, Hay95, HCH91, Mut96, OHN99, PN93, WB97b]. communities [HHHS92, Kai98]. community [CFUCY95, HBA98, OIF95, SN99]. Comparative [BP90, BP91, Ehr91, PK95a, KZ97, Kre94, SS96]. Comparison [BS92, DPHM97, Foo95, GD96, HIL98, Mil96, Mur96, PLPN90, RSN96, SE98, GG98, JG97, Mis96, MB96b, PH94, RPW^{+}99, RKCM92, THJ^{+}95, UC92, WB97b, dVDP^{+}98]. Comparisons [OMF^{+}97, Ham96b, Hay95]. competition [FG96]. Complete [Ano95a, Ano96a]. Complexity [Coc99]. Component [PL97, PTO95]. Composition [EHEW94, AL91, Agl96, AGH^{+}93, AGKB91, Arr96, FG96, HCR94, HKN^{+}91, KHØ95, LLLL99, PTCC98, SV98, SN99, WH98]. compositions [Cor99b]. compounds [BSMK99]. comprehensive [FN99]. compromised [BCJ^{+}97]. computer [AH95, SK95b]. concentration [LHSA98, Ped96a, SKL98]. concentrations [ADS98, BC96, DRR94, Par99]. concept [SJB90]. concerning [BHKH94, MT97a, Tur92]. conclusions [LM99]. concurrent [RGHC96]. condition [BRK97a, PL96, UC92]. conditions [AG98, EMN^{+}94, Möl90, VM98]. configuration [AGKB91]. Confronting [SBSP99]. congeneric [Mau95]. conjugate [MK99a]. Conne [Dem99]. connection [Tay95]. Consequence [MAYC96]. Consequences [BGA97, OBD^{+}97, RP98, SFG96]. conservation [DM99, McD99, Nie99]. consideration [PM96b]. considerations [JWF99]. Constance [SG98]. constraints [HJ99, NBBS95]. consumption [ADS96, BG94, GBB^{+}97, HH93, HRBA91, LS97, MRL92, PPS90b, PPS92, UJA^{+}93, WB97a, WHG98, dSJ95]. contaminant [FN93, FN99, MUV90]. Contaminants [MB96a, CZM98, SK97]. contamination [HFP96]. content [AGH^{+}93, AH96, Arr98, BBG91, BRK97a]. context [KOS^{+}99]. continental [AH95, NN91a]. Continuous [MWRH98, DK96]. contribution [Rit97]. contributions [Ano95a, Ano96a, KP97]. control [Ban95, BFNP97, FH96, HJB90, HN95, RP99, SH95, VR98, Wit99]. convection [BFKR97, MJ97b]. conventional [HM95, Ham96a]. conversion [PPS90a]. Cook [TT96]. cooling [Cas99]. Copenhagen [HA92b, HA92a]. Copepod [HVMI^{+}95, Atk95, FVC94, HHHS92, KB95, NBBS95, PCH91, PHT95, VM98, VHV^{+}98]. Copepoda [FG95b, MLBM98, MH98b, SSH95, SKHM98]. copepods [FG95a, Hay95, KH95, ML95, PILB95, Ric95, TZD^{+}95, THJ^{+}95, dPVV^{+}96]. cormorant [CC93]. cormorants [Gré97b]. correct [Jon90, LL97]. Correcting [TJ97]. correlates [Pep90, WRT95]. correlation [KHK94, Spa96]. Correlations [BSMK99, LP91a]. cost [MH98a]. could [BCJ^{+}97]. Coulter [TZD^{+}95]. counting [MK96, MC98, TZD^{+}95]. countries [Sut99]. counts [Hul95, RSB99, vdML95]. coupling [BJM99]. covariation [MK99b]. cover [CC93]. CPUE [Lar92]. cpues [KZ97]. crab [AL91, GGF94, Hog93, Mun97]. Crangon [Bod96, OHN99]. Crassostrea [GH91]. criss [vdML95]. criss-cross [vdML95]. criteria [Eve93]. criterion [Coo98]. croaker [APD92]. croissance [BB94]. Cross [Mac99, RTWC97, Foo99, vdML95]. Cross-scale [Mac99]. cross-section [Foo99]. Cross-shelf [RTWC97]. cruise [PBSN98]. crustacean [VR98]. ctenophores [Mil95]. Cu [SKL98]. culture [Rym97]. cultured [CLAD97, CAWD97, EHJN91, Jon97, LGKM97, McV97, OMF^{+}97]. Current [GD96, RGHC96, SHB96, Swa97]. currents [RdC95]. curve [Spa95]. curves [Pet98]. Cuv [Nie94]. Cyanea [BMM94]. cycle [SSH95, WW96]. Cyclopoida [FG95b]. cylindrical [Foo95]. cytochrome [BSMK99].

d [KCB99]. dab [FAN99]. dactylopterus [HHB96, KCB99]. Daily [CHH91, HCA93, LHSA98, MF92, hLHM^{+}92, Mok92b, NW93, TA94, Wal94, ZF98]. dams [RSN96]. Danish [KL97, LV94]. Daphnia [SG98]. Data [BBBR98, Bea91, Bel98, CR93, Gud95, Has96, Hee96, HN91a, HN91b, HL97, Kap99, Lar92, LBDH98, Lor91, MM96, Mal90, MC90, Mil92, MUV90, Moh99, Mur96, MBHQ96, NLCS99, PL97, PBSN98, PV93, RP95, SN91, She99, SWB96, Ste96, SS99, WRT95, WB96, vLRvR^{+}96]. database [HL97]. date [CM91, MF92]. Davis [Chr95, OB97]. day [MLCB99, WH98]. deadzone [OM96]. début [AL95]. decadal [MM97, TSP^{+}96]. decade [Ham96b]. decay [MT97b]. December [Dic94, Hem95]. decision [LS98a]. decision-making [LS98a]. decline [HJR96]. declines [VHV^{+}98]. deconvolution [MMHC98]. decreasing [VHV^{+}98]. Deep [Alb94a, Alb94b, EHEW94, HN90, Klo96, MJ97b]. deep-towed [Klo96]. deep-water [HN90]. deepwater [KWK97]. defence [GHM97]. defining [PL92]. Definition [FS93]. definitive [VGHO97]. deflecting [SE97]. Delaroche [KCB99]. delineation [RSB99]. Delphinapterus [HJR96]. delta [Ste96, OR97]. demersal [AH95, EJW98, LR99, Phi98, PR98, RvLDH96, RM96b, SFCP99]. Demography [HRP90, SH95]. densely [AB96]. densities [TL96]. density [Agl96, BR97, DB92, FS93, GWE99, HGOT99, HBC^{+}92, HP99, HHHS92, IMH96, Mis93, MAF95, RvL96, Sam90, TCA97]. density-dependent [GWE99]. dependence [CSW92, MI96, Sam90]. dependent [CHR96, GWE99, KN99, Pet98]. depleted [Bes93]. Depletion [LML99]. deposits [SN99]. depredation [CC93]. depressed [CR99]. deprivation [BBG91]. Depth [MI96, HL97, SS99]. Deriso [Hor92a]. derived [BMS98, HS91, JM99, RW91, Som99, SVJ99]. describe [DJ92]. Design [Bar98, Bas99, KO97]. designed [CC93]. Designing [Foo90a]. Desmarest [APD92]. detect [FN93]. detectability [LR99]. detection [ROH^{+}98, YCL96]. determinant [FL98]. Determination [MK95, TL96, BHS96, Ned90, RCE^{+}93, SP92]. determine [MSG96]. determined [MT97b, RCE^{+}93, UC92]. determining [Foo91, PM96a]. developed [WK99]. Developing [Nie99]. Development [GDJC99, BB94, BWH97, CCW99, DW99, MC98, PM96a, Rad98, Rai93, dG96b]. diadromous [JWF99]. diadromy [McD99]. diameter [RAFC92]. Diamond [Nak90, PS97]. diatom [MGN91]. Diel [Alb95, BWH97, FGS96, MT97b, MGF96, PB97, dPVV^{+}96, ML96]. Diet [Arr96, CDL^{+}94, HGLN95, PTCC98, SV98, Cor99b, GMA98, Gui95, HW97, HRBA91, JJ96a, OVP98]. Dietary [FG96]. diets [KLF96, MM97]. differ [JG93]. difference [Hor92a]. Differences [AS95, AS97, Hou94, Mok92a, OR97]. Different [McD99, Kai98, KH95, KLF96, LML99, Løk98, PPS90a]. differential [Hor92a]. differentiation [SK95b]. digested [MT97b]. digestive [FVC94]. dimensional [MJ96, Ped96b, WT96]. dimensions [MABD92]. dinoflagellate [Rai93]. diploid [OMF^{+}97]. dipterygia [BMM^{+}98]. Direct [BH92, AR90, RR98]. direction [YF96]. discarded [OVP98, SFCP99]. discards [OR97, SPCP97, SFCP99, WB97a]. discharge [Bod96]. discriminant [SAC96, SK95b]. discriminate [Hul95]. discriminating [SBSH97]. Discrimination [Ped96b, RVD90, SWCG94]. discussion [Som99]. Disease [McV97]. Diseases [LM99, ML99, LMW^{+}99, Möl90, MA92b]. disjuncta [VVH91]. disk [BS98]. Dispersal [BDC^{+}94, SS93b]. dispersion [KHK94]. disposal [RRL^{+}92]. dissolved [Laa96, SKL98]. Distinguishing [MF91, HIL98]. Distribution [Ber93, BI92, Cas99, CCC^{+}98, FHA^{+}98, FS97, MJ97a, Mut94, Nie94, NTWW95b, PBSN98, RDHP97, SPL98, WHM97, Agl96, AEH^{+}99, AH92, BMS98, CG97, CBG96, CHØ92, CCS97, CNF90, FG95a, Gar97, HRP90, HP99, HS98, HF94, Hun97, HO96, HHHS92, KO97, KCB99, KWN95, KBC^{+}92, LMW^{+}99, LSP^{+}95, LLLL99, Mac99, MKP96, McD98, MP91, MF92, OF95, O’B99, Orl99, OMN98a, OMN98b, Ric95, RBSV98, RM96b, SVS99, SSKP94, Swa97, Wan97, WH97, dPVV^{+}96]. distributions [DK96, GRDCW98, OBD^{+}97, Pet97, Pet98, RGHC96, Sak97, TL96, WMO98]. disturbance [Kai98]. disturbances [Kim97]. Diurnal [ES92, KN99]. divergence [FE97]. diversity [BCJ^{+}97, GR98b, HBC^{+}92, JPG93, Nie99, RG96, SPCP97, VM98, Ver97]. DNA [BB94, BBG91, BRK97a, McG97, PES99, RG93, UC92]. Do [BH98, LHHN98]. does [MK99a, PTO95]. dogfish [OVP98]. dolphins [NTWW95b]. dome [Jen95]. dome-shaped [Jen95]. domestication [FE97]. don’t [MBT95]. Doppler [GD96, RGHC96]. d’Orbigny [HRP90, LHSA98, LDU98]. Dover [HN91a, HN91b]. down [VR98]. Downs [Cus92]. Dr [Bev91a, Har95]. dredge [FB99, GCM99]. dredger [CWC^{+}97]. dredging [RR98, SN99]. drift [Jen99]. Drifting [HGDC94, BB92]. drilling [DM96]. dry [BRK97a]. Dual [TF96, ET96, Mut96, SW96]. Dual-beam [TF96, ET96, Mut96, SW96]. Dublin [Ano94c]. due [FE97, LP91a]. dumerili [RG93]. duration [MLCB99]. during [Atk95, BMS98, DvWC96, Fos96, Gud98, HM95, HRP90, HGLN95, Hee96, HP99, JJ96a, MA92a, Nak90, NHP^{+}95, Nøt98, OB97, PM96a, PTO95, RA97, Thu97, vdWRvdLF90]. Dutch [Bod96, BC96, DM96, Laa96, dVDP^{+}98]. dynamic [FG95a]. Dynamics [Hou90, Mis93, BW95, FG95a, GBB^{+}97, Hea95, HW90, HN95, Mac99, OHN99, OLO98, Pet98, RP95, RM96a, SB94, SC99, Smi95, VGHO97, Ynd99a].

early [BI92, HP99]. Earth [Ynd99b, Ynd99a]. east [AL91, AH92, BTC93, CZM98, HTG97, HGLN95, JG93, MEH97, OMN98a, OMN98b, RdC95, WHG98, Jak92, LLLL99, MJ97a, NN91a, TIR97, TSP^{+}96, Jør92, MON98]. Eastern [MP99, Swa97, AS95, Ber93, Gar97, MP91, MH98b, Phi98, PR98, PS99, SSH95, WHM97, KLF96, MBHQ96]. Ebro [OR97]. Echo [Knu90, GD96, MAF95, Mis96, MFPT98, MK96, MC98, Ped96b, SFG96, SBH95, SHB96, TF96, WB96]. echo-counting [MK96]. echo-sounder [GD96, Ped96b, SFG96, SBH95, SHB96]. echoes [RWJF90, Rey96, SBSH97]. echogram [PL96]. echosounder [BJM99, SBSH97]. echosounders [WRT95]. Ecological [Hay98, HPDR93, Jen91, Jon97, Rad98, SS93a, dVDP^{+}98]. Ecology [Alb94b, HA92b, HA92a, Alb94a, Cus95b, Eco91, HTG97, HGD92, McD98, SSS98]. Ecosystem [Ano93b, GR98a, RGH93, BBBR98, BE90, Hea95, JG93, KSP96, SSJK98]. Ecosystems [Ano94c, Gre95, Mil95]. eddies [Kim97]. Editor [Fra93]. Editorial [BR92]. Editors [Bra99]. edulis [KL97, RCE^{+}93]. eel [Køi91, Vøl90]. eels [MA99]. Effect [GCM99, ROH^{+}98, Bra95, Lam90, SPCP97, SVJ99, WM99, vdWRvdLF90, BBG91]. effective [SSS99]. Effects [AY99, CG95, GR98a, GKT98, HFP96, HS98, Sin91, Ste99, VVH91, AR90, AGKB91, BH92, BFNP97, DBH98, Daw98, FRK93, GRN96, Mil92, MLCB99, MRN94, PM96b, RAFC92, RR98, Roc98, Rym97, Wan97, WK94, Ye98]. Effet [BBG91]. efficiency [FB99, Gré97b, Som99]. effort [Gré97b, RvL96, RBSV98]. Egg [Hay95, KB95, LN93, Ham96b, ML95, MBR95, Pet97, PHT95, RJ90, ZF98, vdWRvdLF90]. egg-production [Ham96b]. eggs [CCS97, CNF90, FL98, FVC94, HGDC94, MGN91, Nie94, RJ90, Tor91, WH97]. Egidius [Øst90]. EK500 [GD96]. elassodon [PPS92]. Elbe [Möl90]. electron [ARW93]. electronic [CLB92]. electrophoresis [NN91b]. Electrophoretic [BCCP96, SJB90]. element [Par99]. elemental [HCR94]. Eleven [CFUCY95, Foo90b]. eleven-frequency [Foo90b]. ellipse [CB93]. elongatus [SKHM98]. embryonic [PM96a]. embryos [DvWC96]. emigration [LHSA98]. Emmy [Øst90]. Empirical [BHS96]. encrasicolus [BCCP96, JPG93, Nie94]. end [RAFC92]. endurance [WHW99]. energetic [Wan97]. energetics [HTG97]. Energy [PPS90a, PPS92, APD92, AH96, Arr98]. enflata [KLF96]. England [AL91, RRL^{+}92, RW91, TIR97]. English [GK97, Laa96, RPW^{+}99, RM96b, WHM97]. Engraulis [BP91, BCCP96, JPG93, Nie94, Wal94]. enhance [Dem99]. Enhancement [KS90, SJK90, SK90a, SK90b, Rit97]. enrichment [RGH93]. Ensis [RR98]. enterprises [Bia99]. entry [FA97]. Environment [MW97, TR97, BED^{+}93, BMN97, Col98, FRK93, FC97, Hog94, KB95, MB96a, Md95, McG97, PLvdM^{+}96, Ano93b]. Environmental [AG98, WW96, Bas99, Eve93, Jør92, LP91a, Mil92, Möl90, SSKP94, VM98, WRT95, Wan97]. environments [Fog93]. enzyme [UC92]. enzymes [BSMK99]. Epidemiology [MA92b]. Epinephelus [RG93]. equal [Foo95]. equal-area [Foo95]. equation [Jen90]. equatorial [RA97]. equivalence [ML95]. equivalent [Rey98]. Erie [SS95b]. Erin [OHN99]. Erratum [Ano97a, Ano97b, Ano97c, Ano97d, Ano97e, Ano98a, JJ96b, NTWW95a]. errors [LP91a, LDU98]. Escaped [SHKL97, HJL99, JG97, McG97]. escapees [Chr95, MT97a]. Escapement [Nak90, FA97, WM99]. escaping [SPCLT96]. esmarki [Alb94a]. esmarkii [BWH97]. essai [BB94]. essential [Bia99]. Establishment [Neh98, FPW91, PR98]. estimate [HLAM95, JS90, Lar92, LS97, RKCM92, Spa90a, Spa95, ZF98]. estimated [LL97, MRL92, SSFB97]. estimates [Agl96, BH94, CM91, GKT98, GG98, Ham96b, HJL99, HS91, Hay95, HJ99, JM99, JCP91, Mis96, MB96b, PCM96, RSN96, Som99, SFG96, SBH95, Spa90b, Spa90c, Sta94, SFCP99, WA90]. Estimating [KHK94, BG94, BC96, FS93, hLHM^{+}92, Mal90, PK95a, Spa90c]. Estimation [Has96, MH97, RP98, Thu97, HGOT99, Jen90, Kap99, KCB99, Pet97, TA94]. estimator [HN91a, HN91b]. Estimators [SFCP99, Far92]. Estonian [ML99, OLO98]. estuaries [CZM98, HFP96]. estuarine [Lor91, TZD^{+}95]. estuary [CHH91, Gay98, Möl90]. état [BB94]. Euchaeta [Mau95]. Euphausia [MRSF91, MAI96]. Europe [BCJ^{+}97]. European [vLRvR^{+}96, AH95, Bal93, FPW91, FRK93, Gui95, Køi91, MA99, RCE^{+}93, Vøl90, vdML95]. eutrophication [ADS96, GRN96, RvL96]. evacuation [DJ92, Han96, TA94]. evaluate [MSRK99]. Evaluation [CLB92, CCW99, Chr97, Dem99, Gle96, ODR92, Rey96, SBSP99, BP99, CG97, KOS^{+}99, MSC98, MAC96, MK99a, PKMP99, PM96b, PS99, SSS99, BB94]. events [MT97a]. Evidence [HJR96, SBH95, ML96, MBHQ96, Sut99]. evolutionary [Nie99]. ex [NL99]. examination [CM91, HLB91, JCP91]. examinations [BLM99]. example [MK98, SF96]. examples [Sma98]. excluded [SE97]. exhibit [LHHN98]. existing [vLRvR^{+}96]. expectation [HGOT99]. Experiences [BP99, Sut99]. experiment [CHR96, SHM95]. Experimental [FE97, BB94, FB99, Gre95, Ims99, Mut96]. expérimentale [BB94]. experimentally [Som99]. experiments [Ehr91, God95, Mil94, SPCP97, THJ^{+}95, dVDP^{+}98]. Exploitation [OR97, CHØ92, Coo98, GR98b, RM96a]. exploited [ML96, SC96]. exploration [Bel98, Fra97]. export [BFKR97, LPL95]. expose [ZM93]. Extended [She99]. extension [RAFC92]. Extensive [Nøt98]. Extinction [Foo99, JWF99, LML99, WK99]. extraction [dG96a].

Factors [RM96b, Bas99, BEM96, Daw91, LP91a, Mac98, NBBS95, SSKP94, WW96]. Falkland [Hat96, HM99, Hog93]. Far [FH96]. Far-field [FH96]. farm [HRL97]. farmed [BFNP97, FE97, HJL99, JDLH97, JG97, McG97, SHKL97]. farming [DLMS97]. Farne [EHEW94]. Faroe [GHH98, Mar93]. Faroese [HJL99]. fasciatus [MBSM90]. fate [EHEW94]. features [GRDCW98, MRTB95, RDHP97, RA97]. February [Øst90]. Fecundity [Hor93, MBR95, LN93, hLHM^{+}92, Rij91, Ser90]. fed [DJ92, PPS90a]. feed [HJB90]. feed-back [HJB90]. Feeding [PR93, Ped95, Alb95, Atk95, BWH97, CP96, Eco91, FPHM90, Fos96, HTG97, HH93, HKN^{+}91, HGD92, Jen95, Mac99, MK95, OB97, PB97, Rob92, SV98, SC99, TZD^{+}95, THJ^{+}95, Wan97]. female [Rij91]. females [LN93]. ferries [ROH^{+}98]. Fertilization [HCH91]. fiathead [PPS92]. Field [Hay95, LDU98, NW93, FH96, PES99, TA94, THJ^{+}95]. fields [KO97]. filter [MK95]. filter-feeding [MK95]. fimbria [KZ97]. fin [Sam90]. final [PM96a]. Finland [KRKS97]. finmarchicus [BHKH94, GA98, PTO95]. finned [AS95, AS97, GCG96]. Finnmark [FPHM90]. fins [PM96a]. First [BG93, Daw91, Fos96]. first-feeding [Fos96]. first-year [Daw91]. Firth [GMA98, RCM90, TCA97, WHG98]. Fish [Hou90, LM99, MW97, PK95b, SS99, SFCP99, UJA^{+}93, YCL96, Agl96, AEH^{+}99, AY99, AH95, BP90, BP91, BHS96, BB94, BCJ^{+}97, CBG96, CLB92, CCS97, CR93, CvdK96, CHR96, Cus95a, DvWC96, DK96, FG96, FVC94, Foo91, FS93, GR98b, Gis99, Gle96, GBB^{+}97, GMA98, HM95, HRL97, HG96, HS91, HGOT99, HD96a, Hee96, HBA98, HJ99, Hou94, Klo96, KWK97, LP91a, LR99, LR96, hLHM^{+}92, Man93, ML96, MKP96, MAYC96, MAC96, MA92a, MAF95, Mis96, MLCB99, MA92b, MFL98, Orl99, Par99, PK95a, Pat98, PCH91, Ped96b, Pep90, PES99, Pet93, PL96, Pet98, PS97, PLvdM^{+}96, Phi98, PR98, RWJF90, Rey96, RG96, RvLDH96, RM96b, Rym97, Sam91, SV98, SDW^{+}96, SKD96, SFG96, Spa90a, Spa90b, SC96, SK97, SJ96, SK95b, Swa97, TL96]. fish [Thu97, Tur92, WA90, WTB99, WHW99, YF96, ZMHD96]. fished [PBSN98]. Fisheries [LS99, MH96, Oel99, Sim96, SBSP99, AL91, Bax98, Bia99, Bod96, BR93, CCW99, Coc99, CC93, DM99, ET96, EJW98, Fow99, GDJC99, Hat96, HS99, HJB90, Jen91, LS98a, LV94, MH98a, MS90, MB97, NR96, NLCS99, ODR92, Pay99, PL97, Phi98, PM96b, RP99, RP95, SSS99, War94, WT96, Ye98, YWMW97]. Fisheries-Management [SBSP99, LS99, CCW99, SSS99]. Fishery [Chr97, CR97, Chr95, Coo99, DW99, EGB^{+}96, EHEW94, HJL99, Hat96, Jen99, KSP96, MK99a, MSRK99, Moh99, PKMP99, PH94, SRAL99, Sut99, SNSC94, WK94]. fishery-management [Coo99, MK99a, MSRK99, Sut99]. fishes [JWF99, Roc98, Ser90, SW96, Ste99]. Fishing [GR98a, MSPG92, Sam91, Bia99, CC99, Ehr91, Gar97, KZ97, MRN94, Pay99, Roc98]. FishTV [MJ96]. fitting [Mil92]. five [Atk95, Eco91]. fixation [vdWRvdLF90]. fjord [KL97, KS90, LBDH98, MH98b]. fjordic [ADS96]. flat [RCE^{+}93]. flatfish [SKD96]. fleet [MS90]. fleets [CC99]. Flemish [CP96, SRJ98]. flesus [BSMK99, BLM99, DKLM99, Køi99, LMW^{+}99, Möl90]. flexible [MH98a]. flounder [BSMK99, BLM99, CZM98, DKLM99, Køi99, LMW^{+}99, MUV90, Möl90]. flow [HL97]. fluctuation [BEM96]. fluctuations [PLvdM^{+}96, SKS97, Ser90]. fluorosensor [Bar98]. flux [HP99, LPL95]. fluxes [LOÅ97]. following [PR98]. Food [Arr98, Eco91, MW97, MRL92, NHP^{+}95, Alb95, Arr96, BBG91, BWH97, FG96, GBB^{+}97, HGLN95, HRBA91, KB95, LS97, Man93, MT97b, Mun97, OB97, PPS90b, Ser90, Wit96, dSJ95, SH95]. foraging [Gré97b, HTG97]. forbesi [CDL^{+}94, CBR95, Col99, Has96, PBSN98]. forecasting [She91]. forecasts [CKF91, SN91]. Formation [BFKR97]. Formulating [MSRK99]. Forth [WHG98]. forty [Hee96]. forward [YF96]. four [MBR95, PS97]. fractionated [MFL98]. frame [HD90]. framework [LS98a, LS99]. France [DLMS97, FPW91, GH91]. Fraser [LP91b]. French [AL95, Bel98, BBG91, BB94]. frequencies [Foo91, HP95, HJ99, NP98]. Frequency [CSW92, YWMW97, Foo90b, HJ99, LR96, MAC96, MSG96, ROH^{+}98, RBSV98, TL96]. freshwater [HPDR93, Hou94, McG97]. FRG [Möl90]. front [BR97]. frontal [GK97, Kim97]. fronts [Fra97]. FRS [Bev91a]. fulmars [CG97, HTG97]. Fulmarus [CG97]. function [AW92, LGKM97, MH98a]. functioning [Bia99]. functions [MT97b]. Fundy [SOBP99]. furnieri [APD92]. future [GB97, IOEJ97, MH96].

gadoid [Eco91, EJW98, His96]. gadoids [Alb95, GW93, GKT98, MGF96, SE97, SFCP99]. Gadus [AW92, AWEH94, BH98, Bra95, BWH97, CP96, Cas99, DJ92, FG96, FØK94, Han96, HP99, HGD92, Jør90, Jør92, KWN95, ML99, MON98, MT96, NL99, NP98, PU93, PLPN90, RdC95, RP96, SRJ98, SC99, SSFB97, SP96, SMGB97, SS93b, SS96, SNSC94, TA94, WH97, dBTT98]. gahi [HRP90, Hat96, HM99]. Galician [GGFFF98, GCG96]. gall [Rob92]. Gamble [Har95]. gammarus [AL91]. gannet [MM97]. gas [Nøt98, dG96b]. Gascogne [AL95]. gastric [DJ92, Han96, TA94]. Gear [Mac95]. gears [Mil94]. Gelatinous [Ano94c, HGDC94]. gemfish [PS99]. generations [Rym97]. Genetic [DMAG97, McG97, MS96, NN91a, NJN94, RR99, Ver97, FE97, GB97, Nie99]. genetics [Rym97]. genotypic [Ims99, RVD90]. genus [LN93, NN91b]. geographic [BEM96, HS98, McD98, MB97, dBH95]. Geographical [Box97, DKLM99, Wan97]. geometric [MABD92, SS99]. geometry [Foo90a]. Georges [MBSM90, PV93, PLPN90]. geostatistical [Mur96, Pet97, Pet98, WT96]. Geostatistics [Pet93, MSC98]. gerlachei [SSH95]. German [Gre94, GRN96, HCR94, Hic98, Rad98]. gigas [GH91]. Gilbert [PPS92]. gill [Køi91]. gillnet [MH97]. gillnets [TBHL97]. GIS [CC99]. glacialis [CG97]. GLM [Ste96]. global [Mil95]. Globicephala [AS95, AS97]. GMDH [Cla94]. Gobiidae [HCA93]. going [LM99]. golfe [AL95]. gradient [SS99]. gradients [Orl99]. Grand [SCM94]. granulosa [Hog93]. Grapsidae [McD98]. Gray [NTWW95b]. grazing [LPL95, NBBS95]. Great [SS95a, Gré97b]. greater [Ber93, RG93]. Greenland [BRK97b, Chr97, HRL97, HJR96, MJ97a, MJ97b, MB97, NN91a, NJN94, OB97, PR93, Ped95, PK95b, RSB99, Ric95]. grid [SE97, SPCLT96]. groenlandica [HKN^{+}91, KHØ95, LHHN98, MS96, NHP^{+}95]. grossly [LMW^{+}99]. ground [HBA98, Tor91]. Groundfish [SOBP99, Ste96]. grounds [NK99, PK95b]. group [Arr96, BWH97, FG96, GG98, HH93, MJ97a, NN91b, PLPN90]. grouper [RG93]. groups [NFM98, SVS99, SCW96]. Growth [APD92, AH96, FØK94, GA94, GGFFF98, HH93, MBSM90, MON98, SSFB97, SVJ99, TCA97, Wit96, BB92, BB94, Ber93, Bra95, Cla94, DB92, Daw91, DAP97, FAN99, FRK93, GCG96, HLAM95, HN90, JCP91, Jon90, Jør92, KCB99, KHØ95, LPL95, MRSF91, ML95, Mil92, Mil96, MW96, MF92, Mok92a, PPS92, RCE^{+}93, RvL96, SS93b, SS96, Tor90, VPT99, Wal94, NBBS95]. Growth-maturation [MBSM90]. guidelines [DM99]. guillemots [WB97b]. Gulf [ZF98, BS92, Cas99, CFUCY95, CR97, GKT98, HSFAM99, MBSM90, MUCM98, RDHP97, RG93, SC99, SNSC94, Tay95, VGF97]. Gulland [Bev91a]. Gullmar [LBDH98]. gulls [CR97, Mun97]. Gunnar [Øst99]. Gyre [Ric95]. Gyrodinium [Rai93].

habitat [EMN^{+}94, Gre97a, McD98, RA97]. habitats [BP90, BP91, Kai98]. habits [CG97, HKN^{+}91, NHP^{+}95, PR93, Ped95, SV98]. haddock [Alb94b, BWH97, Ehr91, ES92, Hor91, Jak92, JJ96a, JS90, Pat98, PV93, PM96b, SK95b]. haem [BSMK99]. Haemoglobin [NP98]. hake [Gui95, Ham96a, LCP96, PH94]. hakes [PB97]. halibut [AGH^{+}93, BG93, BRK97b, MB97, OB97, PR93, RSB99, SP92]. Hamburg [Har97]. HamSOM [Har97]. hand [Spa95]. Harbour [TBHL97, GS96, NTWW95b]. hard [RW91]. hard-shell [RW91]. harengus [KRKS97, Raj93, Arr96, AH96, BMS98, BH98, Fos96, Gry99, Hog94, HS98, JB91, Mac99, MF91, MF92, Mok92a, Mok92b, Nøt98, RCM90, SPCLT96, Tor91, ZM93, dBH95, dBTT98, dBT98]. Harp [HKN^{+}91, KHØ95, LHHN98, MS96, NHP^{+}95, UJA^{+}93]. Harvest [PS99, Coo99, DW99]. hatch [CM91]. Hatchery [BMN97, HIL98, SL97]. hatchery-reared [HIL98]. hatching [MF92, RCM90]. hatching-date [MF92]. haul [Fry91]. Hauraki [ZF98]. Hawaiian [DW99]. heavy [Lor91]. heavy-metal [Lor91]. held [Ano94c, HA92b, HA92a]. Helicolenus [HHB96, KCB99]. Helminth [BRK97b]. Hematodinium [FHA^{+}98]. Hemigrapsus [McD98]. Henry [LP91b]. Henslow [Mun97]. henslowii [Mun97]. herbivorous [KH95]. herring [Arr96, AH96, BMS98, BH98, CR97, Cor99a, Cus92, DRR94, Foo99, Fos96, Gry99, HD90, Hog94, HBA98, HS98, Hul95, HO96, JS99, JB91, KRKS97, KSP96, Lam90, Mac99, Mis93, MB96b, MFPT98, MF91, MF92, Mok92a, Mok92b, MGN91, Nøt98, PMF^{+}96, Raj93, RCM90, SF96, Spa90c, SRAL99, SPCLT96, Tor90, Tor91, UC92, WWS93, WW96, ZM93, dBH95, dBTT98, dBT98]. High [SL97, HP95, HJ99, MAF95, MFPT98, PTO95, PMF^{+}96, ROH^{+}98]. high-frequency [ROH^{+}98]. high-resolution [MAF95, PMF^{+}96]. Highlights [OLO98]. Hippoglossoides [FAN99, PPS92, WHW99, BRK97b, MB97, OB97, PR93, RSB99]. Hippoglossus [AGH^{+}93]. Hippoglussus [BG93]. Hirtshals [DSO96]. Histological [SRJ98]. Histopathological [BLM99]. Histories [Sin91, SVJ99]. history [CR99, Cus92, Hog93, JB91, Mau95, Nie99, OIF95, PN93, Rob92, Roc98, VVH91]. Homarus [AL91]. Hong [Par99]. hook [MH97]. Hoplostethus [KWK97]. horizontal [MFPT98]. horizontally [SJ96]. horse [BH94, HHA95]. Host [Mac98, NLP97, VGHO97]. Hulbert [Rai93]. hull [Rey98]. hull-mounted [Rey98]. hungry [OBD^{+}97]. hybrid [McG97, MS90]. Hydroacoustic [NL99, DK96, RSN96]. hydroacoustics [SW96]. Hydrobates [vdML95]. hydrocarbons [SK97]. hydrodynamic [SE96]. hydrographic [DSO96, Hic98, SP96]. Hydrography [TSP^{+}96, CvdK96, Gar97, RGHC96, Wit96]. hydrography-related [Wit96]. hydropower [RSN96]. hydrostatic [AW92]. hypotheses [Tur92]. hypothesis [BHKH94, Ims99, NP98].

Iceland [AG98, DMAG97, JS99, LS97, MJ97b]. Icelandic [AG95, BDS96, GA98, Gud98, NJN94, TJ97]. ICES [Bal93, BED^{+}93, KOS^{+}99, LM99, Pat98]. Ichthyoplankton [HF94, OF95, GK97, LSP^{+}95]. ideal [Ham96a]. Identification [LV94, NN91b, RWJF90, dBH95, Bea91, HG96, SDW^{+}96, SAC96, ZMHD96]. identifying [Daw91]. II [Smi95, SK90a]. III [KS90, SH95]. Illex [BI92, GCG96, GMG98, LN93]. image [GMG98, TZD^{+}95]. imaging [MJ96]. immune [BGA97]. Impact [Agl96, Kin92, Ano94c, Chr95, DM96, EHJN91, GHM97, HSFAM99, HS99, McG97, OIF95, Phi98, RBSV98, RGH93, dG96a]. impacts [KSP96, ODR92, PM96b]. Implementation [SBSP99, BP99, LS99]. Implementing [SSS99, KOS^{+}99]. Implications [Hou90, Orl99, SN99, GBB^{+}97, Hou94, JDLH97, MGMS94, McV97, MM97, MB97, PH94, SP96]. Importance [OVP98, Bas99, GMA98, Jon90, LR99]. important [Mac98, Mun97, Spa90b]. Improved [Klo96, Mis96, Spa90b, Foo90a, MWRH98, She99]. Improvement [Coo99, GB97, NLCS99]. improvements [SHB96]. Imsa [Vøl90]. in-season [SRAL99]. incidence [HRL97, HJL99, SL97]. incidents [FN93]. included [Spa90b]. including [WRT95]. incomplete [CR93]. incorporates [VPT99]. Incorporating [Som99, KHK94, SC96]. Increase [Bes93, VR98]. increased [KSP96]. Increasing [MH98a, Coc99]. increment [Mil92]. increments [Mok92b, Wal94]. incubation [vdWRvdLF90]. Index [Ano90a, Ano91a, Ano91b, Ano92a, Ano93a, Ano94a, Ano95b, Ano96b, Ano97f, Ano98b, Ano99a, BRK97a, Gle96, Rad98, She97]. indicators [Neh98, NLP97]. indices [BS92, ES92, KZ97, MJ97a, MGF96, She99]. indigenous [KS90]. indirect [HGOT99, RR98]. Individual [BW95, CSW92, CHR96, FA97, FAN99, MJ96, SK95a, VPT99]. Individual-based [BW95, CHR96]. induced [BEM96, Rey96]. industrial [CR93, HHA95]. industry [Pay99, dG96b]. inevitability [OW95]. infauna [KBC^{+}92]. infaunal [BFNP97]. infection [FHA^{+}98]. infections [DPHM97, Daw98, Mac98]. inferences [LPL95]. inferred [SS93b]. infestation [JG97]. infestations [JDLH97]. inflow [BME96]. Influence [BH94, Gar97, MW97, BH98, ES92, GK97, HGDC94, MGF96, Sam91, SFG96, Ynd99b, Ynd99a]. influences [Jør92, MM97]. information [Ped96b, RKCM92, Spa90c, Wan97]. ingestion [BGA97, PPS90a]. initial [PKMP99, RDHP97, SVJ99]. Inlet [CWC^{+}97, TT96]. inner [MH98b]. inorganic [Laa96]. inputs [DBH98]. inshore [CC99, RM96b]. installations [Kin92]. Instantaneous [KV97]. Instrumental [PL97]. intake [Arr98, MT97b, PPS90b]. integrated [MWRH98, RVD90, Spa90c, vLRvR^{+}96]. integrating [Wan97]. integration [MAF95, Mis96, Rey96, TF96, WB96]. Integrator [Knu90]. intensity [CC99, MFPT98]. inter [MUCM98, PCM96, dBTT98]. inter-annual [MUCM98, PCM96]. inter-cohort [dBTT98]. interaction [Jen91, PV93]. interactions [BTS^{+}97, Jon97, MBSM90, Smi95]. Interannual [PCH91]. Intercalibration [BMM^{+}98]. Intermediate [MW97, Mar93]. Intermediate-Scale [MW97]. intermoult [GGFFF98]. International [Hee96]. Interpretation [MM96]. Interrelationships [KS90]. Intraspecific [JB91]. Introduction [DRP96, GH91, HWY97, Hou90, RSH95, SBSP99, Wal99, Knu90, LM99]. Introductions [Sin91, BCJ^{+}97, Hay98]. invader [Mut94]. invading [UJA^{+}93]. invasion [Gre94, HHB96]. invasions [HKN^{+}91]. invertebrates [Phi98]. investigated [MT97b]. Investigating [GWE99]. Investigation [CC93, Moh99, PHT95, SSD96]. investment [Gré97b]. Ireland [CCC^{+}98, Daw98, MRTB95, OF95]. Irish [EN97, OHN99, BR97, BC96, CDL^{+}94, CBR95, CCS97, Laa96, PHT95, Rai93]. Irminger [MJ97a]. isada [MAI96]. islandica [WK94, Wit96]. Islands [MRSF91, Swa99, DB92, DW99, Hat96, HM99, Hog93, MBR95, OR97]. Isle [OHN99, EN97]. Isles [Bra94, NTWW95b]. isolated [HW90]. isotope [AS97]. isotopes [AS95]. ISVPA [KV97]. ITCH} [Gre95]. iterated [Far92]. IV [Cus95a, SK90b]. ivermectin [BFNP97]. IX [Nak90]. IYGPT [PLPN90].

jack [LCP96]. jacobaeus [HSFAM99]. James [LP91b]. January [Hem95]. Japanese [FPW91, GH91, TSAS96]. japonicus [SKS97]. jellyfish [Mut96]. jellyfishes [BMM94]. John [Bev91a, Har95]. Jordan [PPS92]. Jubilee [BRS96]. July [Ano96e, Øst99]. June [Bev91a, Bla93, Dan97]. juvenile [APD92, BS92, HCA93, HP99, HS98, Ims99, NL99, PU93, SS93b, SS96, dBTT98, dBT98]. juvenile-to-adult [Ims99]. juveniles [BDC^{+}94].

Kattegat [Dan97, NR96]. Kerguelen [DB92]. key [BEM96]. Kiel [SB94]. kinds [McD99]. kisutch [BTS^{+}97]. kittiwake [HW97]. km [RWJF90]. knowledge [Ull96]. Knud [Mad94]. Kokanee [MI96]. Kokemäenjoki [SE98]. Kondakovia [JCP91]. Kong [Par99]. krill [HD96b, MGMS94, MAI96, Mur96]. Kromme [LHSA98]. Krøyer [DPHM97]. Kullenberg [WO93]. Kuroshio [Kim97].

L. [AL91, Alb94b, AL95, AGH^{+}93, AW92, AWEH94, Arr96, AH96, Arr98, BMS98, BDC^{+}94, BBG91, BB94, BG93, BH98, BMM^{+}98, Bod96, BLM99, BTC93, Box97, Bra95, BWH97, CG95, Cas99, CHH91, CNF90, DMAG97, DB92, Daw91, DPHM97, Daw98, DJ92, EN97, FB99, FL98, FG96, FØK94, Fos96, HSFAM99, HH93, HRL97, HJL99, HP99, HW90, Hog94, HS98, HGD92, HN91a, HN91b, Hor93, HCH91, JG97, JB91, JJ96a, Jør90, JPG93, KOS^{+}99, Køi99, KL97, Mac90, MLBM98, Mac99, MSC98, MA99, McG97, MON98, MFGP96, MF91, MF92, Mok92a, Mok92b, NW93, NL99, NLP97, NP98, NTWW95b, ODR92, OMF^{+}97, OHN99, PM96a, Raj93, RCM90, Rij91, RP95, RM96a, Rit97, RdC95, SRJ98, SPCLT96, TA94, TIR97, Tor91, TBHL97, TCA97, VGHO97]. L. [VPT99, Ver97, WK94, Wit96, WH97, dBH95, dBTT98, dBT98, vdWRvdLF90]. laboratory [MT97b, THJ^{+}95]. laboratory-reared [THJ^{+}95]. Labrador [KWN95, ODR92]. Lacepède [NTWW95b]. lactation [NHP^{+}95]. Lagenorhynchus [NTWW95b]. lagoon [DLMS97]. Lake [SS95b, Jen91, Hog94, SG98]. Lakes [SS95a]. land [KS90]. land-locked [KS90]. landings [LL97]. Lantau [Par99]. Large [AH92, ARW93, BMN97, CHØ92, Kap99]. Large-scale [AH92, BMN97]. Larus [Mun97]. larvae [BBG91, BB94, BRK97a, BG93, CCS97, CHR96, Eco91, FPHM90, FØK94, Fos96, HCR94, HD90, Hou94, HGDC94, MF91, Mok92a, MFL98, Nie94, RCM90, SV98, UC92]. larval [BS92, BB92, CCC^{+}98, Fos92, GG98, HS91, PCH91, PES99, SHM95, SH98]. larve [BBG91]. larves [BB94]. late [RM96a]. Lateral [HD96b]. latitudes [PTO95]. Laurec [PK95a]. law [HJB90, MI96]. Lawrence [CR97, Cas99, SC99, SS95a, SNSC94]. layer [IMH96, LLLL99, RTK95]. layers [Ped96b]. least [Far92, Hun97, Jen90]. least-squares [Far92, Jen90]. Lecture [Ano93c]. Lee [Dic94]. legged [Mun97]. Length [Jon90, KN99, DKLM99, Gud95, MC90, MS90, MB97, RCM90, RAFC92, Sam90, Ye98]. length-based [Ye98]. Length-cohort [Jon90]. length-structured [MS90, Sam90]. Lepechin [GA94]. Lepeophtheirus [Box97, CCC^{+}98, DPHM97, Daw98, JG97, MH98b, NLP97, SKHM98]. Lepidorhombus [SPL98]. Lernaeocera [VGHO97]. Lesions [JBC97, CZM98]. lesser [Gau90, OVP98]. lessons [DJ98]. Letter [Bra99]. Letters [Fra93]. leucas [HJR96]. levels [Cor99b, GS96, MH95, MUV90, PTCC98]. lice [BFNP97, Box97, DPHM97, Daw98, JG97, MLBM98, Mac98, NLP97, TIR97]. lidar [Bar98]. Life [OIF95, PN93, SSH95, BI92, CR99, Hog93, JB91, Mau95, Roc98, VVH91]. Life-cycle [SSH95]. life-history [VVH91]. light [WM99]. Limanda [PPS90b]. limitations [Coc99]. line [EGB^{+}96, FS93, GHM97, KZ97]. line-transect [FS93]. linear [FN93, MH97, PV93, SMGB97]. Linearization [Jen90]. lines [RCE^{+}93]. ling [BMM^{+}98]. lining [Løk98]. linked [NR96]. Linnaeus [HIL98, KWN95]. Linnhe [Hea95]. Lipid [HCR94, KH95]. List [Ano95c, Ano96c, Ano96d, Ano97h, Ano97g, Ano98c, Ano98d, Ano99b, Ano95a, Ano96a]. lithodid [Hog93]. littoral [FG96]. live [CSW92, MI96, SP92]. lived [MBR95]. liver [BLM99, CZM98]. lobster [AL91, DW99, FHA^{+}98, GGFFF98, HW90, Pol93, TCA97, VPT99]. local [Pay99, SJK90]. located [BCJ^{+}97, RWJF90]. location [Bra94, Gar97]. locations [MSG96]. Loch [Hea95]. LOCHNESS [DMUR93]. lochs [RGH93]. locked [KS90]. log [KZ97, MH97]. log-linear [MH97]. log-ratio [KZ97]. logistic [MK98]. loliginid [Hat96]. Loliginidae [CDL^{+}94, CBR95, HM99, JM99]. Loligo [CDL^{+}94, CBR95, Col99, Has96, HRP90, Hat96, HM99, LHSA98, LDU98, PBSN98]. Long [AG95, BS98, DSO96, Gay98, Gud98, HD96a, Jør90, Jør92, MW96, PLvdM^{+}96, Phi98, PT98, Tor90, VR98, WH96, WK94, AS95, AS97, Cus95a, DM96, EGB^{+}96, FAN99, FH96, GHM97, KZ97, Løk98, Mil96, Rad98, WA90]. long-finned [AS95, AS97]. long-line [EGB^{+}96, GHM97, KZ97]. long-lining [Løk98]. long-range [WA90]. Long-term [AG95, BS98, DSO96, Gay98, Gud98, HD96a, Jør90, Jør92, MW96, PLvdM^{+}96, Phi98, PT98, Tor90, VR98, WH96, WK94, Cus95a, DM96, FH96, Mil96, Rad98]. longimana [JCP91]. longitudinal [DK96]. loosely [BHS96]. loss [Løk98]. low [LR96, TL96]. low-frequency [LR96, TL96]. lower [BS98]. lozanoi [HCA93]. lucens [BW95]. lupus [FPHM90, PN93, PM96a]. lusci [VGHO97]. luscus [HH93].

m [Mau95]. mackerel [BH94, CG95, Daw91, FL98, HHA95, LCP96, Mac90, MB96b, NW93, RTWC97, WRT95]. macrobenthos [HFP96]. macrocephalus [CHØ92]. macrocosm [FØK94]. macrofauna [BH92, HBC^{+}92]. made [MSRK99]. Maine [MBSM90]. major [ODR92, SKD96]. making [LS98a]. Malformations [DvWC96]. Malin [GRDCW98]. Mallotus [Fos92, GG98, MP91, Nak90, SCM94]. mammals [PTCC98]. Man [EN97, OHN99, BEM96]. man-induced [BEM96]. Management [Fow99, JS99, PL92, SBSP99, AR90, Bas99, Bia99, BP99, CR99, CCW99, Chr97, Coc99, Coo99, DM99, GDJC99, HS99, Hor91, IOEJ97, JDLH97, KOS^{+}99, LS99, MH98a, MK99a, MSRK99, MB97, NLCS99, ODR92, Oel99, PK95a, Pay99, PKMP99, PH94, RP99, SOBP99, SP96, SSS99, SRAL99, Sut99, Vøl90, Wit99]. Manila [SG95]. Mapping [MAF95, Gre97a, Mis96]. March [Bla93, LP91b, NHP^{+}95]. March-May [NHP^{+}95]. Marine [Ano93b, FRK93, NLCS99, Sin91, TR97, AR90, AB96, Bax98, BMN97, BB94, CHR96, FN99, FC97, Hou94, MB96a, PTCC98, PLvdM^{+}96, SS93a, SN99, Sma98, SJB90, SC96, SWCG94, VVH91, dG96a]. marins [BB94]. marinus [Gau90]. mark [HN91a, HN91b, SHM95, SSFB97]. mark-recapture [HN91a, HN91b, SSFB97]. mark-release [SHM95]. market [KSP96]. Mass [MGN91, Har97]. masses [BFKR97, Mis93]. maturation [CBR95, Ims99, MBSM90, SC99]. maturing [CLAD97]. maturity [GGF94, HM99, Jør90, KCB99, MB97, MC99, SRJ98, TJ97]. maturity-at-age [MC99]. maturity-related [TJ97]. maximization [HGOT99]. maximus [BB92, FB99, Ims99]. May [Dic94, LP91b, AEH^{+}99, Dan97, NHP^{+}95]. meal [TA94]. mean [LL97]. means [Gre97a, Klo96]. measured [SW96]. measurement [BJM99, Daw91, Foo90b, PILB95]. measurements [BS98, FS93, GD96, Ham96a, KWK97, LCP96, NL99, ROH^{+}98, TL96, Tra96, UC92, dBH95]. measures [MS90]. Measuring [TZD^{+}95, CLB92]. mechanism [Cor99a]. Mecklenburg [ML99]. Mediterranean [BCCP96, DLMS97, KLF96, MSC98, OR97, PS97, SV98, SVS99, SPCP97]. mediterraneus [VGF97]. Medusae [Mil95, SB94]. megafaunal [Kai98]. megrim [SPL98]. meiobenthos [HHHS92]. Melanogrammus [Alb94b, BWH97, JJ96a, PV93]. melas [AS95, AS97]. membras [KRKS97, Raj93]. memory [Har95]. Mercenaria [RW91]. meristic [JPG93]. Merlangius [BWH97, HH93, HRBA91, MT97b, Rob92]. merlangus [BWH97, HH93, HRBA91, MT97b, Rob92]. Merluccius [Gui95, Ham96a, Tra96]. mesh [GCM99, KSP96, PS97, RAFC92, SPCP97]. meshes [SPCLT96]. meso [Kim97]. meso-scale [Kim97]. mesocosm [dVDP^{+}98]. mesocosms [HGDC94, SVJ99]. mesopelagic [ML96]. mesoplankton [RTK95]. mesoscale [RDHP97]. Mesozooplankton [Smi95, Ber95, Ric95]. Metal [Ped96a, HFP96, Lor91]. metals [Bal93, SKL98]. métamorphose [AL95]. metamorphosis [AL95, MFGP96]. Metazoan [Køi99]. method [ARW93, BG94, BC96, BR93, Daw91, hLHM^{+}92, MB96b, RJ90, She99, SS99, SK95b, TF96, ZF98]. Methods [WK99, Foo91, Gle96, HIL98, LPL95, Løk98, MSRK99, PK95a, RKCM92, SS93a, She91, SHB96, UC92]. Metridia [BW95, SSH95]. Mexico [RDHP97, RG93]. Micro [RBSV98, Hic98]. micro- [Hic98]. Micro-scale [RBSV98]. Micromesistius [MMHC98, OVP98]. micronektonic [AH92].

Micropogonias [APD92]. microsatellite [Nil97]. microscope [ARW93]. microscopic [TZD^{+}95]. microstructure [CM91, MFGP96, MF91, Mok92a, Mok92b, SS93b, SVJ99, ZM93]. microzooplankton [SVS99]. midnight [SS96]. midwater [DMUR93, HD90, Nak90]. migrating [CG95, LGKM97]. Migration [KNI92, PU93, SCM94, God95, HM95, RTWC97, RdC95, WRT95, SK90a]. migrations [TSAS96]. migratory [BH98]. Mini [Ano93b, BFL97, GR98a, HA92b, HA92a]. Mini-Symposium [Ano93b, BFL97, GR98a, HA92b, HA92a]. Minimizing [Rym97]. minimum [MH95]. minke [HGLN95, MRL92, MEH97, MP99, NTWW95b]. mise [BB94]. misreported [Pat98]. mitchilli [HGDC94]. Mitochondrial [RG93]. mixed [HCR94]. mixing [BR97]. mixture [Mal90]. Mnemiopsis [KN94, Kre94, Mut94, VVM95]. MOCNESS [PLPN90]. Model [Har97, AH95, BE90, CR99, DJ92, FB99, Fry91, Gre95, Hor92a, Jen91, Lar92, Mal90, MRL92, MK98, MAYC96, MS90, MK96, MC98, Pat98, Pet97, PV93, Sam90, Spa90b, Spa95, SC96, TA94, VPT99, Ye98, dVDP^{+}98, dSJ95]. modeling [KZ97]. modelled [MSRK99]. Modelling [Ano93b, CB93, FA97, GR98b, Mil92, MRN94, RP95, TA94, VVM95, Wan97, WB95, BBBR98, BW95, MMHC98, PHT95, SN91]. models [Fra97, Han96, HPDR93, Jen95, MH97, PH94, RG96, RGH93, SS95a, vLRvR^{+}96]. mollusc [WK94]. Mollusca [WK94, Wit96]. Molva [BMM^{+}98]. Monhystera [VVH91]. Monitoring [GW93, Rey98, SJ96, Bar98, Bel98, FBG96, FN93, MWRH98]. monitors [FC97]. monkfish [DAP97]. monogeneans [Køi91]. monospecific [SDW^{+}96]. Monthly [WA90]. Moray [GMA98]. morhua [AW92, AWEH94, BH98, Bra95, BWH97, CP96, Cas99, DJ92, FG96, FØK94, Han96, HP99, HGD92, Jør90, Jør92, KWN95, ML99, MON98, MT96, NL99, NP98, PU93, PLPN90, RdC95, RP96, SRJ98, SC99, SSFB97, SP96, SMGB97, SS93b, SS96, SNSC94, TA94, WH97, dBTT98]. morio [RG93]. Morone [SHM95]. morphometric [GGF94, JPG93]. mortality [Chr95, Far92, HN90, HGDC94, Jen90, JS90, KHK94, LP91a, MSPG92, MRSF91, MGN91, OW95, RJ90, Spa90b, Spa90c, SPCLT96, dBTT98, dBT98, SK90b]. moult [GGFFF98, NHP^{+}95]. mounted [GD96, Rey98]. mouth [HHB96]. movement [AH95, SJ96, SS99, WHM97]. Movements [AWEH94, CLAD97, Gau90, MA99, AW92, KWN95]. moving [Mis93]. MSFOR [Spa95]. MSVPA [Spa95]. MSVPA/MSFOR [Spa95]. MtDNA [Nil97]. much [MK99a]. mud [HW90]. Muggiaea [Gre94]. Müller [GG98]. Multi [HJ99, MH95, DMUR93, Fow99, HM95, SFG96]. Multi-annual [MH95]. multi-beam [HM95, SFG96]. Multi-frequency [HJ99]. multi-net [DMUR93]. multi-species [Fow99]. Multiple [AB96, Jen91, Col99, Hat96, MS90, She97]. multiple-fleet [MS90]. multiple-species [MS90]. Multiplicative [SN91, Lar92]. Multiscale [PW95]. Multispecies [BE90, Spa90b, EGB^{+}96, Gis99, MKP96, SK95a]. Multivariate [NFM98, JPG93, MK99a, OB97]. murphyi [HHA95]. mussels [KL97]. Mutual [EHJN91, Gre95]. mykiss [Nie99]. Mysticeti [CHØ92]. Mytilus [KL97].

N [Ric95]. Namibian [Oel99]. nanoplankton [Hic98]. Narragansett [BS98, LS98b, Sma98]. Narrowband [SDW^{+}96]. national [DM99]. native [McG97]. natural [BMN97, DJ92, HCH91, JS90, LP91a, McG97, ROH^{+}98, Spa90b, Spa90c, dBT98]. nauplii [PCH91]. near [Mis96, OM96, Swa99, TT96]. near-surface [TT96]. nearly [KS90]. Necora [GGF94]. needs [BBBR98]. neglect [HW97]. nematode [VVH91]. nematodes [Køi91]. Nephrops [EHEW94, FHA^{+}98, GGFFF98, HW90, MSC98, TCA97, VPT99]. net [CFUCY95, CC93, DMUR93, Jen99, LHSA98, RSN96]. net-zooplankton [CFUCY95]. Netherlands [HH93]. network [SAC96]. networks [HG96]. neural [HG96, SAC96]. Newfoundland [CP96, Dem99, KWN95, MP91, ODR92, RdC95, RP96, Ros98, SCM94]. Newfoundland-Labrador [KWN95]. Nilsson [Alb94a, BWH97]. nimbaria [ML96]. nine [Spa90b]. nitrogen [AS95, AS97]. no [Jen91]. nodules [BLM99]. noise [WB96]. non [CLAD97, HD96a, SMGB97]. non-linear [SMGB97]. non-maturing [CLAD97]. non-target [HD96a]. nonlinear [Jen90]. Nordic [MRF97]. Normalization [Lor91]. North [BED^{+}93, EHEW94, FH96, HCR94, HA92b, HA92a, LV94, MP99, AS97, AGH^{+}93, AH92, AG98, BME96, Ber93, BRK97b, BTC93, BR93, CZM98, HTG97, HGLN95, JG93, MEH97, MM97, MB97, OR97, OMN98a, OMN98b, RTWC97, RdC95, SRJ98, SVS99, WH98, WWS93, dPVV^{+}96, AS95, AWEH94, BHKH94, BME96, BW95, BP96, BH92, Ber93, BWH97, CG97, CR97, CCS97, CKF91, CR93, Coo97, Coo98, CNF90, CvdK96, DM96, DBH98, DvWC96, Eco91, Ehr91, FRK93, Gar97, Gau90, GBB^{+}97, HFP96, Han96, Hay95, HHB96, HD96a, HBC^{+}92, HRBA91, His96, Hor91, Hul95, HHHS92, Jak92, JS90, Jør92, JG93, KOS^{+}99, Kin92, KBC^{+}92, MA99, MON98, MW96, MA92a, Mok92a, MBT95, Neh98, NFM98, OBD^{+}97, Par99, Ped96a, Pep90, Phi98, PR98, PT98, PM96b, RPW^{+}99, RG96, RP98]. North [Rij91, RDvBH91, RP95, RvLDH96, RvL96, RM96a, RBSV98, SV98, SKL98, SKD96, SF96, Spa90a, Spa90b, SV91, Tay95, Tur92, TSP^{+}96, WH96, WHM97, WH98, WK94, dG96a, dG96b, vLRvR^{+}96]. north-east [AH92, BTC93, CZM98, HTG97, HGLN95, JG93, MEH97, OMN98a, OMN98b, RdC95, Jak92, Jør92, MON98]. north-eastern [Ber93]. north-west [BRK97b, BR93, MM97, MB97, SRJ98, WWS93, vLRvR^{+}96]. north-western [BME96, OR97, SVS99, WH98, SV98]. Northeast [SSJK98, RRL^{+}92, Jør90, Mac90, Ned90, Ynd99a]. northeastern [Køi99, MRTB95]. northern [Arr96, BHKH94, BE90, BWH97, CG97, CBG96, Cas99, Eco91, FPHM90, Gui95, HTG97, HN90, Hor93, KSP96, KWN95, MM97, MUCM98, NP98, RDHP97, SPL98, SCM94, Tur92, BCCP96, Dan97]. Northumberland [FH96, FBG96]. Northwest [MBSM90, Nak90]. northwestern [Gau90, MSC98, DW99]. norvegicus [FHA^{+}98, GGFFF98, HW90, MSC98, TCA97, VPT99]. Norway [Alb94a, AGH^{+}93, BMM94, BE90, BWH97, DJ98, EHJN91, FPHM90, FHA^{+}98, God95, GGFFF98, HW90, HN90, KS90, MH98b, NN91a, NP98, SHKL97, SKHM98, SJK90, SK90a, SK90b, Tor91, TCA97, VPT99, Vøl90]. Norwegian [Alb94a, Alb94b, CHØ92, Foo99, HKN^{+}91, HS98, HGD92, HO96, Jen99, KN99, Mac99, Mar93, MF92, Mok92b, Nøt98, PMF^{+}96, SS93b, SS96, SVJ99, Tor90, Tor91, UJA^{+}93, dBH95, dBTT98, dBT98]. note [O’B99]. nourriture [BBG91]. Nova [SRAL99, SS96]. NOWESP [vLRvR^{+}96]. nucleus [ZM93]. numbers [ARW93, LL97, NP98, PM96a, RG96, WA90]. numerical [CHR96]. nursery [BDC^{+}94, HS98]. nutation [Ynd99b, Ynd99a]. Nutrient [WMO98, ADS98, Hic98, RGH93]. nutrients [DLMS97, dVDP^{+}98]. nutrition [KB95]. nutritional [BB94, BRK97a, KB95, UC92]. nutritionnel [BB94]. NW [FH96, MFL98, GCG96, Mun97].

obesus [BJM99]. Objective [HM99]. Observations [FPHM90, BJM99, Mac99, MJ96, NW93, RDHP97]. observed [JBC97, Jen99]. obtained [Mis96, Ped96b, SBH95]. Occurrence [KN94, MH98b, WB97a, CAWD97, FPHM90, SKHM98, TIR97, YWMW97]. occurring [SV98]. Ocean [Har97, MRF97, Ban95, JG97, Mac99, PW95, RTK95, Smi95, FG95b, HTG97, HN95]. oceanic [BFKR97, Ber95, Cas99, JCP91]. Oceanographic [BFL97, Dan97, GRDCW98, MRTB95, MM97, SVS99]. oceanography [HP95, Man93]. October [HA92b, HA92a]. OECD [Sut99]. off [AGKB91, ADS98, Bod96, CP96, CBG96, FBG96, God95, HHA95, HJR96, JS99, MJ97a, MP91, MLCB99, NP98, OF95, PB97, PH94, RRL^{+}92, SKS97, SKHM98, TIR97, UJA^{+}93, VM98, WRT95, WB95, dPVV^{+}96]. offs [MSRK99]. offshore [Kin92, dG96b]. oil [Kin92, dG96b]. Oithona [FG95b]. oligotrophic [WB95]. oligotrophy [WB95]. Ommastrephidae [LN93]. Omnivory [Atk95]. Oncorhynchus [BTS^{+}97, LML99, Nie99, RSN96]. one [WT96]. one-dimensional [WT96]. Onychoteuthidae [JCP91]. Oosterschelde [VGHO97]. Open [JG97, AW92, Ano93c]. Open-ocean [JG97]. Opening [Mau96]. opercularis [BB92]. Optimal [Vøl90]. Optimization [Wit99, FG95a]. optimum [Bel98]. options [MH98a, MK99a]. orange [KWK97]. ordinary [Jen90]. organ [SC99]. organisms [RBSV98]. organization [PL96]. Organochlorine [SK97]. orientation [MAYC96]. origin [ADS98, HRL97, NLP97, SL97, dBH95]. original [SHKL97]. oscillations [LBDH98, RTK95]. Oslo [MH98b]. Ostrea [RCE^{+}93]. other [FG96, TSAS96]. Otolith [Daw91, ARW93, CM91, CB93, MFGP96, MF91, Mok92a, Mok92b, SS93b, SVJ99, Wal94]. otoliths [ZM93]. otter [SNSC94]. otter-trawl [SNSC94]. output [SL97, WWS93]. over-dispersion [KHK94]. overexploitation [PL92]. overfishing [Fow99, MRM^{+}94, Pol93]. overlapping [Rym97, SBSH97]. overview [Kre94, NLCS99, SKD96]. overwintering [HO96, PTO95]. oxygen [ADS96, DRR94, Ims99, PPS90b]. oyster [GH91, RCE^{+}93].

P450 [BSMK99]. Pacific [Nie99, BMN97, CH98, Ham96a, LML99, MP99, McD98, MT97a, MBR95, RSN96, RR99, RA97, SP92, Swa97, Tra96]. pacifica [MAI96]. packed [AB96]. packing [Mis93]. pagurus [AL91]. paired [KZ97]. Pallas [PPS90b]. Pandalus [HN90]. Paper [HA92a]. Papers [HA92b, Ano94c, Hou90]. Paralomis [Hog93]. parameters [Jon90, KHØ95, PK95a, VVH91]. parasite [McV97]. Parasites [LM99, BRK97b, Køi99, LMW^{+}99, Mac90, ML99]. parasitic [JDLH97, Køi91]. parental [Gré97b]. Parrish [Bla93]. Part [KS90, SK90a, Gui95, SJK90, SK90b]. partial [War94]. particle [MT97b]. particles [MT97b]. particular [LHHN98]. partly [MT97b]. partnership [SSS99]. passage [FVC94, RSN96]. Past [GB97, IOEJ97, MH96]. Patagonian [HRP90, HM99]. patches [HW90]. Patchiness [DK96]. Patchy [KO97]. pathological [Gry99]. pathways [KH95]. pattern [KNI92, MT97b, NW93, PU93]. Patterns [Kre94, PES99, RG96, Sma98, dVDP^{+}98, AH92, BDC^{+}94, BEM96, BWH97, FØK94, LSP^{+}95, Mau95, Md95, SJ96, SS95b, SG98]. Pauly [Ims99]. Pb [SKL98]. PCAIV [PL97]. Pecten [BB92, FB99, HSFAM99]. peculiarities [PN93]. peel [RW91]. peels [RCE^{+}93]. pelagic [BP90, BP91, BHS96, BWH97, CBG96, DvWC96, FPHM90, GMA98, Gre95, MKP96, MA92a, MLCB99, PL96, PLPN90, SKD96, SS93b, SS96]. pelagicus [vdML95]. penaeid [Ye98]. Pennant [BMM^{+}98]. Performance [SBSH97, McG97, PK95a, RP99, Wan97]. period [BMS98, DvWC96, Fos96, GGFFF98, HW97, HP99, JJ96a, Mok92a, PTO95, Rey98]. periodicity [CHH91]. periods [Cor99a, FGS96, Rij91]. perspective [LML99]. perspectives [NLCS99]. Peter [Mad94]. petrel [vdML95]. petrochemical [BGA97]. petroleum [SW96]. Phalacrocorax [CC93, Gré97b, Wan97]. phase [MA99, SBSH97]. phenotypic [RVD90, WW96]. philippinarum [SG95]. Phoca [HKN^{+}91, KHØ95, LHHN98, MS96, NHP^{+}95]. Phocoena [GS96, NTWW95b, TBHL97]. phosphate [Bod96, Laa96]. photography [EMN^{+}94]. Photololigo [JM99]. Physeter [CHØ92]. Physical [HN95, MW97, Man93, CH98, GRDCW98, RA97, dG96a]. physico [Col98, LBDH98]. physico-chemical [Col98, LBDH98]. Physics [Hun97, Smi95]. physiological [Daw98, Eve93]. Phytoplankton [GHH98, Bar98, Bel98, DJ98, Gay98, Gud98, LPL95, Neh98, NFM98, ROH^{+}98, RP98, SSD96, Sma98, dVDP^{+}98]. pileus [Mut94]. pilot [AS95, AS97]. pinnatifida [FPW91]. PIROP [HL97]. Pisces [Ber93, HCA93, HHB96, SV98]. piscivorous [GMA98]. piston [Foo95]. Pivotal [Ban95]. Plaice [CNF90, EN97, Hor93, KOS^{+}99, Mil96, MFGP96, MC99, PR98, RJ90, Rij91, RDvBH91, RP95, RvL96, RM96a, WHW99, vdWRvdLF90]. planktivorous [Mil95]. Plankton [FL98, GRDCW98, MWRH98, Sim96, Col98, LSP^{+}95, MH96, MSG96, MUCM98, SH95]. planktonic [AH92, Hea95, Gre95]. Plasma [BTC93]. platessa [CNF90, EN97, Hor93, KOS^{+}99, MFGP96, Rij91, RP95, RM96a, vdWRvdLF90]. platessoides [FAN99, WHW99]. platform [SW96]. Platichthys [BSMK99, BLM99, DKLM99, Køi99, LMW^{+}99, Möl90]. Platinum [BRS96]. Pleurobrachia [Mut94]. Pleuronectes [CZM98, CNF90, EN97, Hor93, KOS^{+}99, MFGP96, Rij91, RP95, RM96a, vdWRvdLF90]. PLITCH [Gre95]. poaching [HS99]. point [BB94, Fos92]. points [Gis99, Jak92, KHK94]. poissons [BB94]. Polar [KH95, GA94]. policies [CR99]. Pollachius [BWH97]. pollock [BS92, PPS90a, SSKP94, Swa99, Tra96, WT96]. pollutants [VVH91]. Pollution [Ano93b]. Pollution-Associated [Ano93b]. Polybius [Mun97]. polychaetes [BFNP97]. polycyclic [SK97]. polymorphisms [BTC93]. polynyas [BFKR97]. Pomatoschistus [HCA93]. Population [HN90, JPG93, OHN99, SB94, Ser90, TCA97, BW95, Ber93, Bod96, BRK97b, CR97, Col99, DB92, FG95a, GA98, HCH91, HN95, Lam90, LP91a, MRL92, MW96, Moh99, MRN94, NP98, PK95a, PBSN98, RW91, RM96a, RR98, Sam90, Smi95, THJ^{+}95, VPT99, VKA98, VVM95, vdML95]. populations [BEM96, BTC93, BCJ^{+}97, DMAG97, Dem99, Gle96, HW90, Jen99, MSC98, McG97, MK99b, McV97, MS96, Mil96, MM97, MC99, Mut94, Pol93, RR99, RM96b, Rym97, SC96, Ver97, WH96]. Porcupine [WMO98]. porpoise [GS96, TBHL97]. porpoises [NTWW95b]. port [CC99, OHN99]. port-based [CC99]. Portuguese [EGB^{+}96, GCM99]. Portunidae [GGF94]. possible [FPW91, GRN96]. post [Daw98, SWB96, WB96]. post-processing [SWB96, WB96]. post-smolt [Daw98]. pot [AL91]. Potential [Chr95, SHB96, DBH98, FG96, MRN94, NLP97, SK95b, War94]. pound [CC93]. pout [Alb94a, BWH97]. poutassou [OVP98]. power [FN93]. prawn [AGKB91, HN90]. pre [SCM94]. pre-spawning [SCM94]. Precautionary [RP99, DM99, KOS^{+}99]. precise [TF96]. precision [CM91, GKT98, JM99, Klo96]. Predation [Han96, SH95, Tor91, Bax98, CHR96, Gre95, Jen95, PM96b, dBTT98]. predation-rate [Jen95]. predator [Nøt98, SC96]. Predators [Ano94c, Mil95]. Predict [Ano93b, HPDR93]. Predicted [OBD^{+}97]. predicting [AH95]. Prediction [Cus95a, She97, Bax98, Cla94, Ull96]. predictions [Han96]. Prefecture [SKS97]. preferences [LHHN98]. preliminary [God95, GK97, Mut96]. preparing [ARW93]. present [CHØ92, GB97, MH96, MW96, WH98]. presentation [Bel98]. prevalence [FHA^{+}98]. prey [AS97, BK96, DJ92, GMA98, LHHN98, MFL98, PPS90a, PCH91, SC96, SVJ99]. Pribilof [Swa99]. Primary [LBDH98, AS97, BC96, MFL98, NR96]. Principal [PL97]. priorities [MB96a]. priors [MK99a]. privation [BBG91]. Probabilistic [SS93a]. problem [FS93, Moh99]. Problems [KWK97, IOEJ97, McD99]. procedure [hLHM^{+}92, MUV90, WT96]. procedures [Bas99, BP99, GDJC99, KOS^{+}99, MC98, Wit99]. process [BFKR97, Som99]. Processes [BFL97, MW97, FA97, Fra97, LBDH98, MSRK99, RTWC97, SVS99, vLRvR^{+}96]. processing [OM96, SWB96, WB96]. produced [BFKR97]. production [Ban95, BMN97, Ber95, BC96, FG95b, HH93, Ham96b, HS91, Hay95, Hor92a, HVMI^{+}95, Kin92, KB95, KL97, LBDH98, ML95, MBR95, MFL98, NR96, PILB95, PHT95, RP98, SS95a, WB95, ZF98]. productivity [BMN97, Gud98, Kim97, RR99, Sak97]. productus [Ham96a, Tra96]. Professor [Ano96e, WO93]. Profiler [GD96, RGHC96]. profiling [McG97]. progeny [McG97]. Programme [vLRvR^{+}96, FN93]. programs [SSFB97, SK95a]. propagation [RR99]. propinquus [SSH95]. Proportion [VHV^{+}98]. proposed [Cor99a]. protected [PR98]. protection [MB96a]. protective [AR90]. Protein [AGH^{+}93]. proximity [Swa99]. Pseudodactylogyrus [Køi91]. puber [GGF94]. purposes [CB93]. purse [HHA95]. purse-seiners [HHA95].

quality [CKF91, Eve93, KB95]. Quantification [Rad98]. Quantitative [Cus95b, dG96b, MSRK99]. quasi [Ham96a]. quasi-ideal [Ham96a]. Québec [CR97]. quicker [MK99a]. quota [SK95a].

radiata [Ped95, SV91]. radionuclides [FS97]. Raitt [Gau90]. Raja [Ped95, SV91]. ranching [IOEJ97]. random [Mur96, SF96]. randomly [Fog93]. range [RWJF90, WA90]. ranges [HTG97]. ranking [MK99a]. RAPD [MEH97, MP99]. RAPD-typing [MP99]. Rapido [HSFAM99]. Rapport [BB94]. rate [HCH91, Jen95, LP91a, MF92, Mok92a, PPS90a, RCE^{+}93, dBT98]. Rates [PPS90b, AH95, Bes93, ES92, Far92, GHM97, Han96, HLAM95, KHK94, KN99, MK95, Sam91, Spa90b]. rather [BRK97a, HJ99]. ratio [KZ97, UC92]. ration [HCA93, PPS92, TA94]. rational [BDS96]. ratios [AS95, AS97, MT96, PES99]. Ray [Ano96e, Ped95, SV91, WH96, WHM97, WH98]. rays [PM96a, WH98]. razor [GCM99, RR98]. Re [JCP91, MAC96]. re-evaluation [MAC96]. Re-examination [JCP91]. readings [BMM^{+}98, LDU98]. reared [BTS^{+}97, BCJ^{+}97, DMAG97, HRL97, HIL98, KS90, SVJ99, SJK90, SK90a, SK90b, THJ^{+}95, YWMW97]. rearing [Dem99]. recapture [HN91a, HN91b, SSFB97]. recognizing [SHB96]. reconstruction [KO97]. record [BG93]. Recorder [MWRH98, LSP^{+}95]. Recording [HM95, LSP^{+}95]. recoveries [KHK94]. Recovery [WTB99, Fos92]. recruit [BS92, PLvdM^{+}96, She97, Ye98]. Recruitment [CBR95, Fog93, Pol93, Hat96, Hou94, Lam90, LP91a, MK99b, MRM^{+}94, MBT95, Pep90, PM96b, PILB95, RKCM92, SMGB97, Spa96, Tur92, SJK90]. Recurrent [Md95]. red [RG93]. redfish [MJ97a, MBSM90, NN91a, NN91b, NJN94, PR93]. reduce [CC93]. Reduced [DRR94, Chr95, Ull96]. reducing [PL92]. reduction [hLHM^{+}92]. reef [Gle96]. reef-fish [Gle96]. Reeve [SG95]. referee [Ano97g]. Referees [Ano90b, Ano91c, Ano91d, Ano92b, Ano93d, Ano94b, Ano95d, Ano98c, Ano95c, Ano96c, Ano96d, Ano97h, Ano98d, Ano99b]. reference [FGS96, Gis99, Jak92]. reflected [RG96]. reflection [Wit96]. regime [Ynd99b]. region [MBSM90, VR98, ZM93]. regional [Hay95, RM96b, Smi95]. regions [JG93]. regression [PV93, She97]. regulating [NBBS95]. regulatory [MS90]. Reinhardtius [BRK97b, MB97, OB97, PR93, RSB99]. relaid [KL97]. related [Ims99, SCM94, SPCLT96, TJ97, Wit96]. Relating [SSKP94]. relation [AWEH94, ADS98, BME96, BMN97, Bod96, Cas99, CvdK96, GRDCW98, HP99, Hic98, MP91, MON98, MFL98, PLvdM^{+}96, RRL^{+}92, RTWC97, RDHP97, RvL96, RBSV98, Rob92, SVS99, Ser90, Tur92, TSP^{+}96, VM98]. Relationship [CR97, IMH96, SKS97, Agl96, Cus95a, DKLM99, Jen95, PES99, SN99]. Relationships [CBG96, DB92, MABD92, MJ97a, MAC96, RA97]. relative [Lar92, MFPT98, NTWW95b, PPS90b, RdC95]. release [Far92, Nøt98, SHM95, SK90b]. releases [SE98]. Relevance [dBTT98]. reliability [MGF96]. remote [Bar98]. remove [WB96]. rents [SK95a]. replace [SHKL97]. replacement [War94]. replicas [RW91]. Report [BB94]. representatives [Pay99]. reproduction [Ber93, THJ^{+}95]. Reproductive [BTS^{+}97, Hou90, RDvBH91, VGF97, BP90, BP91, GGF94, KHØ95, LGKM97, MRN94, SC99, Ser90, WWS93, WW96]. requirements [PPS92]. research [Coo97, Ehr91, Has96, PBSN98, RKCM92, SNSC94]. reservoir [SJ96]. resident [Mut94]. resilience [Pol93]. resilient [WH98]. Resolution [Col99, MAF95, MFPT98, PMF^{+}96]. resolving [Hat96]. Resource [SK95a, PH94, Rit97]. Resources [Sin91, AR90]. respect [CCW99]. response [GR98b, GMA98, LR96, Smi95, SG98]. responses [BK96]. restoration [SH98]. results [AEH^{+}99, BC96]. retrospective [Moh99]. returning [Daw98]. returns [BV96]. revealed [BS98, Coo97, DK96, JPG93, PMF^{+}96]. Review [Bev91b, Bla98, Cus90, Daw94, Lip99, Loc91, Mil99, Hun97, Jen95, Jon97, Man93, MK98, Pet93]. Reviews [CWGC93, Hil90, Hor90, Hor92b, Mar91, MGF93, Pur91a, Pur91b]. revisited [OM96]. Rexea [PS99]. reynaudii [LHSA98, LDU98]. richness [VM98]. Ricker [Spa95]. right [Spa95]. right-hand [Spa95]. rigid [SE97, SPCLT96]. risk [DBH98, Hay98, HPDR93, LS98a, SS93a, WK99, Wit99]. Risks [Ano93b]. Rissa [HW97]. River [Dem99, RSN96, SL97, Vøl90, CLAD97, CAWD97, DK96, LGKM97, McG97, SE98, SHKL97]. riverine [MK96]. RJHB [Ram91]. RNA [BB94, PES99, UC92]. RNA/DNA [PES99, UC92]. Rockall [KCB99]. rockfish [KCB99]. Role [SS96, Ban95, KB95, SB94, WHW99]. Rossfjord [Hog94]. rough [FAN99]. roughy [KWK97]. roundfish [SKD96]. RoxAnn [Gre97a]. rules [RP99]. Russian [Jen99].

S. [DB92, KLF96]. sablefish [KZ97]. sac [HGDC94]. Sætersdal [Øst99]. Sagitta [KLF96]. saida [GA94]. saithe [BWH97, Mis93]. salar [BDC^{+}94, Box97, CLAD97, CAWD97, DMAG97, DPHM97, HRL97, HJL99, HIL98, JG97, McG97, NLP97, ODR92, OMF^{+}97, Rit97, Ver97]. salinity [CBG96, Laa96, VHV^{+}98]. Salmo [BDC^{+}94, Box97, CLAD97, CAWD97, DMAG97, DPHM97, Daw98, HRL97, HJL99, HIL98, JG97, MLBM98, McG97, MH98b, NLP97, ODR92, OMF^{+}97, Rit97, SKHM98, TIR97, Ver97]. salmon [BDC^{+}94, BMN97, BTS^{+}97, BCJ^{+}97, Box97, CLAD97, CAWD97, CR99, DMAG97, DPHM97, Daw98, Dem99, EHJN91, FE97, FRK93, GB97, HRL97, HJL99, HIL98, IOEJ97, JG97, Jen99, Jon97, KNI92, KP97, LGKM97, LML99, McG97, MT97a, MK99b, McV97, MI96, Nil97, NLP97, ODR92, OMF^{+}97, RSN96, RVD90, RR99, Rit97, SHKL97, SE98, SL97, TT96, Ver97, WK99, YWMW97]. salmonicida [WTB99]. salmonids [BFNP97, JDLH97]. salmonis [Box97, CCC^{+}98, DPHM97, Daw98, JG97, MH98b, NLP97, SKHM98]. salps [MK95]. sample [FL98, Mur96]. sampled [OB97]. sampler [DMUR93]. samplers [PLPN90]. samples [HS91]. Sampling [Mil94, GKT98, Jaf96, MA92a, MLCB99, OM96, ROH^{+}98, War94]. sandeel [WHG98]. sandeels [Gau90, WB97b]. sandfish [SKS97]. sandy [BH92]. sanguineus [McD98]. Santander [VM98]. sardine [BP90, PCM96, TSAS96]. Sardinella [BP90]. saxatilis [SHM95]. scad [BTC93, VGF97]. Scale [MW97, AH92, BMN97, BEM96, Kim97, LSP^{+}95, Mac99, RBSV98, dBH95]. scales [AH95]. scallop [Bea91, FB99]. scallops [HSFAM99]. scanning [ARW93, SJ96]. scattering [AY99, AB96, IMH96, LLLL99, SE96, SCW96, YF96]. scavenging [CG97, WB97a]. Schaefer [Jen91]. school [AY99, FGS96]. schooling [Mis96, MB96b]. schools [HM95, HG96, MKP96, MABD92, MA92a, Mis93, MAF95, MFPT98, PMF^{+}96, SFG96, SSKP94, Swa97]. science [LS99]. scientific [GD96]. scintillation [YCL96]. Scomber [CG95, Daw91, FL98, Mac90, NW93]. scombrus [CG95, Daw91, FL98, Mac90, NW93]. Scophthalmus [Ims99]. Scorpaenidae [HHB96]. Scotia [SOBP99, SRAL99]. Scotian [MFL98, SSFB97, SS96]. Scotland [Col99, FHA^{+}98, GMA98, MLBM98, RTWC97, TCA97, WRT95, WHG98]. Scottish [Has96, PBSN98, RAFC92, RGH93, SFCP99, WB97b, YWMW97]. Scyliorhinus [OVP98]. Scyphomedusae [OIF95]. scyphozoan [BMM94]. Sea [Atk95, BP96, EHEW94, FH96, Hay95, HA92b, HA92a, LM99, MLBM98, PHT95, Ric95, RTK95, SSH95, Swa99, AW92, BR97, BFNP97, CBG96, DPHM97, Daw98, Mac98, MH98b, NLP97, Ped96a, RBSV98, RGH93, SKHM98, SE96, TIR97, ZMHD96, AWEH94, Arr96, AH96, Arr98, BHKH94, BK96, BED^{+}93, BME96, BP96, BR97, BH92, Ber93, BV96, BSMK99, BLM99, BG94, BC96, BWH97, CG97, CCS97, CKF91, CR93, Coo97, Coo98, CNF90, CvdK96, DM96, DBH98, DvWC96, Eco91, Ehr91, EN97, FAN99, Fos92, FS97, Gar97, Gau90, GA94, GKT98, GBB^{+}97, Gry99, GG98, HCR94, HFP96, Han96, Har97, HHB96, HD96a, HBC^{+}92, HP99, HIL98, HRBA91, His96, HF94, Hor91, Hor93, Hul95, Hun97, HHHS92, JJ96a, JS90, JG93, KNI92, KOS^{+}99, Kin92, KHØ95, Køi99, KP97]. Sea [KN99, KBC^{+}92, Laa96, LMW^{+}99, LV94, LLLL99, LOÅ97, MJ97a, Mar93, MA99, MK99b, MW96, MA92a, Mok92a, MA92b, Mut94, Neh98, NFM98, Nie94, NHP^{+}95, OHN99, OLO98, OVP98, OBD^{+}97, PN93, Pep90, PLvdM^{+}96, Phi98, PR98, PMF^{+}96, PM96b, Raj93, ROH^{+}98, RPW^{+}99, RG96, RP98, Rij91, RDvBH91, RP95, RvLDH96, RvL96, RM96a, RBSV98, Sak97, SE98, SVS99, SKL98, SKD96, SF96, Smi95, Spa90a, Spa90b, Spa90c, SV91, SK97, Ste99, SSKP94, Thu97, TBHL97, Tur92, TSP^{+}96, VVM95, VHV^{+}98, WH96, WHM97, WH98, WB97a, WTB99, WK94, Ynd99b, dBT98, dG96a, dG96b]. Sea-going [LM99]. seabed [Gre97a, OM96]. Seabird [Løk98, BR97, Gar97, HL97, LS97, OBD^{+}97, RA97]. seabird-depth [HL97]. Seabirds [FC97, TR97, BK96, BGA97, JBC97, OR97, RDHP97, WB97a, WHG98]. Seal [EMN^{+}94, HKN^{+}91, MS96]. seals [KHØ95, LHHN98, NHP^{+}95, UJA^{+}93]. search [MRM^{+}94]. Searching [HHA95]. seas [Tay95, MJ97b, MRF97, RPW^{+}99]. SeaSoar [RGHC96]. season [NBBS95, SRAL99]. Seasonal [CHØ92, DLMS97, GBB^{+}97, GMA98, HGD92, MRSF91, MP91, MUCM98, Ric95, RTK95, SKHM98, SC99, AH96, Arr98, FHA^{+}98, Hay95, LGKM97, Smi95, VGHO97, WB95]. seasons [KLF96, Ye98]. seatrout [DB92]. Sebastes [Ned90, MBSM90, NN91a, NN91b, NJN94, PR93]. Secchi [BS98]. Secchi-disk [BS98]. secondary [Hay95, PILB95]. section [DSO96, Foo99]. sections [ARW93, RW91]. sector [DM96]. sediment [BH92, CZM98, DLMS97, SN99]. sedimentary [BED^{+}93]. sedimentation [FS97]. sedimenting [MGN91]. sediments [Lor91, RRL^{+}92, SK97]. see [MBT95]. segments [vdML95]. seiners [HHA95]. seines [RAFC92]. selected [Hou90]. selection [Hee96, KRKS97]. selective [EJW98, MRN94, RJ90]. Selectivity [SPCP97, AL91, Arr96, Atk95, EGB^{+}96, Fry91, Has96, Mac95, Mil94, MH97, PS97, RAFC92, TZD^{+}95]. semi [GMG98, LR99]. semi-automatic [GMG98]. semi-demersal [LR99]. Senegalese [PL96]. Sensitive [PM96a, WH98]. Sensitivity [FB99, BE90]. separable [KV97]. separation [Daw91]. September [Ano94c, Bev91a, Øst90]. sequential [Moh99]. series [CFUCY95, FN99, Gud95, Hee96, LBDH98, MSPG92, She97, SS99, VM98, dVDP^{+}98]. Seriola [RG93]. serratodentata [KLF96]. Session [Ano94c, MW97]. set [LBDH98, Mal90, TBHL97]. sets [Kap99, vLRvR^{+}96]. setting [Løk98]. several [Phi98, SCW96]. severely [Bes93]. sewage [RRL^{+}92]. sex [MT96, SP92]. Sexual [GGF94, Ims99, Jør90, SRJ98]. shags [Wan97]. shallow [BR97, HH93]. shape [CB93, Mis93, MAF95, Ste99]. shaped [Jen95]. sharks [Cor99b]. shear [Jen95]. shedding [HLB91]. Shelf [Ano94c, Har97, KWN95, MFL98, RdC95, SSJK98, vLRvR^{+}96, AH95, BFKR97, Bal93, JG93, RTWC97, SPL98, GHH98, GRDCW98, SSFB97]. shell [RW91]. shellfish [DLMS97]. Shepherd [PK95a]. Shetland [Mar93, MRSF91, TSP^{+}96]. shifts [BMN97, Tay95, WH98]. ship [GD96, vdML95]. ship-mounted [GD96]. shipping [DBH98]. shoals [AB96, MT96, SDW^{+}96]. shore [RWJF90, CR97]. shore-based [RWJF90]. Short [Roc98, Cus92, DM96, GCG96, MBR95, She91]. short-finned [GCG96]. short-lived [MBR95]. Short-term [Roc98, DM96, She91]. Should [LL97, MSRK99]. showing [BC96]. Shrimp [Chr97, Bod96, Chr95, OHN99, PK95b, SE97, WB97a, Ye98]. side [Spa95]. signal [OM96]. signals [CH98, Jaf96]. significance [Bax98]. silus [Ber93]. silver [Ber93, MA99]. silver-phase [MA99]. similis [FG95b]. Simple [She91, BE90, SC96]. simulated [Moh99]. Simulation [HPDR93, AH95, Coo99, Far92, Gre95, MRL92, Pat98, SF96, dVDP^{+}98]. since [PM96b, Rij91, RvL96, RM96a]. sinensis [CC93]. Single [Gis99, Far92, LR96, Rey96, SHB96, SBSH97, SK95a]. single-chambered [LR96]. single-fish [Rey96]. single-release [Far92]. siphonophores [Mil95]. site [MP91]. sites [HCR94]. situ [BH94, BHS96, BJM99, LSP^{+}95, MM96, NL99]. six [Coo97, LS97]. Size [AL91, CHR96, Hea95, PS97, RJ90, SS95a, SPCLT96, WH97, AGKB91, EGB^{+}96, FPHM90, GCM99, GR98b, HGOT99, KSP96, LN93, MAC96, Mil92, Mis93, MAF95, MSG96, MC99, MFL98, NBBS95, PPS90b, RAFC92, RG96, Spa96, SPCP97, Ste99, SVJ99, TA94, TL96, vdML95, vdWRvdLF90, AL95]. size-at-age [Mil92]. Size-based [SS95a]. Size-dependent [CHR96]. size-fractionated [MFL98]. Size-related [SPCLT96]. Size-selective [RJ90]. Size-specific [WH97]. sized [Mil94]. sizes [EMN^{+}94, HN91a, HN91b]. Skagerrak [ADS96, ADS98, Ber93, BG93, DJ98, DSO96, Dan97, HGD92]. skates [WH98]. skeletal [PM96a]. Skewed [MT96]. slopes [NN91a]. sludge [RRL^{+}92]. small [Alb95, LSP^{+}95, MLCB99]. small-scale [LSP^{+}95]. smelt [Ber93]. smolt [Daw98, KP97, RSN96, SL97]. smolts [Dem99, SE98]. smoothers [FN99]. smoothing [HGOT99]. snapper [ZF98]. soak [FA97]. software [HL97]. solandri [PS99]. Sole [Pet97, AL95, BBG91, BB94, BRK97a, CHH91, HN91a, HN91b, HCH91, Mil96, MW96, PPS90b, PPS92, RDvBH91]. Solea [AL95, BBG91, BB94, CHH91, HN91a, HN91b, HCH91]. Solomon [MBR95]. solution [BE90]. Some [Ano94c, Mal90, MRTB95, Pay99, FG96, Jør92, SVS99, Ser90]. Sonar [GS96, HM95, Jaf96, MAF95, Mis96, MFPT98, PMF^{+}96, RWJF90, SFG96, WA90]. sonic [BJM99]. sorting [SPCLT96]. Sound [NK99, SE96, AB96, CSW92, IMH96, LR96, LLLL99]. sound-scattering [IMH96, LLLL99]. sounder [GD96, Ped96b, SFG96, SBH95, SHB96, ZMHD96]. source [GS96]. South [AGKB91, CWC^{+}97, LHSA98, MRSF91, Tay95, BI92, DJ98, Gar97, HH93, HHA95, MH98b, Phi98, PR98, Rai93, WHG98, Ham96b, HS99, MBR95, Pay99, PB97, Wal94]. south-central [HHA95]. south-east [WHG98]. south-eastern [Gar97, MH98b, Phi98, PR98]. south-west [BI92, HH93, Rai93]. Southampton [RW91]. southerly [RTWC97]. southern [AWEH94, BK96, BH92, CCS97, CNF90, CC93, CvdK96, DAP97, GDJC99, GRDCW98, HLB91, HVMI^{+}95, LLLL99, MMHC98, NFM98, NHP^{+}95, PKMP99, PH94, RBSV98, Sam90, SKHM98, SC99, SNSC94, VM98, VR98, WK94, FG95b, HN95]. southwestern [BP91, Køi99]. sp [FHA^{+}98, Mut94, VVM95]. space [BHKH94, Pet97]. space-time [BHKH94]. spaced [MSG96]. spacing [GCM99]. Spain [GCG96, Mun97, VM98, dPVV^{+}96]. Spanish [PCM96, SPL98]. spat [BB92]. Spatial [BR97, HTG97, LMW^{+}99, MGMS94, MK99b, PL96, RRL^{+}92, SVS99, SS95b, Swa99, BMS98, CC99, DJ98, HP99, MKP96, MSC98, Orl99, Pet98, PV93, RP98, RP95, SFG96, WB97b, vdML95, Cus95a, WH98]. Spatial/Temporal [Cus95a]. spatio [PL96]. spatio-temporal [PL96]. spawned [MF91, RCM90]. spawners [BH98]. Spawning [EN97, KRKS97, Raj93, BMS98, Bra94, CAWD97, CHH91, CBR95, CNF90, Foo99, HBA98, HS98, HO96, JS99, LN93, LHSA98, Mac99, MF92, Mok92b, MT96, NW93, NK99, Nøt98, SCM94, Spa96, SJK90, Tor90, Tor91, WWS93, dBH95, dBTT98, dBT98]. special [FGS96]. Species [EGB^{+}96, SAC96, Agl96, Atk95, BHS96, BK96, CvdK96, EJW98, FG96, Fow99, HG96, HD96a, Hee96, Jen91, KN99, KBC^{+}92, LS97, Mau95, MAYC96, MS90, MBR95, Ned90, Neh98, NFM98, PS97, Phi98, RvLDH96, SV98, Spa90b, SPCP97, VM98, VGHO97, WHM97, WH98, ZMHD96]. species-dependent [KN99]. species-selective [EJW98]. specific [WH97]. specification [RP99]. spectra [GR98b, RG96]. spectrum [GS96, Hea95]. speed [HO96, MLCB99]. speeds [RdC95]. sperm [CHØ92]. Spheres [Foo90b]. spiny [Pol93]. split [BJM99, ET96, MMHC98, MK96, MC98, Ped96b, SBH95, SHB96, SBSH97, SJ96]. split-beam [BJM99, ET96, MK96, MC98, Ped96b, SBH95, SHB96, SBSH97, SJ96]. spotted [OVP98]. spp [Køi91, LML99, NN91a, NJN94, PR93, RSN96]. sprat [Arr96, Arr98, Gry99, HCR94, Mis93]. Sprattus [Arr96, Arr98, Gry99, HCR94]. spring [AG95, Atk95, Foo99, GA98, Gud98, HS98, HO96, Mac99, MF91, MF92, Mok92b, Nøt98, RCM90, RA97, Tor90, Tor91, WWS93, dBH95, dBTT98, dBT98]. spring- [MF91]. spring-spawned [RCM90]. spring-spawning [Foo99, HS98, Mac99, MF92, Mok92b, Nøt98, Tor90, Tor91, WWS93, dBH95, dBTT98, dBT98]. Spurious [LP91a]. square [PS97]. squares [Far92, Jen90]. squid [BR93, BI92, CDL^{+}94, GCG96, Has96, HRP90, HM99, JM99, LHSA98, LDU98, MRN94]. squids [GMG98, JCP91]. St [Cas99, CR97, SC99, SS95a, SNSC94]. stability [FBG96, HBA98]. stabilize [HJB90]. stable [AS95, AS97]. stage [FL98, GG98]. stages [BI92, CCC^{+}98, PM96a]. standard [DM99]. Standardized [Cor99b]. Standardizing [KZ97]. starry [Ped95, SV91]. start [AL95]. state [NBBS95]. States [CZM98, DM99, McD98]. statistical [MC98]. statolith [GCG96, JM99]. statoliths [GMG98, LDU98]. Status [SKD96, BB94, SG98]. stay [PTO95]. steelhead [Nie99]. Steenstrup [CDL^{+}94, CBR95]. Stochastic [VKA98, Spa90c]. Stock [Coo97, SMGB97, Ull96, BDS96, Bea91, BCCP96, CKF91, Cus92, Daw91, FL98, HN91a, HN91b, KP97, MK98, MGMS94, NW93, PK95a, Pet93, PM96b, PS99, RSB99, RVD90, SHKL97, SJB90, SP96, Som99, Spa96, SS96, SJK90, Tur92, WHM97]. stocking [SH98]. Stocks [Hou90, BK96, Bes93, CHØ92, CR93, Coo97, DAP97, Gis99, His96, Hul95, JB91, KNI92, Man93, Pat98, Pep90, SKD96, SK95b, SVJ99]. stomach [MT97b]. storage [SS99]. Storer [MBSM90]. storm [vdML95]. Strait [Chr95, OB97]. stranded [JBC97]. strategies [Chr97, DW99, PL92, PKMP99, Ric95, SSH95]. strategy [Dem99, Nie99, OIF95, PS99, SSS99]. stratified [HCR94]. Stream [Tay95, BDC^{+}94]. streams [AWEH94, CG95]. strength [BH94, BJM99, Foo91, Ham96a, HD96b, IMH96, KWK97, LCP96, MM96, MAYC96, MAC96, MMHC98, MB96b, MAI96, MI96, Mut96, NL99, RP96, Ros98, Ser90, She97, SBH95, Tra96]. strengths [BHS96]. stress [BGA97]. striped [SHM95, SH98]. structure [BCCP96, Ber93, DMAG97, FGS96, GA98, GK97, Hic98, HBA98, Lam90, MKP96, MSC98, Md95, MRN94, NP98, PL96, PM96b, RW91, SN99, WHM97]. structured [MS90, Sam90]. structures [AY99]. studied [FØK94]. studies [BFKR97, BP90, BP91, BRK97b, Far92, Gay98, God95, KS90, LSP^{+}95, NN91a, NJN94, Orl99, RSB99, RP96, SJB90, SJK90, SK90a, SK90b]. Study [DAP97, Far92, FB99, Fos96, GK97, HS99, MKP96, Mut96, RJ90, SG98]. Subarctic [CH98]. submarine [SVS99]. subtropical [Ye98]. success [HW97]. succession [BHKH94, Md95]. suction [RR98]. suggested [Hay98]. suitability [RSB99]. Summary [Foo91]. Summer [WHG98, Gar97, HGLN95, JS99, LS97, Swa97]. summer-spawning [JS99]. sun [SS96]. superba [MRSF91]. supplemented [RR99]. supply [Wit96]. support [Ims99]. suprapopulation [VGHO97]. surface [BP96, CBG96, Mis96, PTO95, TT96]. surrounding [Tay95]. Survey [Hee96, AEH^{+}99, Bal93, BED^{+}93, DK96, Gre97a, Gry99, Has96, KO97, Kap99, KZ97, LMW^{+}99, Mur96, NLCS99, PV93, Rey98, Som99, Ste96, SNSC94, WM99]. surveying [MA92a]. surveys [BMS98, Coo97, GW93, GWE99, HM95, Klo96, KN99, LR99, Nak90, PCM96, RG96, RGHC96, RKCM92, SF96, TF96, War94, WT96]. Survival [SE97, BDC^{+}94, FRK93, PES99]. survive [FVC94]. survivors [She99]. susceptibility [DPHM97]. sustainability [Coo98, Fow99]. sustainable [CCW99, Rit97]. SW [KRKS97, Gre95, SRAL99]. Swedish [PU93]. swept [Agl96]. swept-area [Agl96]. Swimbladder [Køi91, AW92, MAYC96, MMHC98]. swimbladders [LR96]. Swimming [MA92a, MAI96, WHW99, GGF94, HO96, MJ96, Mun97]. Symposium [Ano93b, BFL97, GR98a, HA92b, HA92a]. symptoms [Gry99]. synchronized [BEM96]. Synchrony [MBT95]. Synoptic [Gry99, LMW^{+}99]. synthesis [BSMK99]. system [Bar98, BGA97, Foo90b, GMG98, HVMI^{+}95, HL97, LSP^{+}95, MJ96, MK99a, MRF97, Mut96, NLCS99, SWB96, SSS98, Ynd99a]. systematic [SF96]. Systems [SBSP99, CCW99, HPDR93, LS99, MSRK99, SSS99, Wan97].

TACs [MH95]. tactics [JB91, Sam91]. tag [HLB91, KHK94]. tag-shedding [HLB91]. tagged [Gau90]. tagging [Far92, God95, MBHQ96, RP95]. tags [BRK97b, Mac90, SS99]. Taille [AL95]. Tapes [SG95]. Target [Ham96a, LCP96, MMHC98, MB96b, Mut96, RP96, Tra96, BH94, BHS96, BJM99, ET96, Foo91, HD96a, HD96b, KWK97, MM96, MAYC96, MAC96, MAI96, MI96, NL99, Ros98, SBH95]. Target-strength [Ham96a, LCP96, MB96b, RP96, Tra96, BJM99, MM96]. targeted [SFCP99]. targets [SHB96]. technique [Gre97a, WB96, YCL96]. techniques [SJ96, Wan97]. telemetry [Wan97]. Teleostei [VGF97]. teleosts [SJB90]. temperature [BP96, Bra95, CBG96, DB92, MP91, MON98, OMN98a, OMN98b, PPS90b, RvL96, SKS97, SCM94, SS99, SS93b, Ynd99b]. Temporal [Cus95a, DJ98, Hic98, Jaf96, LS98b, MB97, Box97, Col98, HBA98, MGMS94, MK99b, PL96, RRL^{+}92, RP98, SS95b, WH98]. ten [HD96a]. tendency [BH98]. teneur [BBG91]. term [AG95, BS98, Cus95a, DM96, DSO96, FH96, Gay98, Gud98, HD96a, Jør90, Jør92, Mil96, MW96, PLvdM^{+}96, Phi98, PT98, Rad98, Roc98, She91, Tor90, VR98, WH96, WK94]. Test [dSJ95, AR90, NP98, Sam90]. testing [Coo99]. tests [FE97, Gre95, Pat98]. Thaliacea [MK95]. Thames [CHH91]. Thau [DLMS97]. Their [MW97, AS97, BP96, Col98, ES92, JDLH97, MFL98, NK99, PES99, PM96b, Sak97, Sam91, Tay95]. Theme [Ano94c, MW97]. Theme-Session [Ano94c]. theory [Gre95]. Theragra [BS92, PPS90a, Swa99, Tra96]. thermohaline [MRF97]. thermophilic [Neh98]. thinning [ZM93]. thirty [SKD96]. thirty-year [SKD96]. Three [MJ96, BHS96, CCC^{+}98, Far92, FØK94, MAYC96, MSG96, MC99, Ped96b, Rij91, SVJ99, VGHO97, WHM97]. Three-dimensional [MJ96, Ped96b]. threshold [Rey96]. threshold-induced [Rey96]. thresholds [MRM^{+}94]. throughout [BCJ^{+}97]. Thunnus [BJM99]. tidal [AWEH94, BR97, CG95]. Tilt [HO96]. Time [Gud95, Hee96, VM98, BHKH94, CFUCY95, FA97, FN99, LBDH98, MSPG92, MUV90, MLCB99, Pet97, SS99, dVDP^{+}98]. Time-series [Hee96, VM98, CFUCY95, LBDH98, dVDP^{+}98]. Timing [MJ97b, Bra94, CHH91]. together [PM96b]. Tokyo [OIF95]. tonsa [KB95, THJ^{+}95]. Tool [Ano93b, Bia99, Cla94, HG96, MWRH98]. tooth [GCM99]. top [VR98]. top-down [VR98]. Torungen [DSO96]. total [LL97, Spa90a, SFCP99, Thu97]. tow [MLCB99]. towed [Klo96]. towing [MLCB99]. toxicity [NBBS95]. toxins [HPDR93]. Trace [Par99, Bal93]. tracer [LPL95]. trachurus [BH94, BH94, BTC93, HHA95, VGF97]. tracked [MA99]. Tracking [MFPT98, WRT95, BJM99, ET96]. tractor [CWC^{+}97]. tracts [FVC94]. trade [MSRK99]. trade-offs [MSRK99]. traits [Roc98]. transcarbamylase [Ber95, HLAM95]. transducer [Foo90a, Klo96]. transducers [Foo95, Ham96a, Rey98]. transect [FS93, Køi99]. transects [vdML95]. transferable [SK95a]. transferrin [BTC93]. Transfers [Sin91]. transformation [Har97]. transition [FGS96, Ims99, SOBP99]. Transplantation [God95]. Transplantation-tagging-experiments [God95]. Transport [MW97, RdC95]. trap [DW99]. traps [FA97]. Trawl [Hee96, Kap99, MLCB99, AGKB91, ES92, EJW98, GW93, GKT98, GWE99, Has96, HD90, KN99, LV94, Mil94, MA92a, Nak90, PB97, PV93, RvL96, RBSV98, SE97, Som99, SPCP97, SPCLT96, SNSC94, WM99]. trawler [OR97, WB97a]. trawlers [Ehr91]. trawling [BH94, HSFAM99, Kai98, RBSV98]. trawls [AGKB91, RAFC92]. Trends [HBC^{+}92, KHØ95, RM96a, SKL98, BS98, BC96, Coo97, FN93, HD96a, HHHS92, JG93, Mil96, MUV90, NFM98, OB97, RRL^{+}92, SKD96, WK94, dVDP^{+}98]. Trial [CWC^{+}97, BB94]. tributary [SHM95]. Tributyltin [DBH98]. tridactyla [HW97]. Trieste [CFUCY95, MUCM98, VGF97]. triploid [OMF^{+}97]. Trisopterus [Alb94a, BWH97, HH93]. Trophic [SG95, Cor99b, PTCC98, SB94, SG98]. Trophodynamics [BMM94]. tropical [AR90, JM99, MBR95, PW95, Smi95, Ye98]. Trough [KCB99]. trout [Cus95b, DPHM97, Daw98, Jen91, MLBM98, MH98b, SKHM98, TIR97]. trutta [DB92, DPHM97, Daw98, MLBM98, MH98b, SKHM98, TIR97]. tryptic [UC92]. tuna [BJM99, HLB91, ML96, PKMP99]. Tunicata [MK95]. turbot [Ims99]. turbulence [SWCG94]. Turkish [KN94]. tusk [BMM^{+}98]. twenty [FBG96]. twenty-two [FBG96]. two [BMM94, FBG96, HFP96, HIL98, Jen95, Jen91, KNI92, MSG96, SV98, UC92]. two-frequency [MSG96]. two-species [Jen91]. types [PPS90a]. typing [MEH97, MP99].

UK [Tay95]. ulcerated [WTB99]. ultrasonic [Foo91]. umbo [RCE^{+}93]. unattended [ROH^{+}98]. uncertainties [KSP96]. Uncertainty [SBSP99, DW99, HS91, PL92, SRAL99, Ull96, Wit99]. Undaria [FPW91]. underlying [BEM96]. Understanding [Bia99, TSAS96, WRT95, RTK95, Rym97]. unit [Gré97b]. United [CZM98, DM99, McD98]. universal [Jen95]. unpaired [PM96a]. unrefined [HN91a, HN91b]. unusual [ML96]. unusually [Kap99]. update [MBT95]. Updating [MWRH98]. upper [RTK95]. upwelling [VR98]. USA [RSN96]. Use [BRK97a, Lar92, MSG96, PL97, SH98, HLAM95, Hul95, McG97, MK96, THJ^{+}95]. used [BFNP97, LP91a, LL97, Mil94]. Using [FN99, Spa95, AGKB91, ADS96, Ham96a, Jen90, Kap99, KP97, Løk98, MK98, MMHC98, MJ96, MH97, MB96b, Moh99, NN91b, Nie99, Ped96b, PV93, Rad98, SAC96, SF96, SFG96, Spa90c, SJ96, SK95b, TZD^{+}95, TSAS96, TL96, Wal94, WA90, Ye98, vdML95]. utility [Ham96b]. utilization [BDS96]. utilized [HBA98].

Validation [GMG98, Mok92b, Wal94, LDU98]. Variability [FBG96, GA98, Hou90, RGHC96, WWS93, Col98, DJ98, Dan97, FGS96, GHH98, GKT98, Hic98, Laa96, LS98b, MGMS94, Mis93, MUCM98, Pep90, PB97, PW95, PV93, PCM96, ROH^{+}98, RP98, RDvBH91, Sak97, Sma98, TSP^{+}96, VPT99, dBTT98]. Variable [dBT98, Md95, TA94]. Variables [PL97, CH98, SP96]. Variablity [Cus95a]. variance [Pet97]. Variation [CvdK96, MC99, RCM90, BR97, BCJ^{+}97, Box97, DSO96, DKLM99, Fry91, GBB^{+}97, GMA98, Gud98, HTG97, JB91, KN99, MSPG92, Mac95, MS96, MGF96, MB97, Nil97, PLvdM^{+}96, RG93, SKL98, TJ97]. Variations [BME96, JDLH97, MFPT98, ADS98, ES92, PCH91, RTK95, SRJ98, Wit96, dPVV^{+}96]. varying [Fog93, GW93]. vectors [SS99]. veliger [BB92]. velvet [GGF94]. Venice [HSFAM99]. versus [Mil92]. vertebral [Hul95, NP98, RSB99]. Vertical [AW92, CCS97, FG95a, AEH^{+}99, CNF90, MP91, Ric95, WH97, dPVV^{+}96]. vessel [Coo97, Has96, SFG96]. vessels [SFCP99]. viability [RR99]. vibrational [LR96]. vicinity [SVS99]. video [EMN^{+}94, Gle96, LSP^{+}95]. view [Fos92]. viewing [ARW93]. views [Pay99]. villosus [Fos92, GG98, MP91, Nak90, SCM94]. Vinciguerria [ML96]. virens [BWH97]. virtual [LP91a, VKA98]. visible [LMW^{+}99]. vision [SK95b]. Visual [SP92]. visualization [SWB96]. visualization-based [SWB96]. Volume [Ano90a, Ano90b, Ano91a, Ano91b, Ano91c, Ano91d, Ano92a, Ano92b, Ano93a, Ano93d, Ano94a, Ano94b, Ano95b, Ano95d, Ano96b, Ano98b, Ano99a, Ano97f]. Vosso [SHKL97]. VPA [KV97, LP91a, Spa90b, Spa90c]. vulgaris [LHSA98, LDU98]. vulnerability [CHR96].

Wadden [BV96, MA92b, PLvdM^{+}96, WB97a]. wakes [SE96]. Walbaum [BRK97b, RSB99]. Wales [AGKB91, CWC^{+}97]. walleye [BS92, Swa99, Tra96]. Water [BME96, Har97, LOÅ97, RW91, ADS98, BFKR97, BS98, DLMS97, Hay98, HP99, HN90, Mar93, Mau95, MP91, SS93b]. waters [AG95, AG98, ADS96, Bal93, CHØ92, CDL^{+}94, CBR95, FPW91, GA98, GGFFF98, GCG96, Gud98, Ham96a, HM99, HKN^{+}91, KLF96, Kre94, LLLL99, MRTB95, NJN94, OLO98, Par99, PR93, Ped95, PTO95, PL96, PBSN98, RA97, RCE^{+}93, RP96, Ros98, SKS97, TT96, WB97b, dVDP^{+}98]. way [ZM93]. weather [SG98]. web [GBB^{+}97]. webs [BWH97]. Weddell [SSH95]. weight [BRK97a, LL97, MT97b]. weight-at-age [LL97]. weighted [Far92, MUV90]. weights [SC99]. Well [Hou90, Sut99]. Well-Being [Hou90]. West [WK99, BRK97b, BR93, BI92, Col99, CCC^{+}98, FHA^{+}98, HH93, MRTB95, MM97, MB97, OF95, PU93, PB97, Rai93, SRJ98, Tor91, WWS93, HRL97, HJR96, PR93, Ped95, PK95b, RSB99, vLRvR^{+}96]. Western [RTWC97, AS95, AS97, BMM94, BS92, BME96, Daw98, FPHM90, FL98, KS90, MA99, McD98, NW93, OR97, PLvdM^{+}96, SHKL97, SVS99, SB94, SJK90, SK90a, SK90b, WH98, MP99, SV98]. whale [MRL92, MEH97]. whales [AS95, AS97, Bes93, CHØ92, HGLN95, MP99, NTWW95b, Sam90]. Which [SF96]. whiffiagonis [SPL98]. white [DAP97, NTWW95b, NHP^{+}95, PN93]. white-beaked [NTWW95b]. whitefish [Jen91]. whiting [BWH97, HH93, HRBA91, MMHC98, MT97b, OVP98, Pat98, PM96b, Rob92, Tra96]. whole [Par99]. Wideband [ZMHD96, SAC96]. widely [MSG96]. wild [BMN97, BTS^{+}97, DMAG97, Dem99, EHJN91, FE97, HJL99, HIL98, JG97, Jon97, KP97, MLBM98, McG97, McV97, SJK90]. wind [CvdK96]. winter [CZM98, HRP90, KWN95]. wintering [DRR94]. within [SOBP99]. without [PH94, TA94]. wolffish [FPHM90, GHM97, PN93, PM96a]. Working [Ram91]. works [Sut99]. Workshop [LM99].

year [Daw91, GG98, HW97, SKD96, Ser90, She97]. year-class [She97]. year-class-strength [Ser90]. years [CFUCY95, FBG96, Mut94]. yellow [MA99, Mun97]. yellow- [MA99]. yellow-legged [Mun97]. yellowfin [BJM99, PPS90b]. yield [AR90, Mac95, NR96, Ye98]. yield-per-recruit [Ye98]. yolk [HGDC94]. yolk-sac [HGDC94]. young [AGH^{+}93, AH96, Arr98, SE97].

Zealand [MAYC96, ZF98]. Zn [SKL98]. zone [BC96, FG96, Laa96]. zoobenthos [BEM96]. zooplankters [MJ96]. Zooplankton [Ano94c, Ban95, JG93, LPL95, AG95, AG98, CFUCY95, CSW92, CH98, Cus95a, FH96, GRN96, HP99, HLAM95, HGDC94, Hun97, IMH96, LLLL99, MSWL96, Md95, OLO98, OIF95, PW95, PT98, SSJK98, SWCG94, Sta94, SCW96, SS95b, Swa99, Tay95, VR98, WB95].

References

Andreeva:1996:MSS

[AB96] I. B. Andreeva and A. V. Belousov. Multiple sound scattering by densely packed shoals of marine animals. ICES Journal of Marine Science, 53(2):323–327, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

              academic.oup.com/icesjms/article/53/2/323/792600.

Aure:1996:AES

[ADS96] Jan Aure, Didrik Danielssen, and Roald Sætre. Assessment of eutrophication in Skagerrak coastal waters using oxygen consumption in fjordic basins. ICES Journal of Marine Science, 53(3):589–595, June 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

              com/icesjms/article/53/3/589/625709.

Aure:1998:OSC

[ADS98] Jan Aure, Didrik Danielssen, and Einar Svendsen. The origin of Skagerrak coastal water off arendal in relation to variations in nutrient concentrations. ICES Journal of Marine Science, 55 (4):610–619, August 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

              com/icesjms/article/55/4/610/698142.

Aglen:1999:HVF

[AEH^{+}99] A. Aglen, A. Engås, I. Huse, K. Michalsen, and B. K. Stensholt. How vertical fish distribution may affect survey results. ICES Journal of Marine Science, 56(3):345–360, June 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
56/3/345/623519.

Astthorsson:1995:LTC

[AG95] Olafur S. Astthorsson and Astthor Gislason. Long-term changes in zooplankton biomass in Icelandic waters in spring. ICES Journal of Marine Science, 52(3–4):657–668, June 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/52/3-4/657/841861.

Astthorsson:1998:ECZ

[AG98] Olafur S. Astthorsson and Astthor Gislason. Environmental conditions, zooplankton, and capelin in the waters north of Iceland. ICES Journal of Marine Science, 55(4):808–810, August 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/55/4/808/698245.

Andreasen:1993:PCA

[AGH^{+}93] Thomas Vennø Andreasen, Bjørn Gulliksen, Tore Haug, Einar M. Nilsen, and Einar Ringø. Protein content and amino acid composition of young Atlantic halibut (Hippoglossus hippoglossus L.) captured in the autumn in north Norway. ICES Journal of Marine Science, 50(1):93–100, ???? 1993. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
50/1/93/652893.

Andrew:1991:ETC

[AGKB91] N. L. Andrew, K. J. Graham, S. J. Kennelly, and M. K. Broadhurst. The effects of trawl configuration on the size and composition of catches using benthic prawn trawls off the coast of New South Wales, Australia. ICES Journal of Marine Science, 48 (2):201–209, August 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/48/2/201/644492.

Aglen:1996:IFD

[Agl96] Asgeir Aglen. Impact of fish distribution and species composition on the relationship between acoustic and swept-area estimates of fish density. ICES Journal of Marine Science, 53 (2):501–505, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/53/2/501/792951.

Ange:1992:LSP

[AH92] M. V. Ange and P. M. Hargreaves. Large-scale patterns in the distribution of planktonic and micronektonic biomass in the north-east Atlantic. ICES Journal of Marine Science, 49(4): 403–411, November 1992. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/49/4/403/653320.

Arnold:1995:CSM

[AH95] G. P. Arnold and B. H. Holford. A computer simulation model for predicting rates and scales of movement of demersal fish on the European continental shelf. ICES Journal of Marine Science, 52(6):981–990, December 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/52/6/981/649283.

Arrhenius:1996:GSC

[AH96] F. Arrhenius and S. Hansson. Growth and seasonal changes in energy content of young Baltic Sea herring (Clupea harengus L.). ICES Journal of Marine Science, 53(5):792–801, October 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/53/5/792/704307.

Addison:1991:SCP

[AL91] J. T. Addison and S. R. J. Lovewell. Size composition and pot selectivity in the lobster (Homarus gammarus (L.)) and crab (Cancer pagurus L.) fisheries on the east coast of England. ICES Journal of Marine Science, 48(1):79–90, May 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
48/1/79/696133.

Amara:1995:TAA

[AL95] Rachid Amara and Françoise Lagardère. Taille et âge au début de la métamorphose chez la sole (Solea solea (L.)) du golfe de gascogne. (French) [Size and age at the start of metamorphosis in sole (Solea solea (L.)) from the Bay of Biscay]. ICES Journal of Marine Science, 52(2):247–256, April 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
52/2/247/593908.

Albert:1994:BEN

[Alb94a] Ole Thomas Albert. Biology and ecology of Norway pout (Trisopterus esmarki Nilsson, 1855) in the Norwegian deep. ICES Journal of Marine Science, 51(1):45–61, ???? 1994. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
51/1/45/760301.

Albert:1994:EHM

[Alb94b] Ole Thomas Albert. Ecology of haddock (Melanogrammus aeglefinus L.) in the Norwegian deep. ICES Journal of Marine Science, 51(1):31–44, ???? 1994. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/51/1/31/760298.

Albert:1995:DCF

[Alb95] Ole Thomas Albert. Diel changes in food and feeding of small gadoids on a coastal bank. ICES Journal of Marine Science, 52 (5):873–885, October 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/52/5/873/783892.

Anonymous:1990:IV

[Ano90a] Anonymous. Index to volume 46. ICES Journal of Marine Science, 46(3):343, ???? 1990. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/46/3/343/596428.

Anonymous:1990:RV

[Ano90b] Anonymous. Referees for volume 46. ICES Journal of Marine Science, 46(3):342, ???? 1990. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/46/3/342/596426.

Anonymous:1991:IVa

[Ano91a] Anonymous. Index to volume 47. ICES Journal of Marine Science, 47(3):415–416, ???? 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/47/3/415/601199.

Anonymous:1991:IVb

[Ano91b] Anonymous. Index to volume 48. ICES Journal of Marine Science, 48(3):377–378, November 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/48/3/377/737346.

Anonymous:1991:RVa

[Ano91c] Anonymous. Referees for volume 47. ICES Journal of Marine Science, 47(3):413, ???? 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/47/3/413/601197.

Anonymous:1991:RVb

[Ano91d] Anonymous. Referees for volume 48. ICES Journal of Marine Science, 48(3):375, November 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/48/3/375/737344.

Anonymous:1992:IV

[Ano92a] Anonymous. Index to volume 49. ICES Journal of Marine Science, 49(4):465–466, November 1992. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/49/4/465/653354.

Anonymous:1992:RV

[Ano92b] Anonymous. Referees for volume 49. ICES Journal of Marine Science, 49(4):463, November 1992. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/49/4/463/653349.

Anonymous:1993:IV

[Ano93a] Anonymous. Index to volume 50. ICES Journal of Marine Science, 50(4):509–510, ???? 1993. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/50/4/509/668276.

Anonymous:1993:MSE

[Ano93b] Anonymous. Mini-Symposium on Ecosystem Modelling as a Tool to Predict Pollution-Associated Risks for the Marine Environment. ICES Journal of Marine Science, 50(4):331, ???? 1993. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/50/4/331/668215.

Anonymous:1993:OL

[Ano93c] Anonymous. Open Lecture, 1992. ICES Journal of Marine Science, 50(2):103, ???? 1993. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/50/2/103/774772.

Anonymous:1993:RV

[Ano93d] Anonymous. Referees for volume 50. ICES Journal of Marine Science, 50(4):507, ???? 1993. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/50/4/507/668273.

Anonymous:1994:IV

[Ano94a] Anonymous. Index to volume 51. ICES Journal of Marine Science, 51(4):515, August 1994. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/51/4/515/714270.

Anonymous:1994:RV

[Ano94b] Anonymous. Referees for volume 51. ICES Journal of Marine Science, 51(4):513, August 1994. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/51/4/513/714267.

Anonymous:1994:SPT

[Ano94c] Anonymous. Some of the papers from the theme-session on impact of gelatinous zooplankton predators on coastal and shelf ecosystems held in Dublin on 25 September 1993. ICES Journal of Marine Science, 51(4):345, August 1994. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/51/4/345/714237.

Anonymous:1995:CLC

[Ano95a] Anonymous. Complete list of contributions. ICES Journal of Marine Science, 52(3–4):757–773, June 1995. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/52/3-
4/757/841951.

Anonymous:1995:IV

[Ano95b] Anonymous. Index to volume 52. ICES Journal of Marine Science, 52(6):1013, December 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/52/6/1013/649247.

Anonymous:1995:LR

[Ano95c] Anonymous. List of referees. ICES Journal of Marine Science, 52(3–4):755, June 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.oup.

com/icesjms/article/52/3-4/755/841945.

Anonymous:1995:RV

[Ano95d] Anonymous. Referees for volume 52. ICES Journal of Marine Science, 52(6):1013, December 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/52/6/1013/649251.

Anonymous:1996:CLC

[Ano96a] Anonymous. Complete list of contributions. ICES Journal of Marine Science, 53(2):519–535, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/53/2/519/793007.

Anonymous:1996:IV

[Ano96b] Anonymous. Index to volume 53. ICES Journal of Marine Science, 53(6):1226, December 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/53/6/1226/592510.

Anonymous:1996:LRa

[Ano96c] Anonymous. List of referees. ICES Journal of Marine Science, 53(2):517, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/53/2/517/793000.

Anonymous:1996:LRb

[Ano96d] Anonymous. List of referees. ICES Journal of Marine Science, 53(6):1226, December 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/53/6/1226/592512.

Anonymous:1996:PRB

[Ano96e] Anonymous. Professor Ray Beverton 29 August 1922–23 July 1995. ICES Journal of Marine Science, 53(1):1–9, February 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/53/1/1/731536.

Anonymous:1997:Ea

[Ano97a] Anonymous. Erratum. ICES Journal of Marine Science, 54 (1):156, February 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/54/1/156/782172.

Anonymous:1997:Eb

[Ano97b] Anonymous. Erratum. ICES Journal of Marine Science, 54 (1):157, February 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/54/1/157/782174.

Anonymous:1997:Ec

[Ano97c] Anonymous. Erratum. ICES Journal of Marine Science, 54(3):504, June 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/54/3/504/652502.

Anonymous:1997:Ed

[Ano97d] Anonymous. Erratum. ICES Journal of Marine Science, 54 (5):960, October 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/54/5/960/751577.

Anonymous:1997:Ee

[Ano97e] Anonymous. Erratum. ICES Journal of Marine Science, 54 (5):961, October 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/54/5/961/751580.

Anonymous:1997:IV

[Ano97f] Anonymous. Index to volume 54. ICES Journal of Marine Science, 54(6):1226, December 1997. CODEN ICESEC.

ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/54/6/1226/674056.

Anonymous:1997:LRb

[Ano97g] Anonymous. List of referee. ICES Journal of Marine Science, 54(6):1226, December 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/54/6/1226/674064.

Anonymous:1997:LRa

[Ano97h] Anonymous. List of referees. ICES Journal of Marine Science, 54(4):739, August 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/54/4/739/607520.

Anonymous:1998:E

[Ano98a] Anonymous. Erratum. ICES Journal of Marine Science, 55 (6):1153, December 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/55/6/1153/639003.

Anonymous:1998:IV

[Ano98b] Anonymous. Index to volume 55. ICES Journal of Marine Science, 55(6):1153, December 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/55/6/1153/639005.

Anonymous:1998:LRa

[Ano98c] Anonymous. List of referees. ICES Journal of Marine Science, 55(4):824, August 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/55/4/824/698269.

Anonymous:1998:LRb

[Ano98d] Anonymous. List of referees. ICES Journal of Marine Science, 55(6):1152, December 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/55/6/1152/638999.

Anonymous:1999:IV

[Ano99a] Anonymous. Index to volume 56. ICES Journal of Marine Science, 56(6):1074, December 1999. CODEN ICESEC.

ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/56/6/1074/658115.

Anonymous:1999:LR

[Ano99b] Anonymous. List of referees. ICES Journal of Marine Science, 56(6):1073–1074, December 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/56/6/1073/658113.

Abud:1992:GEB

[APD92] Eddie O. Aristizabal Abud, Leszek B. Prenski, and Gustavo R. Daleo. Growth and energy budget in juvenile croaker (Micropogonias furnieri (Desmarest 1823)). ICES Journal of Marine Science, 49(1):65–68, February 1992. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/49/1/65/732785.

Alcala:1990:DTE

[AR90] A. C. Alcala and G. R. Russ. A direct test of the effects of protective management on abundance and yield of tropical marine resources. ICES Journal of Marine Science, 47(1):40–47, ???? 1990. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/47/1/40/638514.

Arrhenius:1996:DCF

[Arr96] Fredrik Arrhenius. Diet composition and food selectivity of 0group herring (Clupea harengus L.) and sprat (Sprattus sprattus (L.)) in the northern Baltic Sea. ICES Journal of Marine Science, 53(4):701–712, August 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/53/4/701/799732.

Arrhenius:1998:FIS

[Arr98] F. Arrhenius. Food intake and seasonal changes in energy content of young Baltic Sea sprat (Sprattus sprattus L.). ICES Journal of Marine Science, 55(2):319–324, April 1998. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
55/2/319/781774.

Ashford:1993:MPL

[ARW93] J. R. Ashford, K. Robinson, and M. G. White. A method for preparing large numbers of otolith sections for viewing by scanning electron microscope. ICES Journal of Marine Science, 50(2):227–229, ???? 1993. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/50/2/227/774842.

Abend:1995:DRS

[AS95] Alan G. Abend and Tim D. Smith. Differences in ratios of stable isotopes of nitrogen in long-finned pilot whales (Globicephala melas) in the western and eastern North Atlantic. ICES Journal of Marine Science, 52(5):837–841, October 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
52/5/837/783873.

Abend:1997:DSI

[AS97] Alan G. Abend and Tim D. Smith. Differences in stable isotope ratios of carbon and nitrogen between long-finned pilot whales (Globicephala melas) and their primary prey in the western north Atlantic. ICES Journal of Marine Science, 54 (3):500–503, June 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/54/3/500/652500.

Atkinson:1995:OFS

[Atk95] A. Atkinson. Omnivory and feeding selectivity in five copepod species during spring in the Bellingshausen Sea, Antarctica. ICES Journal of Marine Science, 52(3–4):385–396, June 1995. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/52/3-4/385/841540.

Arnold:1992:VMC

[AW92] G. P. Arnold and M. Greer Walker. Vertical movements of cod (Gadus morhua L.) in the open sea and the hydrostatic function of the swimbladder. ICES Journal of Marine Science, 49(3): 357–372, August 1992. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/49/3/357/819860.

Arnold:1994:MCG

[AWEH94] G. P. Arnold, M. Greer Walker, L. S. Emerson, and B. H. Holford. Movements of cod (Gadus morhua L.) in relation to the tidal streams in the southern North Sea. ICES Journal of Marine Science, 51(2):207–232, ???? 1994. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/51/2/207/674484.

Alvarez:1999:EFS

[AY99] A. Alvarez and Z. Ye. Effects of fish school structures on acoustic scattering. ICES Journal of Marine Science, 56(3):361–369, June 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/56/3/361/623522.

Balls:1993:IBS

[Bal93] P. Balls. ICES baseline survey of trace metals in European shelf waters. ICES Journal of Marine Science, 50(4):435–444, ???? 1993. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/50/4/435/668241.

Banse:1995:ZPR

[Ban95] Karl Banse. Zooplankton: Pivotal role in the control of ocean production: I. Biomass and production. ICES Journal of Marine Science, 52(3–4):265–277, June 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https:/

/academic.oup.com/icesjms/article/52/3-4/265/841465.

Barbini:1998:DAL

[Bar98] R. Barbini. Design and application of a lidar fluorosensor system for remote monitoring of phytoplankton. ICES Journal of Marine Science, 55(4):793–802, August 1998. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
55/4/793/698235.

Basson:1999:IEF

[Bas99] M. Basson. The importance of environmental factors in the design of management procedures. ICES Journal of Marine Science, 56(6):933–942, December 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/56/6/933/658149.

Bax:1998:SPP

[Bax98] N. J. Bax. The significance and prediction of predation in marine fisheries. ICES Journal of Marine Science, 55(6):997– 1030, December 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/55/6/997/639013.

Beaumont:1992:AVL

[BB92] A. R. Beaumont and D. A. Barnes. Aspects of veliger larval growth and byssus drifting of the spat of Pecten maximus and Aequipecten (Chlamys) opercularis. ICES Journal of Marine Science, 49(4):417–423, November 1992. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/49/4/417/653330.

Bergeron:1994:RAA

[BB94] Jean-Pierre Bergeron and Michel Boulhic. Rapport ARN/ADN et évaluation de l’état nutritionnel et de la croissance des larves de poissons marins: un essai de mise au point expérimentale chez la sole (Solea solea L.). (French) [the RNA /DNA Report and assessment of nutritional status and growth of marine fish larvae: an experimental development trial in sole (Solea solea L.)]. ICES Journal of Marine Science, 51(2):181–190, ???? 1994. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/51/2/181/674481.

Baretta:1998:DNE

[BBBR98] J. W. Baretta, J. G. Baretta-Bekker, and P. Ruardij. Data needs for ecosystem modelling. ICES Journal of Marine Science, 55(4):756–766, August 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/55/4/756/698209.

Bergeron:1991:EPN

[BBG91] Jean-Pierre Bergeron, Michel Boulhic, and Robert Galois. Effet de la privation de nourriture sur la teneur en ADN de la larve de sole (Solea solea L.). (French) [Effect of food deprivation on DNA content of sole larvae (Solea solea L.)]. ICES Journal of Marine Science, 48(1):127–134, May 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/48/1/127/696094.

Bot:1996:MEP

[BC96] Peter V. M. Bot and Franciscus Colijn. A method for estimating primary production from chlorophyll concentrations with results showing trends in the Irish Sea and the Dutch coastal zone. ICES Journal of Marine Science, 53(6):945–950, December 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/53/6/945/592531.

Bembo:1996:EAS

[BCCP96] D. G. Bembo, G. R. Carvalho, N. Cingolani, and T. J. Pitcher. Electrophoretic analysis of stock structure in Northern Mediterranean anchovies, Engraulis encrasicolus. ICES Journal of Marine Science, 53(1):115–128, February 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
53/1/115/731545.

Bourke:1997:AVP

[BCJ^{+}97] E. A. Bourke, J. Coughlan, H. Jansson, P. Galvin, and T. F. Cross. Allozyme variation in populations of Atlantic salmon located throughout Europe: diversity that could be compromised by introductions of reared fish. ICES Journal of Marine Science, 54(6):974–985, December 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/54/6/974/674080.

Beall:1994:DPS

[BDC^{+}94] E. Beall, J. Dumas, D. Claireaux, L. Barrière, and C. Marty. Dispersal patterns and survival of Atlantic salmon (Salmo salar L.) juveniles in a nursery stream. ICES Journal of Marine Science, 51(1):1–9, ???? 1994. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/51/1/1/760277.

Baldursson:1996:RUI

[BDS96] Fridrik M. Baldursson, Ásgeir Daníelsson, and Gunnar Stefánsson. On the rational utilization of the Icelandic cod stock. ICES Journal of Marine Science, 53(4):643–658, August 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/53/4/643/799705.

Bax:1990:MAB

[BE90] Nicholas Bax and Jens-Eric Eliassen. Multispecies analysis in Balsfjord, northern Norway: solution and sensitivity analysis of a simple ecosystem model. ICES Journal of Marine Science, 47(2):175–204, ???? 1990. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/47/2/175/611326.

Beaumont:1991:ADS

[Bea91] A. R. Beaumont. Allozyme data and scallop stock identification. ICES Journal of Marine Science, 47(3):333–338, ???? 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/47/3/333/601174.

Basford:1993:INS

[BED^{+}93] D. J. Basford, A. Eleftheriou, I. M. Davies, G. Irion, and T. Soltwedel. The ICES North Sea benthos survey: the sedimentary environment. ICES Journal of Marine Science, 50 (1):71–80, ???? 1993. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/50/1/71/652888.

Belin:1998:FPM

[Bel98] Catherine Belin. French phytoplankton monitoring: an exploration of optimum data presentation. ICES Journal of Marine Science, 55(4):705–710, August 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/55/4/705/698162.

Beukema:1996:GSS

[BEM96] J. J. Beukema, K. Essink, and H. Michaelis. The geographic scale of synchronized fluctuation patterns in zoobenthos populations as a key to underlying factors: climatic or maninduced. ICES Journal of Marine Science, 53(6):964–971, December 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/53/6/964/592537.

Bergstad:1993:DPS

[Ber93] O. A. Bergstad. Distribution, population structure, growth, and reproduction of the greater silver smelt, Argentina silus (Pisces, Argentinidae), of the Skagerrak and the north-eastern North Sea. ICES Journal of Marine Science, 50(2):129–143, ???? 1993. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/50/2/129/774787.

Bergeron:1995:ATA

[Ber95] Jean-Pierre Bergeron. Aspartate transcarbamylase activity for the assessment of mesozooplankton production: new aspects from oceanic areas. ICES Journal of Marine Science, 52(3– 4):305–313, June 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.oup.

com/icesjms/article/52/3-4/305/841491.

Best:1993:IRS

[Bes93] Peter B. Best. Increase rates in severely depleted stocks of baleen whales. ICES Journal of Marine Science, 50(2):169–186, ???? 1993. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/50/2/169/774803.

Beverton:1991:DJA

[Bev91a] R. J. H. Beverton. Dr John Alan Gulland, FRS 26 September 1926–24 June 1990. ICES Journal of Marine Science, 47 (3):411–412, ???? 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/47/3/411/601195.

Beverton:1991:R

[Bev91b] R. J. H. Beverton. Review. ICES Journal of Marine Science, 48(1):135, May 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/48/1/135/696098.

Backhaus:1997:FEW

[BFKR97] Jan O. Backhaus, Hermann Fohrmann, Jochen Kämpf, and Angelo Rubino. Formation and export of water masses produced in Arctic shelf polynyas — process studies of oceanic convection. ICES Journal of Marine Science, 54(3):366–382, June 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/54/3/366/652473.

Bewers:1997:MSA

[BFL97] J. M. Bewers, L. Føyn, and H. Loeng. Mini-Symposium on Arctic Oceanographic Processes, 1995. ICES Journal of Marine Science, 54(3):281, June 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/54/3/281/652444.

Black:1997:EIU

[BFNP97] Kenneth D. Black, Stuart Fleming, Thomas D. Nickell, and Paula M. F. Pereira. The effects of ivermectin, used to control sea lice on caged farmed salmonids, on infaunal polychaetes. ICES Journal of Marine Science, 54(2):276–279, April 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
54/2/276/620559.

Bergstad:1993:FRA

[BG93] O. A. Bergstad and J. D. M. Gordon. First record of Atlantic halibut (Hippoglussus hippoglussus (L.)) larvae from the Skagerrak. ICES Journal of Marine Science, 50(2):231–232, ???? 1993. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/50/2/231/774845.

Bogstad:1994:MEC

[BG94] Bjarte Bogstad and Harald Gjøsæter. A method for estimating the consumption of capelin by cod in the Barents Sea. ICES Journal of Marine Science, 51(3):273–280, ???? 1994. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
51/3/273/620190.

Briggs:1997:CPI

[BGA97] Kenneth T. Briggs, M. Eric Gershwin, and Daniel W. Anderson. Consequences of petrochemical ingestion and stress on the immune system of seabirds. ICES Journal of Marine Science, 54(4):718–725, August 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/54/4/718/607510.

Bergman:1992:DEB

[BH92] M. J. N. Bergman and M. Hup. Direct effects of beamtrawling on macrofauna in a sandy sediment in the southern North Sea.

ICES Journal of Marine Science, 49(1):5–11, February 1992. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
49/1/5/732782.

Barange:1994:ITS

[BH94] Manuel Barange and Ian Hampton. Influence of trawling on in situ estimates of Cape horse mackerel (Trachurus trachurus capensis) target strength. ICES Journal of Marine Science, 51 (1):121–126, ???? 1994. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/51/1/121/760288.

Bergstad:1998:DHC

[BH98] O. A. Bergstad and Å. S. Høines. Do herring (Clupea harengus L.) spawners influence the migratory tendency of cod (Gadus morhua L.)? ICES Journal of Marine Science, 55(2):325–328, April 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/55/2/325/781778.

Backhaus:1994:HCS

[BHKH94] Jan O. Backhaus, Ingo H. Harms, Michael Krause, and Michael R. Heath. An hypothesis concerning the space-time succession of Calanus finmarchicus in the northern North Sea. ICES Journal of Marine Science, 51(2):169–180, ???? 1994. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
51/2/169/674475.

Barange:1996:EDS

[BHS96] Manuel Barange, Ian Hampton, and Michael Soule. Empirical determination of in situ target strengths of three loosely aggregated pelagic fish species. ICES Journal of Marine Science, 53(2):225–232, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/53/2/225/792429.

Brunetti:1992:DAE

[BI92] Norma E. Brunetti and Marcela L. Ivanovic. Distribution and abundance of early life stages of squid (Illex argentinus) in the south-west Atlantic. ICES Journal of Marine Science, 49 (2):175–183, May 1992. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/49/2/175/683886.

Biais:1999:UFF

[Bia99] G. Biais. Understanding the functioning of fishing enterprises: an essential tool in fisheries management. ICES Journal of Marine Science, 56(6):1044–1050, December 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
56/6/1044/658105.

Bertrand:1999:SAT

[BJM99] A. Bertrand, E. Josse, and J. Massé. In situ acoustic target-strength measurement of bigeye (Thunnus obesus) and yellowfin tuna (Thunnus albacares) by coupling split-beam echosounder observations and sonic tracking. ICES Journal of Marine Science, 56(1):51–60, February 1999. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
56/1/51/652201.

Barrett:1996:RRC

[BK96] Robert T. Barrett and Yuri V. Krasnov. Recent responses to changes in stocks of prey species by seabirds breeding in the southern Barents Sea. ICES Journal of Marine Science, 53 (4):713–722, August 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/53/4/713/799739.

Blaxter:1993:BBP

[Bla93] J. H. S. Blaxter. Basil B. Parrish (25 March 1922–28 June 1993). ICES Journal of Marine Science, 50(3):329–330, ???? 1993. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/50/3/329/685322.

Blaxter:1998:R

[Bla98] J. H. S. Blaxter. Review. ICES Journal of Marine Science, 55(5):978, October 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/55/5/978/691016.

Bogovski:1999:HEL

[BLM99] S. Bogovski, T. Lang, and S. Mellergaard. Histopathological examinations of liver nodules in flounder (Platichthys flesus L.) from the Baltic Sea. ICES Journal of Marine Science, 56 (2):148–151, April 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/56/2/148/622445.

Basford:1996:VBN

[BME96] D. J. Basford, D. C. Moore, and A. S. Eleftheriou. Variations in benthos in the north-western North Sea in relation to the inflow of Atlantic water, 1980–1984. ICES Journal of Marine Science, 53(6):957–963, December 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/53/6/957/592536.

Baamstedt:1994:TTS

[BMM94] Ulf Båmstedt, M. B. Martinussen, and S. Matsakis. Trophodynamics of the two scyphozoan jellyfishes, Aurelia aurita and Cyanea capillata, in western Norway. ICES Journal of Marine Science, 51(4):369–382, August 1994. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/51/4/369/714243.

Bergstad:1998:IAR

[BMM^{+}98] O. A. Bergstad, J. V. Magnusson, J. Magnusson, N.-R. Hareide, and J. Reinert. Intercalibration of age readings of ling (Molva molva L.) blue ling (Molva dipterygia Pennant, 1784) and tusk (Brosme brosme L.). ICES Journal of Marine Science, 55 (2):309–318, April 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/55/2/309/781768.

Beamish:1997:HWP

[BMN97] R. J. Beamish, C. Mahnken, and C. M. Neville. Hatchery and wild production of Pacific salmon in relation to large-scale, natural shifts in the productivity of the marine environment. ICES Journal of Marine Science, 54(6):1200–1215, December 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
54/6/1200/674023.

Bailey:1998:CSD

[BMS98] M. C. Bailey, C. D. Maravelias, and E. J. Simmonds. Changes in the spatial distribution of autumn spawning herring (Clupea harengus L.) derived from annual acoustic surveys during the period 1984–1996. ICES Journal of Marine Science, 55 (3):545–555, June 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/55/3/545/655092.

Boddeke:1996:CBS

[Bod96] R. Boddeke. Changes in the brown shrimp (Crangon crangon L.) population off the Dutch coast in relation to fisheries and phosphate discharge. ICES Journal of Marine Science, 53(6): 995–1002, December 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/53/6/995/592548.

Boxaspen:1997:GTV

[Box97] K. Boxaspen. Geographical and temporal variation in abundance of salmon lice (Lepeophtheirus salmonis) on salmon (Salmo salar L.). ICES Journal of Marine Science, 54(6): 1144–1147, December 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/54/6/1144/673988.

Bakun:1990:CSC

[BP90] Andrew Bakun and Richard H. Parrish. Comparative studies of coastal pelagic fish reproductive habitats: the Brazilian sardine (Sardinella aurita). ICES Journal of Marine Science, 46 (3):269–283, ???? 1990. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/46/3/269/596402.

Bakun:1991:CSC

[BP91] Andrew Bakun and Richard H. Parrish. Comparative studies of coastal pelagic fish reproductive habitats: the anchovy (Engraulis anchoita) of the southwestern Atlantic. ICES Journal of Marine Science, 48(3):343–361, November 1991. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
48/3/343/737304.

Becker:1996:SST

[BP96] Gerd A. Becker and Manfred Pauly. Sea surface temperature changes in the North Sea and their causes. ICES Journal of Marine Science, 53(6):887–898, December 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
53/6/887/592518.

Butterworth:1999:EEI

[BP99] D. S. Butterworth and A. E. Punt. Experiences in the evaluation and implementation of management procedures. ICES Journal of Marine Science, 56(6):985–998, December 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
56/6/985/658160.

Blaxter:1992:E

[BR92] John Blaxter and John Ramster. Editorial. ICES Journal of Marine Science, 49(1):1–2, February 1992. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/49/1/1/732751.

Brodziak:1993:MAS

[BR93] J. K. T. Brodziak and A. A. Rosenberg. A method to assess squid fisheries in the north-west Atlantic. ICES Journal of Marine Science, 50(2):187–194, ???? 1993. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/50/2/187/774813.

Begg:1997:SVS

[BR97] Graham S. Begg and James B. Reid. Spatial variation in seabird density at a shallow sea tidal mixing front in the Irish Sea. ICES Journal of Marine Science, 54(4):552–565, August 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/54/4/552/607444.

Brander:1994:LTC

[Bra94] K. M. Brander. The location and timing of cod spawning around the British Isles. ICES Journal of Marine Science, 51(1):71–89, ???? 1994. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/51/1/71/760307.

Brander:1995:ETG

[Bra95] K. M. Brander. The effect of temperature on growth of Atlantic cod (Gadus morhua L.). ICES Journal of Marine Science, 52 (1):1–10, February 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/52/1/1/699719.

Brander:1999:LE

[Bra99] Keith Brander. Letter to the Editors. ICES Journal of Marine Science, 56(1):125, February 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/56/1/125/652189.

Bergeron:1997:UCR

[BRK97a] Jean-Pierre Bergeron, Jeannine Person-Le Ruyet, and Constantin Koutsikopoulos. Use of carbon rather than dry weight to assess the DNA content and nutritional condition index of sole larvae. ICES Journal of Marine Science, 54(1):148– 151, February 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/54/1/148/782163.

Boje:1997:HPB

[BRK97b] J. Boje, F. Riget, and M. Køie. Helminth parasites as biological tags in population studies of Greenland halibut (Reinhardtius hippoglossoides (Walbaum)), in the north-west Atlantic. ICES Journal of Marine Science, 54(5):886–895, October 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/54/5/886/751540.

Blaxter:1996:PJ

[BRS96] J. H. S. Blaxter, J. W. Ramster, and S. J. Smith. Platinum jubilee. ICES Journal of Marine Science, 53(5):749, October 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/53/5/749/704261.

Bailey:1992:CLA

[BS92] Kevin M. Bailey and Stella M. Spring. Comparison of larval, age-0 juvenile and age-2 recruit abundance indices of walleye pollock, Theragra chalcogramma, in the western Gulf of Alaska.

ICES Journal of Marine Science, 49(3):297–304, August 1992. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
49/3/297/819775.

Borkman:1998:LTT

[BS98] David G. Borkman and Theodore J. Smayda. Long-term trends in water clarity revealed by secchi-disk measurements in lower Narragansett Bay. ICES Journal of Marine Science, 55(4): 668–679, August 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/55/4/668/698154.

Bogovski:1999:CBC

[BSMK99] S. Bogovski, B. Sergeyev, V. Muzyka, and S. Karlova. Correlations between cytochrome P450, haem synthesis enzymes, and aromatic compounds in flounder (Platichthys flesus) from the Baltic Sea. ICES Journal of Marine Science, 56(2):152–156, April 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/56/2/152/622449.

Borges:1993:PTP

[BTC93] M. F. Borges, R. J. Turner, and J. Casey. Plasma transferrin polymorphisms in scad (Trachurus trachurus L.) populations from the north-east Atlantic. ICES Journal of Marine Science, 50(3):299–301, ???? 1993. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/50/3/299/685313.

Berejikian:1997:RBI

[BTS^{+}97] B. A. Berejikian, E. P. Tezak, S. L. Schroder, C. M. Knudsen, and J. J. Hard. Reproductive behavioral interactions between wild and captively reared Coho salmon (Oncorhynchus kisutch). ICES Journal of Marine Science, 54(6):1040–1050, December 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/54/6/1040/673915.

Boddeke:1996:ARW

[BV96] R. Boddeke and B. Vingerhoed. The anchovy returns to the Wadden Sea. ICES Journal of Marine Science, 53(6):1003– 1007, December 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/53/6/1003/592453.

Batchelder:1995:IBM

[BW95] Harold P. Batchelder and Robert Williams. Individual-based modelling of the population dynamics of Metridia lucens in the North Atlantic. ICES Journal of Marine Science, 52(3– 4):469–482, June 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.oup.

com/icesjms/article/52/3-4/469/841602.

Bromley:1997:DFP

[BWH97] P. J. Bromley, T. Watson, and J. R. G. Hislop. Diel feeding patterns and the development of food webs in pelagic 0-group cod (Gadus morhua L.), haddock (Melanogrammus aeglefinus L.), whiting (Merlangius merlangus L.), saithe (Pollachius virens L.), and Norway pout (Trisopterus esmarkii Nilsson) in the northern North Sea. ICES Journal of Marine Science, 54(5): 846–853, October 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/54/5/846/751500.

Castonguay:1999:DCA

[Cas99] M. Castonguay. Distribution changes of Atlantic cod (Gadus morhua L.) in the northern Gulf of St Lawrence in relation to an oceanic cooling. ICES Journal of Marine Science, 56 (3):333–344, June 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/56/3/333/623516.

Carr:1997:OSC

[CAWD97] J. W. Carr, J. M. Anderson, F. G. Whoriskey, and T. Dilworth. The occurrence and spawning of cultured Atlantic salmon (Salmo salar) in a Canadian river. ICES Journal of Marine Science, 54(6):1064–1073, December 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/54/6/1064/673924.

Chauvelon:1993:MOS

[CB93] P. Chauvelon and P. Bach. Modelling the otolith shape as an ellipse: an attempt for back-calculation purposes. ICES Journal of Marine Science, 50(2):121–128, ???? 1993. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
50/2/121/774781.

Castillo:1996:RBS

[CBG96] J. Castillo, M. A. Barbieri, and A. Gonzalez. Relationships between sea surface temperature, salinity, and pelagic fish distribution off northern Chile. ICES Journal of Marine Science, 53(2):139–146, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/53/2/139/791578.

Collins:1995:RMS

[CBR95] M. A. Collins, G. M. Burnell, and P. G. Rodhouse. Recruitment, maturation, and spawning of Loligo forbesi Steenstrup (Cephalopoda: Loliginidae) in Irish waters. ICES Journal of Marine Science, 52(1):127–137, February 1995. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
52/1/127/699745.

Cornelisse:1993:ICN

[CC93] Kenneth J. Cornelisse and Kim Diget Christensen. Investigation of a cover net designed to reduce southern cormorant (Phalacrocorax carbo sinensis) fisheries depredation in a pound net. ICES Journal of Marine Science, 50(3):279–284, ???? 1993. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
50/3/279/685308.

Caddy:1999:SAF

[CC99] J. F. Caddy and F. Carocci. The spatial allocation of fishing intensity by port-based inshore fleets: a GIS application. ICES Journal of Marine Science, 56(3):388–403, June 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
56/3/388/623530.

Costelloe:1998:DLS

[CCC^{+}98] M. Costelloe, J. Costelloe, N. Coghlan, G. O’Donohoe, and B. O’Connor. Distribution of the larval stages of Lepeophtheirus salmonis in three bays on the west coast of Ireland. ICES Journal of Marine Science, 55(2):181–187, April 1998.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
55/2/181/781702.

Conway:1997:VDF

[CCS97] D. V. P. Conway, S. H. Coombs, and C. Smith. Vertical distribution of fish eggs and larvae in the Irish Sea and southern North Sea. ICES Journal of Marine Science, 54(1):136– 147, February 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/54/1/136/782158.

Chesson:1999:EFM

[CCW99] J. Chesson, H. Clayton, and B. Whitworth. Evaluation of fisheries-management systems with respect to sustainable development. ICES Journal of Marine Science, 56(6):980– 984, December 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/56/6/980/658158.

Collins:1994:DSL

[CDL^{+}94] M. A. Collins, S. De Grave, C. Lordan, G. M. Burnell, and P. G. Rodhouse. Diet of the squid Loligo forbesi Steenstrup (Cephalopoda: Loliginidae) in Irish waters. ICES Journal of Marine Science, 51(3):337–344, ???? 1994. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/51/3/337/620227.

Cataletto:1995:EYT

[CFUCY95] B. Cataletto, E. Feoli, S. Fonda Umani, and Sun Cheng-Yong. Eleven years of time-series analysis on the net-zooplankton community in the Gulf of Trieste. ICES Journal of Marine Science, 52(3–4):669–678, June 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https:/

/academic.oup.com/icesjms/article/52/3-4/669/841870.

Castonguay:1995:ETS

[CG95] Martin Castonguay and Denis Gilbert. Effects of tidal streams on migrating Atlantic mackerel, Scomber scombrus L. ICES Journal of Marine Science, 52(6):941–954, December 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
52/6/941/649271.

Camphuysen:1997:EDS

[CG97] Kees (C. J.) Camphuysen and Stefan Garthe. An evaluation of the distribution and scavenging habits of northern fulmars (Fulmarus glacialis) in the North Sea. ICES Journal of Marine Science, 54(4):654–683, August 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/54/4/654/607481.

Conversi:1998:CSB

[CH98] Alessandra Conversi and Sultan Hameed. Common signals between physical and atmospheric variables and zooplankton biomass in the Subarctic Pacific. ICES Journal of Marine Science, 55(4):739–747, August 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/55/4/739/698190.

Child:1991:DPT

[CHH91] A. R. Child, B. R. Howell, and R. G. Houghton. Daily periodicity and timing of the spawning of sole, Solea solea (L.), in the Thames estuary. ICES Journal of Marine Science, 48(3): 317–323, November 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/48/3/317/737260.

Christensen:1992:SDE

[CHØ92] Ivar Christensen, Tore Haug, and Nils Øien. Seasonal distribution, exploitation and present abundance of stocks of large baleen whales (Mysticeti) and sperm whales (Physeter macrocephalus) in Norwegian and adjacent waters. ICES Journal of Marine Science, 49(3):341–355, August 1992. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
49/3/341/819813.

Christensen:1995:PBI

[Chr95] Steen Christensen. Potential bioeconomic impact of reduced mortality of codend escapees of the shrimp fishery in the Davis Strait. ICES Journal of Marine Science, 52(5):843–851, October 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/52/5/843/783876.

Cowan:1996:SDV

[CHR96] James H. Cowan, Jr., Edward D. Houde, and Kenneth A. Rose. Size-dependent vulnerability of marine fish larvae to predation: an individual-based numerical experiment. ICES Journal of Marine Science, 53(1):23–37, February 1996. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
53/1/23/731547.

Christensen:1997:EMS

[Chr97] Steen Christensen. Evaluation of management strategies — a bioeconomic approach applied to the Greenland shrimp fishery. ICES Journal of Marine Science, 54(3):412–426, June 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
54/3/412/652481.

Cook:1991:QNS

[CKF91] R. M. Cook, P. A. Kunzlik, and R. J. Fryer. On the quality of North Sea cod stock forecasts. ICES Journal of Marine Science, 48(1):1–13, May 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/48/1/1/696081.

Claereboudt:1994:GAT

[Cla94] Michel R. Claereboudt. GMDH algorithm as a tool for bivalve growth analysis and prediction. ICES Journal of Marine Science, 51(4):439–445, August 1994. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/51/4/439/714255.

Carr:1997:MNM

[CLAD97] J. W. Carr, G. L. Lacroix, J. M. Anderson, and T. Dilworth. Movements of non-maturing cultured Atlantic salmon (Salmo salar) in a Canadian river. ICES Journal of Marine Science, 54 (6):1082–1085, December 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/54/6/1082/673934.

Chaput:1992:EEF

[CLB92] G. J. Chaput, C. H. LeBlanc, and C. Bourque. Evaluation of an electronic fish measuring board. ICES Journal of Marine Science, 49(3):335–339, August 1992. CODEN ICESEC.

ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/49/3/335/819805.

Campana:1991:APA

[CM91] Steven E. Campana and Erlend Moksness. Accuracy and precision of age and hatch date estimates from otolith microstructure examination. ICES Journal of Marine Science, 48(3): 303–316, November 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/48/3/303/737254.

Coombs:1990:PEP

[CNF90] S. H. Coombs, J. H. Nichols, and C. A. Fosh. Plaice eggs (Pleuronectes platessa L.) in the southern North Sea: abundance, spawning area, vertical distribution, and buoyancy. ICES Journal of Marine Science, 47(2):133–139, ???? 1990. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
47/2/133/611296.

Cochrane:1999:CFL

[Coc99] K. L. Cochrane. Complexity in fisheries and limitations in the increasing complexity of fisheries management. ICES Journal of Marine Science, 56(6):917–926, December 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
56/6/917/658143.

Colijn:1998:TVP

[Col98] Franciscus Colijn. The temporal variability of plankton and their physico-chemical environment. ICES Journal of Marine Science, 55(4):557–561, August 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/55/4/557/698129.

Collins:1999:RMC

[Col99] M. Collins. Resolution of multiple cohorts in the Loligo forbesi population from the west of Scotland. ICES Journal of Marine Science, 56(4):500–509, August 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/56/4/500/624653.

Cook:1997:STS

[Coo97] R. M. Cook. Stock trends in six North Sea stocks as revealed by an analysis of research vessel surveys. ICES Journal of Marine Science, 54(5):924–933, October 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/54/5/924/751564.

Cook:1998:SCE

[Coo98] R. M. Cook. A sustainability criterion for the exploitation of North Sea cod. ICES Journal of Marine Science, 55(6): 1061–1070, December 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/55/6/1061/638973.

Cooke:1999:IFM

[Coo99] J. G. Cooke. Improvement of fishery-management advice through simulation testing of harvest algorithms. ICES Journal of Marine Science, 56(6):797–810, December 1999. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
56/6/797/658119.

Corten:1999:PMB

[Cor99a] A. Corten. A proposed mechanism for the Bohuslän herring periods. ICES Journal of Marine Science, 56(2):207–220, April 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/56/2/207/622489.

Cortes:1999:SDC

[Cor99b] E. Cortés. Standardized diet compositions and trophic levels of sharks. ICES Journal of Marine Science, 56(5):707–717, October 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/56/5/707/691331.

Casas:1996:RCF

[CP96] J. M. Casas and J. Paz. Recent changes in the feeding of cod (Gadus morhua) off the Flemish Cap, Newfoundland 1989– 1993. ICES Journal of Marine Science, 53(5):750–756, October 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/53/5/750/704267.

Cook:1993:ANS

[CR93] R. M. Cook and S. A. Reeves. Assessment of North Sea industrial fish stocks with incomplete catch-at-age data. ICES Journal of Marine Science, 50(4):425–434, ???? 1993. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
50/4/425/668239.

Chapdelaine:1997:RBC

[CR97] Gilles Chapdelaine and Jean-François Rail. Relationship between cod fishery activities and the population of herring gulls on the North Shore of the Gulf of St Lawrence, Québec, Canada. ICES Journal of Marine Science, 54(4):708–713, August 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/54/4/708/607498.

Cass:1999:LHM

[CR99] A. Cass and B. Riddell. A life history model for assessing alternative management policies for depressed Chinook salmon. ICES Journal of Marine Science, 56(4):414–421, August 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
56/4/414/624623.

Chu:1992:FDS

[CSW92] D. Chu, T. K. Stanton, and P. H. Wiebe. Frequency dependence of sound backscattering from live individual zooplankton. ICES Journal of Marine Science, 49(1):97–106, February 1992. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
49/1/97/732792.

Cushing:1990:R

[Cus90] D. H. Cushing. Review. ICES Journal of Marine Science, 47 (2):246–247, ???? 1990. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/47/2/246/611367.

Cushing:1992:SHD

[Cus92] D. H. Cushing. A short history of the Downs stock of herring. ICES Journal of Marine Science, 49(4):437–443, November 1992. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/49/4/437/653339.

Cushing:1995:LTR

[Cus95a] D. H. Cushing. The long-term relationship between zooplankton and fish: IV. Spatial/temporal variablity and prediction. ICES Journal of Marine Science, 52(3–4):611–626, June 1995. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/52/3-4/611/841782.

Cushing:1995:QEB

[Cus95b] D. H. Cushing. Quantitative ecology and the brown trout. ICES Journal of Marine Science, 52(1):151–152, February 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/52/1/151/699765.

Corten:1996:VAS

[CvdK96] A. Corten and G. van de Kamp. Variation in the abundance of southern fish species in the southern North Sea in relation to hydrography and wind. ICES Journal of Marine Science, 53(6): 1113–1119, December 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/53/6/1113/592480.

Cotter:1997:TTD

[CWC^{+}97] A. J. R. Cotter, P. Walker, P. Coates, W. Cook, and P. J. Dare. Trial of a tractor dredger for cockles in Burry Inlet, South Wales. ICES Journal of Marine Science, 54(1):72–83, February 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/54/1/72/782192.

Chadwick:1993:R

[CWGC93] Michael Chadwick, M. J. Waldock, R. J. Gowen, and D. Calamari. Reviews. ICES Journal of Marine Science, 50(3): 325–327, ???? 1993. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/50/3/325/685321.

Chang:1998:ABL

[CZM98] S. Chang, V. S. Zdanowicz, and R. A. Murchelano. Associations between liver lesions in winter flounder (Pleuronectes americanus) and sediment chemical contaminants from northeast United States estuaries. ICES Journal of Marine Science, 55(5):954–969, October 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/55/5/954/691009.

Danielssen:1997:OVS

[Dan97] D. Danielssen. Oceanographic variability in the Skagerrak and Northern Kattegat, May–June, 1990. ICES Journal of Marine Science, 54(5):753–773, October 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/54/5/753/751469.

Duarte:1997:SGS

[DAP97] Rafael Duarte, Manuela Azevedo, and Pilar Pereda. Study of the growth of southern black and white monkfish stocks. ICES Journal of Marine Science, 54(5):866–874, October 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
54/5/866/751531.

Dawson:1991:OMM

[Daw91] Wendy A. Dawson. Otolith measurement as a method of identifying factors affecting first-year growth and stock separation of mackerel (Scomber scombrus L.). ICES Journal of Marine Science, 47(3):303–317, ???? 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/47/3/303/601170.

Dawe:1994:R

[Daw94] Earl G. Dawe. Review. ICES Journal of Marine Science, 51(1):131, ???? 1994. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/51/1/131/760293.

Dawson:1998:PES

[Daw98] L. H. J. Dawson. The physiological effects of salmon lice (Lepeophtheirus salmonis) infections on returning post-smolt sea trout (Salmo trutta L.) in western Ireland, 1996. ICES Journal of Marine Science, 55(2):193–200, April 1998. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
55/2/193/781710.

Davaine:1992:RBT

[DB92] P. Davaine and E. Beall. Relationships between temperature, population density, and growth in a seatrout population (S. trutta L.) of the Kerguelen Islands. ICES Journal of Marine Science, 49(4):445–451, November 1992. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/49/4/445/653343.

deBarros:1995:IGO

[dBH95] Pedro de Barros and Jens Christian Holst. Identification of geographic origin of Norwegian spring-spawning herring (Clupea harengus L.) based on measurements of scale anuli. ICES Journal of Marine Science, 52(5):863–872, October 1995. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
52/5/863/783886.

Davies:1998:TIN

[DBH98] Ian M. Davies, Susan K. Bailey, and Melanie J. C. Harding. Tributyltin inputs to the North Sea from shipping activities, and potential risk of biological effects. ICES Journal of Marine Science, 55(1):34–43, February 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/55/1/34/615943.

deBarros:1998:VNM

[dBT98] Pedro de Barros and Reidar Toresen. Variable natural mortality rate of juvenile Norwegian spring-spawning herring (Clupea harengus L.) in the Barents Sea. ICES Journal of Marine Science, 55(3):430–442, June 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/55/3/430/655054.

deBarros:1998:RCG

[dBTT98] Pedro de Barros, E. Mümtaz Tirasin, and Reidar Toresen. Relevance of cod (Gadus morhua L.) predation for inter-cohort variability in mortality of juvenile Norwegian spring-spawning herring (Clupea harengus L.). ICES Journal of Marine Science, 55(3):454–466, June 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/55/3/454/655067.

Dempson:1999:EAS

[Dem99] J. Dempson. Evaluation of an alternative strategy to enhance salmon populations: Cage rearing wild smolts from Conne River, Newfoundland. ICES Journal of Marine Science, 56 (4):422–432, August 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/56/4/422/624625.

deGroot:1996:PIM

[dG96a] S. J. de Groot. The physical impact of marine aggregate extraction in the North Sea. ICES Journal of Marine Science, 53(6): 1051–1053, December 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/53/6/1051/592468.

deGroot:1996:QAD

[dG96b] S. J. de Groot. Quantitative assessment of the development of the offshore oil and gas industry in the North Sea. ICES Journal of Marine Science, 53(6):1045–1050, December 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
53/6/1045/592465.

Dickson:1994:ALM

[Dic94] R. R. Dickson. Arthur Lee: 17 May 1920–28 December 1993. ICES Journal of Marine Science, 51(2):235–237, ???? 1994. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
51/2/235/674486.

DosSantos:1992:MDG

[DJ92] Jorge Dos Santos and Malcolm Jobling. A model to describe gastric evacuation in cod (Gadus morhua L.) fed natural prey. ICES Journal of Marine Science, 49(2):145–154, May 1992. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
49/2/145/683865.

Dahl:1998:TSV

[DJ98] Einar Dahl and Tore Johannessen. Temporal and spatial variability of phytoplankton and chlorophyll a: lessons from the south coast of Norway and the Skagerrak. ICES Journal of Marine Science, 55(4):680–687, August 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/55/4/680/698155.

Duncan:1996:PLF

[DK96] Annie Duncan and Jan Kubecka. Patchiness of longitudinal fish distributions in a river as revealed by a continuous hydroacoustic survey. ICES Journal of Marine Science, 53(2):161–165, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/53/2/161/791613.

Drevs:1999:GVA

[DKLM99] T. Drevs, V. Kadakas, T. Lang, and S. Mellergaard. Geographical variation in the age/length relationship in Baltic flounder (Platichthys flesus). ICES Journal of Marine Science, 56 (2):134–137, April 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/56/2/134/622433.

DeCasabianca:1997:SCN

[DLMS97] M.-L. De Casabianca, T. Laugier, and E. Marinho-Soriano. Seasonal changes of nutrients in water and sediment in a Mediterranean lagoon with shellfish farming activity (Thau Lagoon, France). ICES Journal of Marine Science, 54(5): 905–916, October 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/54/5/905/751552.

Daan:1996:STL

[DM96] Rogier Daan and Maarten Mulder. On the short-term and long-term impact of drilling activities in the Dutch sector of the North Sea. ICES Journal of Marine Science, 53(6):1036– 1044, December 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/53/6/1036/592463.

Darcy:1999:APA

[DM99] G. H. Darcy and G. C. Matlock. Application of the precautionary approach in the national standard guidelines for conservation and management of fisheries in the United States. ICES Journal of Marine Science, 56(6):853–859, December 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
56/6/853/658132.

Danielsdottir:1997:GSW

[DMAG97] A. K. Danielsdottir, G. Marteinsdottir, F. Arnason, and S. Gudjonsson. Genetic structure of wild and reared Atlantic salmon (Salmo salar L.) populations in Iceland. ICES Journal of Marine Science, 54(6):986–997, December 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
54/6/986/674084.

Dunn:1993:LNM

[DMUR93] J. Dunn, R. B. Mitchell, G. G. Urquhart, and B. J. Ritchie. LOCHNESS — a new multi-net midwater sampler. ICES Journal of Marine Science, 50(2):203–212, ???? 1993. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
50/2/203/774823.

Dawson:1997:CSS

[DPHM97] L. H. J. Dawson, A. W. Pike, D. F. Houlihan, and A. H. McVicar. Comparison of the susceptibility of sea trout (Salmo trutta L.) and Atlantic salmon (Salmo salar L.) to sea lice (Lepeophtheirus salmonis (Krøyer, 1837)) infections. ICES Journal of Marine Science, 54(6):1129–1139, December 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
54/6/1129/673976.

dePuelles:1996:DVV

[dPVV^{+}96] M. L. Fernández de Puelles, L. Valdés, M. Varela, M. T. Alvarez-Ossorio, and N. Halliday. Diel variations in the vertical distribution of copepods off the north coast of Spain. ICES Journal of Marine Science, 53(1):97–106, February 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
53/1/97/731595.

Daan:1996:I

[DRP96] Niels Daan, Katherine Richardson, and John G. Pope. Introduction. ICES Journal of Marine Science, 53(6):879–883, December 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/53/6/879/592513.

Dommasnes:1994:ROC

[DRR94] Are Dommasnes, Francisco Rey, and Ingolf Røttingen. Reduced oxygen concentrations in herring wintering areas. ICES Journal of Marine Science, 51(1):63–69, ???? 1994. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/51/1/63/760303.

dosSantos:1995:TFC

[dSJ95] Jorge dos Santos and Malcolm Jobling. Test of a food consumption model for the Atlantic cod. ICES Journal of Marine Science, 52(2):209–219, April 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/52/2/209/593899.

Danielssen:1996:LTH

[DSO96] Didrik S. Danielssen, Einar Svendsen, and Marek Ostrowski. Long-term hydrographic variation in the Skagerrak based on the section torungen–hirtshals. ICES Journal of Marine Science, 53(6):917–925, December 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/53/6/917/592524.

deVries:1998:PTN

[dVDP^{+}98] I. de Vries, R. N. M. Duin, J. C. H. Peeters, F. J. Los, M. Bokhorst, and R. W. P. M. Laane. Patterns and trends in nutrients and phytoplankton in Dutch coastal waters: comparison of time-series analysis, ecological model simulation, and mesocosm experiments. ICES Journal of Marine Science, 55 (4):620–634, August 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/55/4/620/698143.

Dethlefsen:1996:MNS

[DvWC96] Volkert Dethlefsen, Hein von Westernhagen, and Patricia Cameron. Malformations in North Sea pelagic fish embryos during the period 1984–1995. ICES Journal of Marine Science, 53(6):1024–1035, December 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/53/6/1024/592462.

DiNardo:1999:AUD

[DW99] G. T. DiNardo and J. A. Wetherall. Accounting for uncertainty in the development of harvest strategies for the Northwestern Hawaiian Islands lobster trap fishery. ICES Journal of Marine Science, 56(6):943–951, December 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/56/6/943/658151.

Economou:1991:FFE

[Eco91] Alcibiades N. Economou. Food and feeding ecology of five gadoid larvae in the northern North Sea. ICES Journal of Marine Science, 47(3):339–351, ???? 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/47/3/339/601176.

Erzini:1996:SSS

[EGB^{+}96] Karim Erzini, Jorge M. S. Gonçalves, Luis Bentes, Pedro G. Lino, and João Cruz. Species and size selectivity in a Portuguese multispecies artisanal long-line fishery. ICES Journal of Marine Science, 53(5):811–819, October 1996. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
53/5/811/704315.

Evans:1994:CFC

[EHEW94] S. M. Evans, J. E. Hunter, Elizal, and R. I. Wahju. Composition and fate of the catch and bycatch in the Farne Deep (North Sea) Nephrops fishery. ICES Journal of Marine Science, 51 (2):155–168, ???? 1994. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/51/2/155/674474.

Egidius:1991:MIW

[EHJN91] Emmy Egidius, Lars P. Hansen, Bror Jonsson, and Gunnar Nævdal. Mutual impact of wild and cultured Atlantic salmon in Norway. ICES Journal of Marine Science, 47(3):404–410, ???? 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/47/3/404/601194.

Ehrich:1991:CFE

[Ehr91] Siegfried Ehrich. Comparative fishing experiments by research trawlers for cod and haddock in the North Sea. ICES Journal of Marine Science, 47(3):275–283, ???? 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/47/3/275/601164.

Engaas:1998:SST

[EJW98] A. Engås, T. Jørgensen, and C. W. West. A species-selective trawl for demersal gadoid fisheries. ICES Journal of Marine Science, 55(5):835–845, October 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/55/5/835/690931.

Estep:1994:SSH

[EMN^{+}94] Kenneth W. Estep, Ferren MacIntyre, Thomad T. Noji, Boonchai Stensholt, and Torger Øritsland. Seal sizes and habitat conditions assessed from aerial photography and video analysis. ICES Journal of Marine Science, 51(3):253–261, ???? 1994. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
51/3/253/620179.

Ellis:1997:SPP

[EN97] Tim Ellis and Richard D. M. Nash. Spawning of plaice Pleuronectes platessa L. around the Isle of Man, Irish Sea. ICES Journal of Marine Science, 54(1):84–92, February 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
54/1/84/782197.

Engaas:1992:DVB

[ES92] Arill Engås and Aud Void Soldal. Diurnal variations in bottom trawl catch rates of cod and haddock and their influence on abundance indices. ICES Journal of Marine Science, 49 (1):89–95, February 1992. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/49/1/89/732790.

Ehrenberg:1996:ADB

[ET96] John E. Ehrenberg and Thomas C. Torkelson. Application of dual-beam and split-beam target tracking in fisheries acoustics. ICES Journal of Marine Science, 53(2):329–334, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
53/2/329/792613.

Everts:1993:BEQ

[Eve93] J. Everts. Biomagnification and environmental quality criteria: a physiological approach. ICES Journal of Marine Science, 50 (4):333–335, ???? 1993. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/50/4/333/668217.

Fogarty:1997:MCP

[FA97] Michael J. Fogarty and Julian T. Addison. Modelling capture processes in individual traps: entry, escapement and soak time. ICES Journal of Marine Science, 54(2):193–205, April 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
54/2/193/620540.

Fossen:1999:BCI

[FAN99] I. Fossen, O. T. Albert, and E. M. Nilssen. Back-calculated individual growth of long rough dab (Hippoglossoides platessoides) in the Barents Sea. ICES Journal of Marine Science, 56(5):689–696, October 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/56/5/689/691326.

Farebrother:1992:SST

[Far92] R. W. Farebrother. A simulation study of three iterated weighted least-squares estimators of mortality rates from single-release tagging studies. ICES Journal of Marine Science, 49(2):185–190, May 1992. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/49/2/185/683890.

Fifas:1999:EMS

[FB99] S. Fifas and P. Berthou. An efficiency model of a scallop (Pecten maximus, L.) experimental dredge: Sensitivity study.

ICES Journal of Marine Science, 56(4):489–499, August 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
56/4/489/624647.

Frid:1996:VSB

[FBG96] C. L. J. Frid, J. B. Buchanan, and P. R. Garwood. Variability and stability in benthos: twenty-two years of monitoring off Northumberland. ICES Journal of Marine Science, 53(6): 978–980, December 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/53/6/978/592541.

Furness:1997:SMM

[FC97] R. W. Furness and Kees (C. J.) Camphuysen. Seabirds as monitors of the marine environment. ICES Journal of Marine Science, 54(4):726–737, August 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/54/4/726/607516.

Fleming:1997:ETG

[FE97] I. A. Fleming and S. Einum. Experimental tests of genetic divergence of farmed from wild Atlantic salmon due to domestication. ICES Journal of Marine Science, 54(6):1051–1063, December 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/54/6/1051/673920.

Fiksen:1995:VDP

[FG95a] Øyvind Fiksen and Jarl Giske. Vertical distribution and population dynamics of copepods by dynamic optimization. ICES Journal of Marine Science, 52(3–4):483–503, June 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/52/3-4/483/841611.

Fransz:1995:POS

[FG95b] H. G. Fransz and S. R. Gonzalez. The production of Oithona similis (Copepoda: Cyclopoida) in the Southern Ocean. ICES Journal of Marine Science, 52(3–4):549–555, June 1995. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/52/3-4/549/841684.

Fjosne:1996:DCP

[FG96] Karen Fjøsne and Jakob Gjøsæter. Dietary composition and the potential of food competition between 0-group cod (Gadus morhua L.) and some other fish species in the littoral zone. ICES Journal of Marine Science, 53(5):757–770, October 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
53/5/757/704274.

Freon:1996:DVS

[FGS96] P. Fréon, F. Gerlotto, and M. Soria. Diel variability of school structure with special reference to transition periods. ICES Journal of Marine Science, 53(2):459–464, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
53/2/459/792869.

Frid:1996:FFC

[FH96] Christopher L. J. Frid and Niette V. Huliselan. Far-field control of long-term changes in Northumberland (NW North Sea) coastal zooplankton. ICES Journal of Marine Science, 53(6): 972–977, December 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/53/6/972/592539.

Field:1998:DSP

[FHA^{+}98] R. H. Field, J. M. Hills, R. J. A. Atkinson, S. Magill, and A. M. Shanks. Distribution and seasonal prevalence of Hematodinium sp. infection of the Norway lobster (Nephrops norvegicus) around the west coast of Scotland. ICES Journal of Marine Science, 55(5):846–858, October 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/55/5/846/690933.

Findlay:1998:PSC

[FL98] Malcolm Findlay and Jon Lansley. Plankton sample colour as a determinant of the abundance of stage I eggs in the western mackerel (Scomber scombrus L.) stock. ICES Journal of Marine Science, 55(1):141–144, February 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/55/1/141/615940.

Fryer:1993:PCM

[FN93] R. J. Fryer and M. D. Nicholson. The power of a contaminant monitoring programme to detect linear trends and incidents. ICES Journal of Marine Science, 50(2):161–168, ???? 1993. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
50/2/161/774799.

Fryer:1999:USC

[FN99] R. J. Fryer and M. D. Nicholson. Using smoothers for comprehensive assessments of contaminant time series in marine biota. ICES Journal of Marine Science, 56(5):779–790, October 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
56/5/779/691346.

Fogarty:1993:RRV

[Fog93] Michael J. Fogarty. Recruitment in randomly varying environments. ICES Journal of Marine Science, 50(3):247–260, ???? 1993. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/50/3/247/685300.

Folkvord:1994:GPT

[FØK94] Arild Folkvord, Victor Øiestad, and Per Gunnar Kvenseth. Growth patterns of three cohorts of Atlantic cod larvae (Gadus morhua L.) studied in a macrocosm. ICES Journal of Marine Science, 51(3):325–336, ???? 1994. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/51/3/325/620225.

Foote:1990:DIT

[Foo90a] Kenneth G. Foote. Designing an improved transducer array geometry. ICES Journal of Marine Science, 46(2):129–132, ???? 1990. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/46/2/129/645336.

Foote:1990:SCE

[Foo90b] Kenneth G. Foote. Spheres for calibrating an eleven-frequency acoustic measurement system. ICES Journal of Marine Science, 46(3):284–286, ???? 1990. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/46/3/284/596405.

Foote:1991:SMD

[Foo91] Kenneth G. Foote. Summary of methods for determining fish target strength at ultrasonic frequencies. ICES Journal of Marine Science, 48(2):211–217, August 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/48/2/211/644496.

Foote:1995:CEA

[Foo95] Kenneth G. Foote. Comparison of equal-area cylindrical and circular piston transducers. ICES Journal of Marine Science, 52(1):67–75, February 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/52/1/67/699798.

Foote:1999:ECS

[Foo99] K. G. Foote. Extinction cross-section of Norwegian springspawning herring. ICES Journal of Marine Science, 56(5): 606–612, October 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/56/5/606/691310.

Fossum:1992:RBS

[Fos92] P. Fossum. The recovery of the Barents Sea capelin (Mallotus villosus) from a larval point of view. ICES Journal of Marine Science, 49(2):237–243, May 1992. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/49/2/237/683931.

Fossum:1996:SFF

[Fos96] Petter Fossum. A study of first-feeding herring (Clupea harengus L.) larvae during the period 1985–1993. ICES Journal of Marine Science, 53(1):51–59, February 1996. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
53/1/51/731580.

Fowler:1999:MMS

[Fow99] C. W. Fowler. Management of multi-species fisheries: from overfishing to sustainability. ICES Journal of Marine Science, 56(6):927–932, December 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/56/6/927/658147.

Falk-Petersen:1990:OOS

[FPHM90] I.-B. Falk-Petersen, T. Haug, and E. Moksness. Observations on the occurrence, size, and feeding of pelagic larvae of the common wolffish (Anarhichas lupus) in western Finnmark, northern Norway. ICES Journal of Marine Science, 46(2):148–154, ???? 1990. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/46/2/148/645350.

Floch:1991:JBA

[FPW91] J. Y. Floc’h, R. Pajot, and I. Wallentinus. The Japanese brown alga Undaria pinnatifida on the coast of France and its possible establishment in European waters. ICES Journal of Marine Science, 47(3):379–390, ???? 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/47/3/379/601185.

Francis:1993:LE

[Fra93] R. I. C. C. Francis. Letters to the Editor. ICES Journal of Marine Science, 50(4):505, ???? 1993. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/50/4/505/668269.

Franks:1997:NME

[Fra97] Peter J. S. Franks. New models for the exploration of biological processes at fronts. ICES Journal of Marine Science, 54 (2):161–167, April 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/54/2/161/620532.

Friedland:1993:MSN

[FRK93] K. D. Friedland, D. G. Reddin, and J. F. Kocik. Marine survival of North American and European Atlantic salmon: effects of growth and environment. ICES Journal of Marine Science, 50 (4):481–492, ???? 1993. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/50/4/481/668258.

Fryer:1991:MBH

[Fry91] R. J. Fryer. A model of between-haul variation in selectivity. ICES Journal of Marine Science, 48(3):281–290, November 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/48/3/281/737244.

Foote:1993:DPE

[FS93] K. G. Foote and G. Stefánsson. Definition of the problem of estimating fish abundance over an area from acoustic linetransect measurements of density. ICES Journal of Marine Science, 50(4):369–381, ???? 1993. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/50/4/369/668228.

Foyn:1997:DSR

[FS97] Lars Føyn and Ingrid Sværen. Distribution and sedimentation of radionuclides in the Barents Sea. ICES Journal of Marine Science, 54(3):333–340, June 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/54/3/333/652462.

Flinkman:1994:CCE

[FVC94] J. Flinkman, I. Vuorinen, and M. Christiansen. Calanoid copepod eggs survive passage through fish digestive tracts. ICES Journal of Marine Science, 51(1):127–129, ???? 1994. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
51/1/127/760290.

Gjosaeter:1994:GPC

[GA94] Harald Gjøsæter and Adnan M. Ajiad. Growth of polar cod, Boreogadus saida (Lepechin), in the Barents Sea. ICES Journal of Marine Science, 51(1):115–120, ???? 1994. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
51/1/115/760285.

Gislason:1998:VPS

[GA98] Astthor Gislason and Olafur S. Astthorsson. Variability in the population structure of Calanus finmarchicus in Icelandic waters in spring. ICES Journal of Marine Science, 55(4):811–813, August 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/55/4/811/698249.

Garthe:1997:IHF

[Gar97] Stefan Garthe. Influence of hydrography, fishing activity, and colony location on summer seabird distribution in the southeastern North Sea. ICES Journal of Marine Science, 54(4): 566–577, August 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/54/4/566/607446.

Gauld:1990:MLS

[Gau90] James A. Gauld. Movements of lesser sandeels (Ammodytes marinus Raitt) tagged in the northwestern North Sea. ICES Journal of Marine Science, 46(3):229–231, ???? 1990. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
46/3/229/596390.

Gayoso:1998:LTP

[Gay98] Ana María Gayoso. Long-term phytoplankton studies in the Bahía Blanca estuary, Argentina. ICES Journal of Marine Science, 55(4):655–660, August 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/55/4/655/698149.

Gjoen:1997:PPF

[GB97] H. M. Gjøen and H. B. Bentsen. Past, present, and future of genetic improvement in salmon aquaculture. ICES Journal of Marine Science, 54(6):1009–1014, December 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
54/6/1009/673889.

Greenstreet:1997:SVC

[GBB^{+}97] Simon P. R. Greenstreet, Andrew D. Bryant, Niall Broekhuizen, Stephen J. Hall, and Mike R. Heath. Seasonal variation in the consumption of food by fish in the North Sea and implications for food web dynamics. ICES Journal of Marine Science, 54 (2):243–266, April 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/54/2/243/620553.

Gonzalez:1996:AGS

[GCG96] Angel F. González, Bernardino G. Castro, and Angel Guerra. Age and growth of the short-finned squid Illex coindetii in Galician waters (NW Spain) based on statolith analysis. ICES Journal of Marine Science, 53(5):802–810, October 1996. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
53/5/802/704311.

Gaspar:1999:ETS

[GCM99] M. B. Gaspar, M. Castro, and C. C. Monteiro. Effect of tooth spacing and mesh size on the catch of the Portuguese clam and razor clam dredge. ICES Journal of Marine Science, 56(1): 103–110, February 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/56/1/103/652185.

Griffiths:1996:CAB

[GD96] Gwyn Griffiths and José I. Diaz. Comparison of acoustic backscatter measurements from a ship-mounted Acoustic Doppler Current Profiler and an EK500 scientific echo-sounder. ICES Journal of Marine Science, 53(2):487–491, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
53/2/487/792929.

Geromont:1999:DAM

[GDJC99] H. F. Geromont, J. A. A. De Oliveira, S. J. Johnston, and C. L. Cunningham. Development and application of management procedures for fisheries in southern Africa. ICES Journal of Marine Science, 56(6):952–966, December 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
56/6/952/658153.

Gundersen:1998:CBA

[GG98] Agnes C. Gundersen and Harald Gjøæster. A comparison between abundance estimates of the Barents Sea capelin (Mallotus villosus Müller) at the larval, 0-group and 1-group stage, for the year classes 1981–1994. ICES Journal of Marine Science, 55 (1):95–101, February 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/55/1/95/615954.

Gonzalez-Gurriaran:1994:SMV

[GGF94] Eduardo González-Gurriarán and Juan Freire. Sexual maturity in the velvet swimming crab Necora puber (Brachyura, Portunidae): morphometric and reproductive analyses. ICES Journal of Marine Science, 51(2):133–145, ???? 1994. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
51/2/133/674470.

Gonzalez-Gurriaran:1998:GMI

[GGFFF98] E. González-Gurriarán, J. Freire, A. C. Fariña, and A. Fernández. Growth at moult and intermoult period in the Norway lobster Nephrops norvegicus from Galician waters. ICES Journal of Marine Science, 55(5):924–940, October 1998. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
55/5/924/691002.

Grizel:1991:IFJ

[GH91] Henri Grizel and Maurice Héral. Introduction into France of the Japanese oyster (Crassostrea gigas). ICES Journal of Marine Science, 47(3):399–403, ???? 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/47/3/399/601192.

Gaard:1998:PVF

[GHH98] Eilif Gaard, Bogi Hansen, and Signar P. Heinesen. Phytoplankton variability on the Faroe Shelf. ICES Journal of Marine Science, 55(4):688–696, August 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/55/4/688/698157.

Godo:1997:BDB

[GHM97] Olav Rune Godø, Irene Huse, and Kathrine Michalsen. Bait defence behaviour of wolffish and its impact on long-line catch rates. ICES Journal of Marine Science, 54(2):273–275, April 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/54/2/273/620556.

Gislason:1999:SMR

[Gis99] H. Gislason. Single and multispecies reference points for Baltic fish stocks. ICES Journal of Marine Science, 56(5):

571–583, October 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/56/5/571/691305.

Grioche:1997:PSI

[GK97] A. Grioche and P. Koubbi. A preliminary study of the influence of a coastal frontal structure on ichthyoplankton assemblages in the English Channel. ICES Journal of Marine Science, 54 (1):93–104, February 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/54/1/93/782199.

Godo:1998:ETS

[GKT98] Olav Rune Godø, William A. Karp, and Atle Totland. Effects of trawl sampling variability on precision of acoustic abundance estimates of gadoids from the Barents Sea and the Gulf of Alaska. ICES Journal of Marine Science, 55(1):86–94, February 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/55/1/86/615952.

Gledhill:1996:EVA

[Gle96] C. Gledhill. Evaluation of video and acoustic index methods for assessing reef-fish populations. ICES Journal of Marine Science, 53(2):483–485, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/53/2/483/792920.

Greenstreet:1998:SVI

[GMA98] Simon P. R. Greenstreet, Julie A. McMillan, and Eric Armstrong. Seasonal variation in the importance of pelagic fish in the diet of piscivorous fish in the Moray Firth, NE Scotland: a response to variation in prey abundance? ICES Journal of Marine Science, 55(1):121–133, February 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/55/1/121/615933.

Gonzalez:1998:VSA

[GMG98] Angel F. González, William K. Macy III, and Angel Guerra. Validation of a semi-automatic image analysis system to age squids and its application to age Illex coindetii statoliths. ICES Journal of Marine Science, 55(3):535–544, June 1998.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
55/3/535/655089.

Godo:1995:TTE

[God95] Olav Rune Godø. Transplantation-tagging-experiments in preliminary studies of migration of cod off Norway. ICES Journal of Marine Science, 52(6):955–962, December 1995. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
52/6/955/649274.

Gislason:1998:MSE

[GR98a] H. Gislason and J. Rice. Mini-Symposium on Ecosystem Effects of Fishing, 1996. ICES Journal of Marine Science, 55(3):329, June 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/55/3/329/655005.

Gislason:1998:MRS

[GR98b] H. Gislason and J. Rice. Modelling the response of size and diversity spectra of fish assemblages to changes in exploitation. ICES Journal of Marine Science, 55(3):362–370, June 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
55/3/362/655029.

Gowen:1998:PDR

[GRDCW98] R. J. Gowen, R. Raine, M. Dickey-Collas, and M. White. Plankton distributions in relation to physical oceanographic features on the southern Malin Shelf, August 1996. ICES Journal of Marine Science, 55(6):1095–1111, December 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
55/6/1095/638984.

Greve:1994:GBI

[Gre94] Wulf Greve. The 1989 German Bight invasion of Muggiaea atlantica. ICES Journal of Marine Science, 51(4):355–358, August 1994. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/51/4/355/714239.

Greve:1995:MPC

[Gre95] Wulf Greve. Mutual predation causes bifurcations in pelagic ecosystems: The simulation model PLITCH (PLanktonic swITCH ), experimental tests, and theory. ICES Journal of Marine Science, 52(3–4):505–510, June 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/52/3-
4/505/841623.

Greenstreet:1997:AAS

[Gre97a] S. Greenstreet. An assessment of the acoustic survey technique, RoxAnn, as a means of mapping seabed habitat. ICES Journal of Marine Science, 54(5):939–959, October 1997. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
54/5/939/751574.

Gremillet:1997:CPU

[Gré97b] David Grémillet. Catch per unit effort, foraging efficiency, and parental investment in breeding great cormorants (Phalacrocorax carbo carbo). ICES Journal of Marine Science, 54(4): 635–644, August 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/54/4/635/607474.

Greve:1996:BCZ

[GRN96] Wulf Greve, Frank Reiners, and Jutta Nast. Biocoenotic changes of the zooplankton in the German Bight: the possible effects of eutrophication and climate. ICES Journal of Marine Science, 53(6):951–956, December 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/53/6/951/592535.

Grygiel:1999:SSP

[Gry99] W. Grygiel. Synoptic survey of pathological symptoms in herring (Clupea harengus) and sprat (Sprattus sprattus) in the Baltic Sea. ICES Journal of Marine Science, 56(2):169–174, April 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/56/2/169/622461.

Goodson:1996:SCH

[GS96] A. D. Goodson and C. R. Sturtivant. Sonar characteristics of the harbour porpoise (Phocoena phocoena): source levels and spectrum. ICES Journal of Marine Science, 53(2):465–472, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/53/2/465/792885.

Gudmundsson:1995:TSA

[Gud95] Gudmundur Gudmundsson. Time series analysis of catchat-length data. ICES Journal of Marine Science, 52(5):781– 795, October 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/52/5/781/783834.

Gudmundsson:1998:LTV

[Gud98] Kristinn Gudmundsson. Long-term variation in phytoplankton productivity during spring in Icelandic waters. ICES Journal of Marine Science, 55(4):635–643, August 1998. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
55/4/635/698145.

Guichet:1995:DEH

[Gui95] R. Guichet. The diet of European hake (Merluccius merluccius) in the northern part of the Bay of Biscay. ICES Journal of Marine Science, 52(1):21–31, February 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/52/1/21/699771.

Godo:1993:MCA

[GW93] Olav Rune Godø and Vidar G. Wespestad. Monitoring changes in abundance of gadoids with varying availability to trawl and acoustic surveys. ICES Journal of Marine Science, 50(1):39–51, ???? 1993. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/50/1/39/652881.

Godo:1999:IDD

[GWE99] O. R. Godø, S. J. Walsh, and A. Engås. Investigating density-dependent catchability in bottom-trawl surveys. ICES Journal of Marine Science, 56(3):292–298, June 1999. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
56/3/292/623503.

Heip:1992:PMSb

[HA92a] C. Heip and W. Arntz. Paper from the Mini-Symposium on the Benthic Ecology of the North Sea held in Copenhagen on 9 October 1990. ICES Journal of Marine Science, 49(2):125, May 1992. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/49/2/125/683849.

Heip:1992:PMSa

[HA92b] C. Heip and W. Arntz. Papers from the Mini-Symposium on the Benthic Ecology of the North Sea held in Copenhagen on 9 October 1990. ICES Journal of Marine Science, 49(1):3, February 1992. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/49/1/3/732778.

Hamano:1996:TSM

[Ham96a] A. Hamano. Target-strength measurements of Pacific hake (Merluccius productus) in Canadian waters using quasi-ideal and conventional beam transducers. ICES Journal of Marine Science, 53(2):273–279, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/53/2/273/792515.

Hampton:1996:AEP

[Ham96b] Ian Hampton. Acoustic and egg-production estimates of South African anchovy biomass over a decade: comparisons, accuracy, and utility. ICES Journal of Marine Science, 53(2):493–500, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/53/2/493/792943.

Hansson:1996:PRN

[Han96] S. Hansson. Predation rates by North Sea cod (Gadus morhua) — predictions from models on gastric evacuation and bioenergetics. ICES Journal of Marine Science, 53(1):107–114, February 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/53/1/107/731538.

Harris:1995:MDJ

[Har95] Roger Harris. In memory of Dr John Gamble. ICES Journal of Marine Science, 52(3–4):263, June 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https:/

/academic.oup.com/icesjms/article/52/3-4/263/841456.

Harms:1997:WMT

[Har97] Ingo H. Harms. Water mass transformation in the Barents Sea — application of the Hamburg Shelf Ocean Model (HamSOM). ICES Journal of Marine Science, 54(3):351–365, June 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
54/3/351/652468.

Hastie:1996:ETC

[Has96] Lee C. Hastie. Estimation of trawl codend selectivity for squid (Loligo forbesi), based on Scottish research vessel survey data. ICES Journal of Marine Science, 53(4):741–744, August 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
53/4/741/799763.

Hatfield:1996:TRM

[Hat96] Emma M. C. Hatfield. Towards resolving multiple recruitment into loliginid fisheries: Loligo gahi in the Falkland Islands fishery. ICES Journal of Marine Science, 53(3):565–575, June 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/53/3/565/625705.

Hay:1995:EPS

[Hay95] Steve Hay. Egg production and secondary production of common North sea copepods: Field estimates with regional and seasonal comparisons. ICES Journal of Marine Science, 52 (3–4):315–327, June 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.oup.

com/icesjms/article/52/3-4/315/841498.

Hayes:1998:ERA

[Hay98] K. R. Hayes. Ecological risk assessment for ballast water introductions: a suggested approach. ICES Journal of Marine Science, 55(2):201–212, April 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/55/2/201/781716.

Hoines:1998:STS

[HBA98] Å. S. Høines, O. A. Bergstad, and O. T. Albert. The structure and temporal stability of the fish community on a coastal bank utilized as a spawning ground by herring. ICES Journal of Marine Science, 55(2):271–288, April 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/55/2/271/781749.

Heip:1992:TBD

[HBC^{+}92] C. Heip, D. Basford, J. A. Craeymeersch, J. M. Dewarumez, J. Dorjes, P. de Wilde, G. Duineveld, A. Eleftheriou, P. M. J. Herman, U. Niermann, P. Kingston, A. Kunitzer, E. Rachor, H. Rumohr, K. Soetaert, and T. Soltwedel. Trends in biomass, density and diversity of North Sea macrofauna. ICES Journal of Marine Science, 49(1):13–22, February 1992. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
49/1/13/732769.

Hamerlynck:1993:DRJ

[HCA93] O. Hamerlynck, A. Cattrijsse, and R. V. Arellano. Daily ration of juvenile Pomatoschistus lozanoi de Buen (Pisces: Gobiidae). ICES Journal of Marine Science, 50(4):471–480, ???? 1993. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
50/4/471/668253.

Howell:1991:FRN

[HCH91] B. R. Howell, A. R. Child, and R. G. Houghton. Fertilization rate in a natural population of the common sole (Solea solea L.). ICES Journal of Marine Science, 48(1):53–59, May 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
48/1/53/696118.

Haakanson:1994:LEC

[HCR94] J. L. Håkanson, S. H. Coombs, and P. Ré. Lipid and elemental composition of sprat (Sprattus sprattus) larvae at mixed and stratified sites in the German Bight of the North Sea. ICES Journal of Marine Science, 51(2):147–154, ???? 1994.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
51/2/147/674472.

Heath:1990:AMF

[HD90] M. Heath and J. Dunn. Avoidance of a midwater frame trawl by herring larvae. ICES Journal of Marine Science, 47(2):140–147, ???? 1990. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/47/2/140/611305.

Heessen:1996:LTT

[HD96a] Henk J. L. Heessen and Niels Daan. Long-term trends in ten non-target North Sea fish species. ICES Journal of Marine Science, 53(6):1063–1078, December 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/53/6/1063/592471.

Hewitt:1996:LTS

[HD96b] Roger P. Hewitt and David A. Demer. Lateral target strength of Antarctic krill. ICES Journal of Marine Science, 53(2): 297–302, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/53/2/297/792552.

Heath:1995:SSD

[Hea95] Michael R. Heath. Size spectrum dynamics and the planktonic ecosystem of loch linnhe. ICES Journal of Marine Science, 52 (3–4):627–642, June 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.oup.

com/icesjms/article/52/3-4/627/841842.

Heessen:1996:TSD

[Hee96] Henk J. L. Heessen. Time-series data for a selection of forty fish species caught during the international bottom trawl survey. ICES Journal of Marine Science, 53(6):1079–1084, December 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/53/6/1079/592473.

Hempel:1995:ABJ

[Hem95] Gotthilf Hempel. Adolf Bückmann (17 January 1900–7 December 1993). ICES Journal of Marine Science, 52(2):257–259, April 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/52/2/257/593916.

Horstman:1994:IDA

[HF94] Katherine R. Horstman and Julie M. Fives. Ichthyoplankton distribution and abundance in the Celtic Sea. ICES Journal of Marine Science, 51(4):447–460, August 1994. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
51/4/447/714257.

Hall:1996:EMC

[HFP96] John A. Hall, Christopher L. J. Frid, and Roger K. Proudfoot. Effects of metal contamination on macrobenthos of two North Sea estuaries. ICES Journal of Marine Science, 53(6): 1014–1023, December 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/53/6/1014/592459.

Haralabous:1996:ANN

[HG96] J. Haralabous and S. Georgakarakos. Artificial neural networks as a tool for species identification of fish schools. ICES Journal of Marine Science, 53(2):173–180, April 1996. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
53/2/173/792280.

Hop:1992:SFE

[HGD92] Haakon Hop, Jakob Gjøsæter, and Didrik S. Danielssen. Seasonal feeding ecology of cod (Gadus morhua L.) on the Norwegian Skagerrak coast. ICES Journal of Marine Science, 49(4): 453–461, November 1992. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/49/4/453/653346.

Houde:1994:DMI

[HGDC94] E. D. Houde, J. C. Gamble, S. E. Dorsey, and J. H. Cowan, Jr. Drifting mesocosms: the influence of gelatinous zooplankton on mortality of bay anchovy, Anchoa mitchilli, eggs and yolksac larvae. ICES Journal of Marine Science, 51(4):383–394, August 1994. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/51/4/383/714245.

Haug:1995:DFA

[HGLN95] Tore Haug, Harald Gjøsæter, Ulf Lindstrøm, and Kjell T. Nilssen. Diet and food availability for north-east Atlantic minke whales (Balaenoptera acutorostrata), during the summer of 1992. ICES Journal of Marine Science, 52(1):77–86, February 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/52/1/77/699805.

Hedgepeth:1999:EMS

[HGOT99] J. B. Hedgepeth, V. F. Gallucci, F. O’Sullivan, and R. E. Thorne. An expectation maximization and smoothing approach for indirect acoustic estimation of fish size and density. ICES Journal of Marine Science, 56(1):36–50, February 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
56/1/36/652198.

Hamerlynck:1993:GFP

[HH93] Olivier Hamerlynck and Kris Hostens. Growth, feeding, production, and consumption in 0-group bib (Trisopterus luscus L.) and whiting (Merlangius merlangus L.) in a shallow coastal area of the south-west Netherlands. ICES Journal of Marine Science, 50(1):81–91, ???? 1993. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/50/1/81/652891.

Hancock:1995:SBC

[HHA95] Jim Hancock, Paul J. B. Hart, and Tarcisio Antezana. Searching behaviour and catch of horse mackerel (Trachurus murphyi) by industrial purse-seiners off south-central Chile. ICES Journal of Marine Science, 52(6):991–1004, December 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
52/6/991/649287.

Heessen:1996:INS

[HHB96] Henk J. L. Heessen, John R. G. Hislop, and Trevor W. Boon. An invasion of the North Sea by blue-mouth, Helicolenus dactylopterus (Pisces, Scorpaenidae). ICES Journal of Marine Science, 53(5):874–877, October 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/53/5/874/704344.

Huys:1992:MNS

[HHHS92] R. Huys, P. M. J. Herman, C. H. R. Heip, and K. Soetaert. The meiobenthos of the North Sea: density, biomass trends and distribution of copepod communities. ICES Journal of Marine Science, 49(1):23–44, February 1992. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/49/1/23/732774.

Hickel:1998:TVM

[Hic98]  Wolfgang Hickel.  Temporal variability of micro- and

nanoplankton in the German Bight in relation to hydrographic structure and nutrient changes. ICES Journal of Marine Science, 55(4):600–609, August 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/55/4/600/698139.

Hilden:1990:R

[Hil90] Mikael Hildén. Reviews. ICES Journal of Marine Science, 46 (3):340–341, ???? 1990. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/46/3/340/596424.

Hiilivirta:1998:CTM

[HIL98] P. Hiilivirta, E. Ikonen, and J. Lappalainen. Comparison of two methods for distinguishing wild from hatchery-reared salmon (Salmo salar Linnaeus, 1758) in the Baltic Sea. ICES Journal of Marine Science, 55(6):981–986, December 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
55/6/981/639008.

Hislop:1996:CNS

[His96] John R. G. Hislop. Changes in North Sea gadoid stocks. ICES Journal of Marine Science, 53(6):1146–1156, December 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
53/6/1146/592490.

Horne:1999:MFE

[HJ99] J. K. Horne and J. M. Jech. Multi-frequency estimates of fish abundance: constraints of rather high frequencies. ICES Journal of Marine Science, 56(2):184–199, April 1999. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
56/2/184/622475.

Horwood:1990:FBC

[HJB90] J. W. Horwood, O. L. R. Jacobs, and D. J. Ballance. A feedback control law to stabilize fisheries. ICES Journal of Marine Science, 47(1):57–64, ???? 1990. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/47/1/57/638528.

Hansen:1999:IEF

[HJL99] L. P. Hansen, J. A. Jacobsen, and R. A. Lund. The incidence of escaped farmed Atlantic salmon, Salmo salar L., in the Faroese fishery and estimates of catches of wild salmon. ICES Journal of Marine Science, 56(2):200–206, April 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
56/2/200/622481.

Heide-Jorgensen:1996:EDB

[HJR96] M. P. Heide-Jørgensen and Randall R. Reeves. Evidence of a decline in beluga, Delphinapterus leucas, abundance off West Greenland. ICES Journal of Marine Science, 53(1): 61–72, February 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/53/1/61/731583.

Haug:1991:HSP

[HKN^{+}91] Tore Haug, Ari B. Krøyer, Kjell T. Nilssen, Karl I. Ugland, and Paul E. Aspholm. Harp seal (Phoca groenlandica) invasions in Norwegian coastal waters: age composition and feeding habits. ICES Journal of Marine Science, 48(3):363– 371, November 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/48/3/363/737308.

Huttmann:1997:NSS

[HL97] F. Hüttmann and A. R. Lock. A new software system for the PIROP database: data flow and an approach for a seabirddepth analysis. ICES Journal of Marine Science, 54(4): 518–523, August 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/54/4/518/607433.

Hernandez-Leon:1995:UAT

[HLAM95] Santiago Hernández-León, Carlos Almeida, and Irene Montero. The use of aspartate transcarbamylase activity to estimate growth rates in zooplankton. ICES Journal of Marine Science, 52(3–4):377–383, June 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/52/3-4/377/841535.

Hearn:1991:ETS

[HLB91] William S. Hearn, George M. Leigh, and Raymond J. H. Beverton. An examination of a tag-shedding assumption, with application to southern bluefin tuna. ICES Journal of Marine Science, 48(1):41–51, May 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/48/1/41/696115.

Lo:1992:DFR

[hLHM^{+}92] N. Chyan huei Lo, John R. Hunter, H. Geoffrey Moser, Paul E. Smith, and Richard D. Methot, Jr. The daily fecundity reduction method: a new procedure for estimating adult fish biomass. ICES Journal of Marine Science, 49(2):209–215, May 1992. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/49/2/209/683907.

Hafsteinsson:1995:RMB

[HM95] M. T. Hafsteinsson and O. A. Misund. Recording the migration behaviour of fish schools by multi-beam sonar during conventional acoustic surveys. ICES Journal of Marine Science, 52(6): 915–924, December 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/52/6/915/649261.

Hatfield:1999:OAM

[HM99] E. M. C. Hatfield and A. W. A. Murray. Objective assessment of maturity in the Patagonian squid Loligo gahi (Cephalopoda: Loliginidae) from Falkland Islands waters. ICES Journal of Marine Science, 56(5):746–756, October 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
56/5/746/691340.

Hopkins:1990:PBD

[HN90] C. C. E. Hopkins and E. M. Nilssen. Population biology of the deep-water prawn (Pandalus borealis) in Balsfjord, northern Norway: I. Abundance, mortality, and growth, 1979–1983. ICES Journal of Marine Science, 47(2):148–166, ???? 1990. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
47/2/148/611309.

Horwood:1991:MREa

[HN91a] J. W. Horwood and M. D. Nicholson. A mark-recapture estimator of stock sizes from unrefined data applied to a stock of Dover sole (Solea solea L.). ICES Journal of Marine Science, 48(1):33–39, May 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/48/1/33/696112.

Horwood:1991:MREb

[HN91b] J. W. Horwood and M. D. Nicholson. A mark-recapture estimator of stock sizes from unrefined data applied to a stock of Dover sole (Solea solea L.). ICES Journal of Marine Science, 48(2):251, August 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/48/2/251/644534.

Huntley:1995:PCP

[HN95] Mark E. Huntley and Pearn P. Niiler. Physical control of population dynamics in the Southern Ocean. ICES Journal of Marine Science, 52(3–4):457–468, June 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https:/

/academic.oup.com/icesjms/article/52/3-4/457/841591.

Huse:1996:TAD

[HO96] Ingvar Huse and Egil Ona. Tilt angle distribution and swimming speed of overwintering Norwegian spring spawning herring. ICES Journal of Marine Science, 53(5):863–873, October 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/53/5/863/704339.

Hoggarth:1993:LHL

[Hog93] Daniel D. Hoggarth. The life history of the lithodid crab, Paralomis granulosa, in the Falkland Islands. ICES Journal of Marine Science, 50(4):405–424, ???? 1993. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/50/4/405/668237.

Hognestad:1994:LRH

[Hog94] Per T. Hognestad. The Lake Rossfjord herring (Clupea harengus L.) and its environment. ICES Journal of Marine Science, 51(3):281–292, ???? 1994. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/51/3/281/620195.

Horwood:1990:R

[Hor90] J. W. Horwood. Reviews. ICES Journal of Marine Science, 46(3):340, ???? 1990. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/46/3/340/596421.

Horwood:1991:ABM

[Hor91] Joseph Horwood. An approach to better management: the North Sea haddock. ICES Journal of Marine Science, 47 (3):318–332, ???? 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/47/3/318/601171.

Horbowy:1992:DAD

[Hor92a] Jan Horbowy. The differential alternative to the deriso difference production model. ICES Journal of Marine Science, 49 (2):167–174, May 1992. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/49/2/167/683881.

Horwood:1992:R

[Hor92b] Joseph Horwood. Reviews. ICES Journal of Marine Science, 49 (3):377–378, August 1992. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/49/3/377/819873.

Horwood:1993:FBP

[Hor93] J. W. Horwood. Fecundity and biomass of plaice (Pleuronectes platessa L.) in the northern Celtic Sea. ICES Journal of Marine Science, 50(3):315–323, ???? 1993. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/50/3/315/685317.

Houde:1990:RVI

[Hou90] E. D. Houde. Reproductive variability: Implications for the dynamics and well-being of fish stocks. introduction to selected papers. ICES Journal of Marine Science, 47(2):248, ???? 1990. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
47/2/248/611373.

Houde:1994:DBM

[Hou94] E. D. Houde. Differences between marine and freshwater fish larvae: implications for recruitment. ICES Journal of Marine Science, 51(1):91–97, ???? 1994. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/51/1/91/760310.

Holliday:1995:BOH

[HP95] D. V. Holliday and R. E. Pieper. Bioacoustical oceanography at high frequencies. ICES Journal of Marine Science, 52(3– 4):279–296, June 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.oup.

com/icesjms/article/52/3-4/279/841473.

Helle:1999:RSD

[HP99] K. Helle and M. Pennington. The relation of the spatial distribution of early juvenile cod (Gadus morhua L.) in the Barents Sea to zooplankton density and water flux during the period 1978–1984. ICES Journal of Marine Science, 56(1): 15–27, February 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/56/1/15/652192.

Hommen:1993:SMP

[HPDR93] U. Hommen, H.-J. Poethke, U. Dülmer, and H. T. Ratte. Simulation models to predict ecological risk of toxins in freshwater systems. ICES Journal of Marine Science, 50(4):337–347, ???? 1993. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/50/4/337/668219.

Hislop:1991:DFC

[HRBA91] J. R. G. Hislop, A. P. Robb, M. A. Bell, and D. W. Armstrong. The diet and food consumption of whiting (Merlangius merlangus) in the North Sea. ICES Journal of Marine Science, 48 (2):139–156, August 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/48/2/139/644457.

Hansen:1997:IRA

[HRL97] L. P. Hansen, D. G. Reddin, and R. A. Lund. The incidence of reared Atlantic salmon (Salmo salar L.) of fish farm origin at West Greenland. ICES Journal of Marine Science, 54(1): 152–155, February 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/54/1/152/782167.

Hatfield:1990:DDP

[HRP90] E. M. C. Hatfield, P. G. Rodhouse, and J. Porebski. Demography and distribution of the Patagonian squid (Loligo gahi d’Orbigny) during the austral winter. ICES Journal of Marine Science, 46(3):306–312, ???? 1990. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/46/3/306/596412.

Hauser:1991:UEP

[HS91] J. W. Hauser and M. P. Sissenwine. The uncertainty in estimates of the production of larval fish derived from samples of larval abundance. ICES Journal of Marine Science, 48(1):23– 32, May 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/48/1/23/696108.

Holst:1998:EJN

[HS98] J. C. Holst and A. Slotte. Effects of juvenile nursery on geographic spawning distribution in Norwegian spring-spawning herring (Clupea harengus L.). ICES Journal of Marine Science, 55(6):987–996, December 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/55/6/987/639009.

Hauck:1999:CSA

[HS99] M. Hauck and N. A. Sweijd. A case study of abalone poaching in South Africa and its impact on fisheries management. ICES Journal of Marine Science, 56(6):1024–1032, December 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
56/6/1024/658101.

Hall-Spencer:1999:IRT

[HSFAM99] J. M. Hall-Spencer, C. Froglia, R. J. A. Atkinson, and P. G. Moore. The impact of Rapido trawling for scallops, Pecten jacobaeus (L.), on the benthos of the Gulf of Venice. ICES Journal of Marine Science, 56(1):111–124, February 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
56/1/111/652187.

Hamer:1997:SVF

[HTG97] K. C. Hamer, D. R. Thompson, and C. M. Gray. Spatial variation in the feeding ecology, foraging ranges, and breeding energetics of northern fulmars in the north-east Atlantic Ocean. ICES Journal of Marine Science, 54(4):645–653, August 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
54/4/645/607477.

Hulme:1995:UVC

[Hul95] T. J. Hulme. The use of vertebral counts to discriminate between North Sea herring stocks. ICES Journal of Marine Science, 52(5):775–779, October 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/52/5/775/783812.

Hunt:1997:PZD

[Hun97] George L. Hunt, Jr. Physics, zooplankton, and the distribution of least auklets in the Bering Sea — a review. ICES Journal of Marine Science, 54(4):600–607, August 1997. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
54/4/600/607455.

Hutchings:1995:CPS

[HVMI^{+}95] L. Hutchings, H. M. Verheye, B. A. Mitchell-Innes, W. T. Peterson, J. A. Huggett, and S. J. Painting. Copepod production in the southern Benguela system. ICES Journal of Marine Science, 52(3–4):439–455, June 1995. CODEN ICESEC.

ISSN 1054-3139 (print), 1095-9289 (electronic). URL https:/

/academic.oup.com/icesjms/article/52/3-4/439/841582.

Hill:1990:DNL

[HW90] A. E. Hill and R. G. White. The dynamics of Norway lobster (Nephrops norvegicus L.) populations on isolated mud patches. ICES Journal of Marine Science, 46(2):167–174, ???? 1990. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
46/2/167/645359.

Harris:1997:BSD

[HW97] M. P. Harris and S. Wanless. Breeding success, diet, and brood neglect in the kittiwake (Rissa tridactyla) over an 11year period. ICES Journal of Marine Science, 54(4):615–623, August 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/54/4/615/607462.

Hansen:1997:I

[HWY97] L. P. Hansen, M. L. Windsor, and A. F. Youngson. Introduction. ICES Journal of Marine Science, 54(6):963–964, December 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/54/6/963/674068.

Iida:1996:RBA

[IMH96] Kohji Iida, Tohru Mukai, and DooJin Hwang. Relationship between acoustic backscattering strength and density of zooplankton in the sound-scattering layer. ICES Journal of Marine Science, 53(2):507–512, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/53/2/507/792958.

Imsland:1999:SMT

[Ims99] A. K. Imsland. Sexual maturation in turbot (Scophthalmus maximus) is related to genotypic oxygen affinity: experimental support for Pauly’s juvenile-to-adult transition hypothesis. ICES Journal of Marine Science, 56(3):320–325, June 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
56/3/320/623511.

Isaksson:1997:ASR

[IOEJ97] A. Isaksson, S. Oskarsson, S. M. Einarsson, and J. Jonasson. Atlantic salmon ranching: past problems and future management. ICES Journal of Marine Science, 54(6):1188–1199, December 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/54/6/1188/674019.

Jaffe:1996:TSB

[Jaf96] Jules S. Jaffe. Temporal sampling of backscattered sonar signals. ICES Journal of Marine Science, 53(2):359–361, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/53/2/359/792667.

Jakobsen:1992:BRP

[Jak92] Tore Jakobsen. Biological reference points for North-East Arctic cod and haddock. ICES Journal of Marine Science, 49 (2):155–166, May 1992. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/49/2/155/683875.

Jennings:1991:IVL

[JB91] S. Jennings and R. J. H. Beverton. Intraspecific variation in the life history tactics of Atlantic herring (Clupea harengus L.) stocks. ICES Journal of Marine Science, 48(1):117–125, May 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/48/1/117/696087.

Jauniaux:1997:LOS

[JBC97] T. Jauniaux, L. Brosens, and F. Coignoul. Lesions observed on stranded seabirds along the Belgian coast from 1992 to 1995. ICES Journal of Marine Science, 54(4):714–717, August 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
54/4/714/607503.

Jarre:1991:REG

[JCP91] A. Jarre, M. R. Clarke, and D. Pauly. Re-examination of growth estimates in oceanic squids: the case of Kondakovia longimana (Onychoteuthidae). ICES Journal of Marine Science, 48(2):195–200, August 1991. CODEN ICESEC.

ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/48/2/195/644489.

Jackson:1997:VPC

[JDLH97] D. Jackson, S. Deady, Y. Leahy, and D. Hassett. Variations in parasitic caligid infestations on farmed salmonids and implications for their management. ICES Journal of Marine Science, 54(6):1104–1112, December 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/54/6/1104/673951.

Jensen:1990:LOL

[Jen90] A. L. Jensen. Linearization with ordinary least-squares estimation, and nonlinear estimation of mortality using the catch equation. ICES Journal of Marine Science, 46(2):175–177, ???? 1990. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/46/2/175/645367.

Jensen:1991:MSF

[Jen91] A. L. Jensen. Multiple species fisheries with no ecological interaction: two-species Schaefer model applied to lake trout and lake whitefish. ICES Journal of Marine Science, 48(2):167–171, August 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/48/2/167/644472.

Jenkinson:1995:RTR

[Jen95] Ian R. Jenkinson. A review of two recent predation-rate models: the dome-shaped relationship between feeding rate and shear rate appears universal. ICES Journal of Marine Science, 52 (3–4):605–610, June 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.oup.

com/icesjms/article/52/3-4/605/841766.

Jensen:1999:CND

[Jen99] A. Jensen. Cessation of the Norwegian drift net fishery: changes observed in Norwegian and Russian populations of Atlantic salmon. ICES Journal of Marine Science, 56(1):84–95, February 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/56/1/84/652212.

Jossi:1993:ZTU

[JG93] J. W. Jossi and J. R. Goulet, Jr. Zooplankton trends: US northeast shelf ecosystem and adjacent regions differ from north-east Atlantic and North Sea. ICES Journal of Marine Science, 50 (3):303–313, ???? 1993. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/50/3/303/685316.

Jacobsen:1997:OOI

[JG97] J. A. Jacobsen and E. Gaard. Open-ocean infestation by salmon lice (Lepeophtheirus salmonis): comparison of wild and escaped farmed Atlantic salmon (Salmo salar L.). ICES Journal of Marine Science, 54(6):1113–1119, December 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
54/6/1113/673961.

Jiang:1996:DHM

[JJ96a] W. Jiang and T. Jørgensen. The diet of haddock (Melanogrammus aeglefinus L.) in the Barents Sea during the period 1984– 1991. ICES Journal of Marine Science, 53(1):11–21, February 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/53/1/11/731542.

Jiang:1996:E

[JJ96b] W. Jiang and T. Jorgensen. Erratum. ICES Journal of Marine Science, 53(3):641–642, June 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/53/3/641/625717.

Jackson:1999:APS

[JM99] G. D. Jackson and N. A. Moltschaniwskyj. Analysis of precision in statolith derived age estimates of the tropical squid Photololigo (Cephalopoda: Loliginidae). ICES Journal of Marine Science, 56(2):221–227, April 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/56/2/221/622491.

Jones:1990:LCA

[Jon90] R. Jones. Length-cohort analysis: The importance of choosing the correct growth parameters. ICES Journal of Marine Science, 46(2):133–139, ???? 1990. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/46/2/133/645340.

Jonsson:1997:REB

[Jon97] B. Jonsson. A review of ecological and behavioural interactions between cultured and wild Atlantic salmon. ICES Journal of Marine Science, 54(6):1031–1039, December 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
54/6/1031/673908.

Jorgensen:1990:LTC

[Jør90] Terje Jørgensen. Long-term changes in age at sexual maturity of Northeast Arctic cod (Gadus morhua L.). ICES Journal of Marine Science, 46(3):235–248, ???? 1990. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/46/3/235/596393.

Jorgensen:1992:LTC

[Jør92] Terje Jørgensen. Long-term changes in growth of North-east Arctic cod (Gadus morhua) and some environmental influences. ICES Journal of Marine Science, 49(3):263–278, August 1992. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
49/3/263/819745.

Junquera:1993:PDB

[JPG93] S. Junquera and G. Perez-Gándaras. Population diversity in Bay of Biscay anchovy (Engraulis encrasicolus L. 1758) as revealed by multivariate analysis of morphometric and meristic characters. ICES Journal of Marine Science, 50(4):383–391, ???? 1993. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/50/4/383/668229.

Jones:1990:ENM

[JS90] R. Jones and A. M. Shanks. An estimate of natural mortality for North Sea haddock. ICES Journal of Marine Science, 47 (1):99–103, ???? 1990. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/47/1/99/638554.

Jakobsson:1999:MSS

[JS99] J. Jakobsson and G. Stefánsson. Management of summerspawning herring off Iceland. ICES Journal of Marine Science, 56(6):827–833, December 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/56/6/827/658124.

Jonsson:1999:ECD

[JWF99] B. Jonsson, R. S. Waples, and K. D. Friedland. Extinction considerations for diadromous fishes. ICES Journal of Marine Science, 56(4):405–409, August 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/56/4/405/624618.

Kaiser:1998:CMB

[Kai98] M. Kaiser. Changes in megafaunal benthic communities in different habitats after trawling disturbance. ICES Journal of Marine Science, 55(3):353–361, June 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/55/3/353/655021.

Kappenman:1999:TSB

[Kap99] R. F. Kappenman. Trawl survey based abundance estimation using data sets with unusually large catches. ICES Journal of Marine Science, 56(1):28–35, February 1999. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
56/1/28/652195.

Kleppel:1995:EPN

[KB95] G. S. Kleppel and C. A. Burkart. Egg production and the nutritional environment of Acartia tonsa: the role of food quality in copepod nutrition. ICES Journal of Marine Science, 52(3– 4):297–304, June 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.oup.

com/icesjms/article/52/3-4/297/841483.

Kunitzer:1992:BIN

[KBC^{+}92] A. Kunitzer, D. Basford, J. A. Craeymeersch, J. M. Dewarumez, J. Dorjes, G. C. A. Duineveld, A. Eleftheriou, C. Heip, P. Herman, P. Kingston, U. Niermann, E. Rachor, H. Rumohr, and P. A. J. de Wilde. The benthic infauna of the North Sea: species distribution and assemblages. ICES Journal of Marine Science, 49(2):127–143, May 1992. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
49/2/127/683853.

Kelly:1999:AEG

[KCB99] C. J. Kelly, P. L. Connolly, and J. J. Bracken. Age estimation, growth, maturity, and distribution of the bluemouth rockfish Helicolenus d. dactylopterus (Delaroche 1809) from the Rockall Trough. ICES Journal of Marine Science, 56(1):61–74, February 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/56/1/61/652204.

Kattner:1995:PHC

[KH95] G. Kattner and W. Hagen. Polar herbivorous copepods — different pathways in lipid biosynthesis. ICES Journal of Marine Science, 52(3–4):329–335, June 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https:/

/academic.oup.com/icesjms/article/52/3-4/329/841503.

Kitada:1994:EMR

[KHK94] Shuichi Kitada, Kazuhiko Hiramatsu, and Hirohisa Kishino. Estimating mortality rates from tag recoveries: incorporating over-dispersion, correlation, and change points. ICES Journal of Marine Science, 51(3):241–251, ???? 1994. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
51/3/241/620173.

Kjellqwist:1995:TAC

[KHØ95] Sten Anders Kjellqwist, Tore Haug, and Torger Øritsland. Trends in age-composition, growth and reproductive parameters of Barents Sea harp seals, Phoca groenlandica. ICES Journal of Marine Science, 52(2):197–208, April 1995. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
52/2/197/593896.

Kimura:1997:BPM

[Kim97] S. Kimura. Biological productivity of meso-scale eddies caused by frontal disturbances in the Kuroshio. ICES Journal of Marine Science, 54(2):179–192, April 1997. CODEN ICESEC.

ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/54/2/179/620538.

Kingston:1992:IOO

[Kin92] P. F. Kingston. Impact of offshore oil production installations on the benthos of the North Sea. ICES Journal of Marine Science, 49(1):45–53, February 1992. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/49/1/45/732779.

Kristensen:1997:PRB

[KL97] Per Sand Kristensen and Hans Lassen. The production of relaid blue mussels (Mytilus edulis L.) in a Danish fjord. ICES Journal of Marine Science, 54(5):854–865, October 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
54/5/854/751506.

Kehayias:1996:DCS

[KLF96] G. Kehayias, J. Lykakis, and N. Fragopoulu. The diets of the chaetognaths Sagitta enflata, S. serratodentata atlantica and S. bipunctata at different seasons in Eastern Mediterranean coastal waters. ICES Journal of Marine Science, 53 (5):837–846, October 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/53/5/837/704332.

Kloser:1996:IPA

[Klo96] Rudy J. Kloser. Improved precision of acoustic surveys of benthopelagic fish by means of a deep-towed transducer. ICES Journal of Marine Science, 53(2):407–413, April 1996. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
53/2/407/792762.

Kideys:1994:OMA

[KN94] A. E. Kideys and U. Niermann. Occurrence of Mnemiopsis along the Turkish coast. ICES Journal of Marine Science, 51 (4):423–427, August 1994. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/51/4/423/714249.

Korsbrekke:1999:LSD

[KN99] K. Korsbrekke and O. Nakken. Length and species-dependent diurnal variation of catch rates in the Norwegian Barents Sea bottom-trawl surveys. ICES Journal of Marine Science, 56 (3):284–291, June 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/56/3/284/623500.

Kallio-Nyberg:1992:MPT

[KNI92] I. Kallio-Nyberg and E. Ikonen. Migration pattern of two salmon stocks in the Baltic Sea. ICES Journal of Marine Science, 49(2):191–198, May 1992. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/49/2/191/683893.

Knudsen:1990:BEI

[Knu90] Hans P. Knudsen. The Bergen echo integrator: an introduction. ICES Journal of Marine Science, 47(2):167–174, ???? 1990. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
47/2/167/611318.

Kalikhman:1997:PDF

[KO97] I. Kalikhman and I. Ostrovsky. Patchy distribution fields: survey design and adequacy of reconstruction. ICES Journal of Marine Science, 54(5):809–818, October 1997. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
54/5/809/751488.

Koie:1991:SNA

[Køi91] Marianne Køie. Swimbladder nematodes (Anguillicola spp.) and gill monogeneans (Pseudodactylogyrus spp.) parasitic on the European eel (Anguilla anguilla). ICES Journal of Marine Science, 47(3):391–398, ???? 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/47/3/391/601187.

Koie:1999:MPF

[Køi99] M. Køie. Metazoan parasites of flounder Platichthys flesus (L.) along a transect from the southwestern to the northeastern Baltic Sea. ICES Journal of Marine Science, 56(2):157–163, April 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/56/2/157/622455.

Kell:1999:EMP

[KOS^{+}99] L. T. Kell, C. M. O’Brien, M. T. Smith, T. K. Stokes, and B. D. Rackham. An evaluation of management procedures for implementing a precautionary approach in the ICES context for North Sea plaice (Pleuronectes platessa L.). ICES Journal of Marine Science, 56(6):834–845, December 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
56/6/834/658127.

Koljonen:1997:AUS

[KP97] M.-L. Koljonen and J. J. Pella. The advantage of using smolt age with allozymes for assessing wild stock contributions to Atlantic salmon catches in the Baltic Sea. ICES Journal of Marine Science, 54(6):1015–1030, December 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/54/6/1015/673898.

Kremer:1994:PAM

[Kre94] Patricia Kremer. Patterns of abundance for Mnemiopsis in US coastal waters: a comparative overview. ICES Journal of Marine Science, 51(4):347–354, August 1994. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/51/4/347/714238.

Kaaria:1997:SBS

[KRKS97] J. Kääriä, M. Rajasilta, M. Kurkilahti, and M. Soikkeli. Spawning bed selection by the Baltic herring (Clupea harengus membras) in the Archipelago of SW Finland. ICES Journal of Marine Science, 54(5):917–923, October 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/54/5/917/751558.

Kristiansen:1990:ESC

[KS90] T. S. Kristiansen and T. Svåsand. Enhancement studies of coastal cod in western Norway. part III. Interrelationships between reared and indigenous cod in a nearly land-locked fjord. ICES Journal of Marine Science, 47(1):23–29, ???? 1990.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
47/1/23/638505.

Kuikka:1996:IIC

[KSP96] S. Kuikka, P. Suuronen, and R. Parmanne. The impacts of increased codend mesh size on the northern Baltic herring fishery: ecosystem and market uncertainties. ICES Journal of Marine Science, 53(4):723–730, August 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/53/4/723/799746.

Kizner:1997:ISV

[KV97] Z. I. Kizner and D. A. Vasilyev. Instantaneous separable VPA (ISVPA). ICES Journal of Marine Science, 54(3):399–411, June 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/54/3/399/652478.

Kloser:1997:PAT

[KWK97] R. J. Kloser, A. Williams, and J. A. Koslow. Problems with acoustic target strength measurements of a deepwater fish, orange roughy (Hoplostethus atlanticus, Collett). ICES Journal of Marine Science, 54(1):60–71, February 1997. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
54/1/60/782187.

Kulka:1995:RCW

[KWN95] David W. Kulka, J. S. Wroblewski, and Savi Narayanan. Recent changes in the winter distribution and movements of northern Atlantic cod (Gadus morhua Linnaeus, 1758) on the Newfoundland-Labrador shelf. ICES Journal of Marine Science, 52(6):889–902, December 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/52/6/889/649254.

Kimura:1997:SSA

[KZ97] Daniel K. Kimura and Harold H. Zenger, Jr. Standardizing sablefish (Anoplopoma fimbria) long-line survey abundance indices by modeling the log-ratio of paired comparative fishing cpues. ICES Journal of Marine Science, 54(1):48–59, February 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/54/1/48/782184.

Laane:1996:CCV

[Laa96] R. Laane. Changes and causes of variability in salinity and dissolved inorganic phosphate in the Irish Sea, English Channel, and Dutch coastal zone. ICES Journal of Marine Science, 53 (6):933–944, December 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/53/6/933/592530.

Lambert:1990:EPS

[Lam90] Timothy C. Lambert. The effect of population structure on recruitment in herring. ICES Journal of Marine Science, 47 (2):249–255, ???? 1990. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/47/2/249/611378.

Large:1992:UMM

[Lar92] P. A. Large. Use of a multiplicative model to estimate relative abundance from commercial CPUE data. ICES Journal of Marine Science, 49(3):253–262, August 1992. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/49/3/253/819740.

Lindahl:1998:PPC

[LBDH98] Odd Lindahl, Andrea Belgrano, Lennart Davidsson, and Bodil Hernroth. Primary production, climatic oscillations, and physico-chemical processes: the Gullmar Fjord time-series data set (1985–1996). ICES Journal of Marine Science, 55(4): 723–729, August 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/55/4/723/698169.

Lillo:1996:TSM

[LCP96] S. Lillo, J. Cordova, and A. Paillaman. Target-strength measurements of hake and jack mackerel. ICES Journal of Marine Science, 53(2):267–271, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/53/2/267/792504.

Lipinski:1998:FVA

[LDU98] M. R. Lipinski, M. D. Durholtz, and L. G. Underhill. Field validation of age readings from the statoliths of chokka squid (Loligo vulgaris reynaudii d’Orbigny, 1845) and an assessment of associated errors. ICES Journal of Marine Science, 55 (2):240–257, April 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/55/2/240/781734.

Lacroix:1997:ASC

[LGKM97] G. L. Lacroix, B. J. Galloway, D. Knox, and D. MacLatchy. Absence of seasonal changes in reproductive function of cultured Atlantic salmon migrating into a Canadian river. ICES Journal of Marine Science, 54(6):1086–1091, December 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
54/6/1086/673938.

Lindstrom:1998:DHS

[LHHN98] U. Lindstrøm, A. Harbitz, T. Haug, and K. T. Nilssen. Do harp seals Phoca groenlandica exhibit particular prey preferences? ICES Journal of Marine Science, 55(5):941–953, October 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/55/5/941/691006.

Lipinski:1998:DNE

[LHSA98] M. R. Lipinski, I. Hampton, W. H. H. Sauer, and C. J. Augustyn. Daily net emigration from a spawning concentration of chokka squid (Loligo vulgaris reynaudii d’Orbigny, 1845) in Kromme Bay, South Africa. ICES Journal of Marine Science, 55(2):258–270, April 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/55/2/258/781741.

Lipinski:1999:R

[Lip99] Marek Lipinski. Review. ICES Journal of Marine Science, 56 (2):247–250, April 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/56/2/247/622499.

Lewy:1997:STL

[LL97] Peter Lewy and Hans Lassen. Should total landings be used to correct estimated catch in numbers or mean weight-at-age? ICES Journal of Marine Science, 54(3):462–470, June 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
54/3/462/652488.

Liao:1999:BDZ

[LLLL99] Cheng-Hsin Liao, Kuo-Tien Lee, Ming-Anne Lee, and Hsueh-Jung Lu. Biomass distribution and zooplankton composition of the sound-scattering layer in the waters of southern East China Sea. ICES Journal of Marine Science, 56(5):766–778, October 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/56/5/766/691344.

Lang:1999:BIS

[LM99] T. Lang and S. Mellergaard. The BMB/ICES Sea-going Workshop “Fish Diseases and Parasites in the Baltic Sea” — introduction and conclusions. ICES Journal of Marine Science, 56 (2):129–133, April 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/56/2/129/622429.

Lichatowich:1999:DEP

[LML99] J. Lichatowich, L. Mobrand, and L. Lestelle. Depletion and extinction of Pacific salmon (Oncorhynchus spp.): a different perspective. ICES Journal of Marine Science, 56(4):467–472, August 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/56/4/467/624635.

Lang:1999:SDG

[LMW^{+}99] T. Lang, S. Mellergaard, W. Wosniok, V. Kadakas, and K. Neumann. Spatial distribution of grossly visible diseases and parasites in flounder (Platichthys flesus) from the Baltic Sea: a synoptic survey. ICES Journal of Marine Science, 56(2):138–147, April 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/56/2/138/622440.

Laptikhovsky:1993:ESF

[LN93] Vladimir V. Laptikhovsky and Chingis M. Nigmatullin. Egg size, fecundity, and spawning in females of the genus Illex (Cephalopoda: Ommastrephidae). ICES Journal of Marine Science, 50(4):393–403, ???? 1993. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/50/4/393/668232.

Loeng:1997:WFT

[LOÅ97] Harald Loeng, Vladimir Ozhigin, and Bjørn Ådlandsvik. Water fluxes through the Barents Sea. ICES Journal of Marine Science, 54(3):310–317, June 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/54/3/310/652454.

Lockwood:1991:R

[Loc91] Stephen J. Lockwood. Review. ICES Journal of Marine Science, 47(3):376, ???? 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/47/3/376/601181.

Lokkeborg:1998:SCB

[Løk98] Svein Løkkeborg. Seabird by-catch and bait loss in longlining using different setting methods. ICES Journal of Marine Science, 55(1):145–149, February 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/55/1/145/615942.

Loring:1991:NHM

[Lor91] D. H. Loring. Normalization of heavy-metal data from estuarine and coastal sediments. ICES Journal of Marine Science, 48(1):101–115, May 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/48/1/101/696084.

Lapointe:1991:SCB

[LP91a] Michael F. Lapointe and Randall M. Peterman. Spurious correlations between fish recruitment and environmental factors due to errors in the natural mortality rate used in virtual population analysis (VPA). ICES Journal of Marine Science, 48 (2):219–228, August 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/48/2/219/644500.

Lucas:1991:JHF

[LP91b] Cyril E. Lucas and Basil B. Parrish. James Henry Fraser 18 May 1909–14 March 1990. ICES Journal of Marine Science, 48(1):137–138, May 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/48/1/137/696103.

Landry:1995:ZGP

[LPL95] M. R. Landry, W. K. Peterson, and C. J. Lorenzen. Zooplankton grazing, phytoplankton growth, and export flux: inferences from chlorophyll tracer methods. ICES Journal of Marine Science, 52(3–4):337–345, June 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https:/

/academic.oup.com/icesjms/article/52/3-4/337/841508.

Lewis:1996:VRS

[LR96] Thomas N. Lewis and Peter H. Rogers. The vibrational response of single-chambered fish swimbladders to low-frequency sound. ICES Journal of Marine Science, 53(2):285–287, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/53/2/285/792533.

Lawson:1999:IDA

[LR99] G. L. Lawson and G. A. Rose. The importance of detectability to acoustic surveys of semi-demersal fish. ICES Journal of Marine Science, 56(3):370–380, June 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/56/3/370/623524.

Lilliendahl:1997:ESF

[LS97] K. Lilliendahl and J. Solmundsson. An estimate of summer food consumption of six seabird species in Iceland. ICES Journal of Marine Science, 54(4):624–630, August 1997. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
54/4/624/607465.

Lane:1998:FRA

[LS98a] D. E. Lane and R. L. Stephenson. A framework for risk analysis in fisheries decision-making. ICES Journal of Marine Science, 55(1):1–13, February 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/55/1/1/615930.

Li:1998:TVC

[LS98b] Yaqin Li and Theodore J. Smayda. Temporal variability of chlorophyll in Narragansett Bay, 1973–1990. ICES Journal of Marine Science, 55(4):661–667, August 1998. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
55/4/661/698151.

Lane:1999:FMS

[LS99] D. E. Lane and R. L. Stephenson. Fisheries-management science: a framework for the implementation of fisheriesmanagement systems. ICES Journal of Marine Science, 56(6): 1059–1066, December 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/56/6/1059/658110.

Lenz:1995:IRV

[LSP^{+}95] J. Lenz, D. Schnack, D. Petersen, J. Kreikemeier, B. Hermann, S. Mees, and K. Wieland. The Ichthyoplankton Recorder: a video recording system for in situ studies of small-scale plankton distribution patterns. ICES Journal of Marine Science, 52 (3–4):409–417, June 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.oup.

com/icesjms/article/52/3-4/409/841556.

Lewy:1994:IDN

[LV94] Peter Lewy and Morten Vinther. Identification of Danish North Sea trawl fisheries. ICES Journal of Marine Science, 51(3): 263–272, ???? 1994. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/51/3/263/620186.

Misund:1992:SBF

[MA92a] Ole Arve Misund and Asgeir Aglen. Swimming behaviour of fish schools in the North Sea during acoustic surveying and pelagic trawl sampling. ICES Journal of Marine Science, 49 (3):325–334, August 1992. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/49/3/325/819798.

Moller:1992:EFD

[MA92b] H. Möller and K. Anders. Epidemiology of fish diseases in the Wadden Sea. ICES Journal of Marine Science, 49(2):199–208, May 1992. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/49/2/199/683900.

McCleave:1999:MYS

[MA99] J. D. McCleave and G. P. Arnold. Movements of yellow- and silver-phase European eels (Anguilla anguilla L.) tracked in the western North Sea. ICES Journal of Marine Science, 56 (4):510–536, August 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/56/4/510/624659.

Misund:1992:RBG

[MABD92] Ole Arve Misund, Asgeir Aglen, Arvid K. Beltestad, and John Dalen. Relationships between the geometric dimensions and biomass of schools. ICES Journal of Marine Science, 49(3): 305–315, August 1992. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/49/3/305/819779.

MacKenzie:1990:CPB

[Mac90] K. MacKenzie. Cestode parasites as biological tags for mackerel (Scomber scombrus L.) in the Northeast Atlantic. ICES Journal of Marine Science, 46(2):155–166, ???? 1990. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
46/2/155/645355.

MacLennan:1995:GSV

[Mac95] David N. MacLennan. Gear selectivity and the variation of yield. ICES Journal of Marine Science, 52(5):827–836, October 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/52/5/827/783868.

McClatchie:1996:RER

[MAC96] Sam McClatchie, Jon Alsop, and Roger F. Coombs. A reevaluation of relationships between fish size, acoustic frequency, and target strength. ICES Journal of Marine Science, 53(5): 780–791, October 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/53/5/780/704302.

MacKinnon:1998:HFI

[Mac98] B. M. MacKinnon. Host factors important in sea lice infections. ICES Journal of Marine Science, 55(2):188–192, April 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
55/2/188/781706.

Mackinson:1999:CSO

[Mac99] S. Mackinson. Cross-scale observations on distribution and behavioural dynamics of ocean feeding Norwegian springspawning herring (Clupea harengus L.). ICES Journal of Marine Science, 56(5):613–626, October 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/56/5/613/691313.

Madsen:1994:KPA

[Mad94] K. Popp Madsen. Knud Peter Andersen: 26 August 1925–13 April 1992. ICES Journal of Marine Science, 51(2):239–240, ???? 1994. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/51/2/239/674489.

Misund:1995:MSS

[MAF95] Ole Arve Misund, Asgeir Aglen, and Erik Frønæs. Mapping the shape, size, and density of fish schools by echo integration and a high-resolution sonar. ICES Journal of Marine Science, 52 (1):11–20, February 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/52/1/11/699731.

Miyashita:1996:SBT

[MAI96] Kazushi Miyashita, Ichiro Aoki, and Tadashi Inagaki. Swimming behaviour and target strength of isada krill (Euphausia pacifica). ICES Journal of Marine Science, 53(2):303–308, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/53/2/303/792561.

Maller:1990:SAM

[Mal90] R. A. Maller. Some aspects of a mixture model for estimating the boundary of a set of data. ICES Journal of Marine Science, 46(2):140–147, ???? 1990. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/46/2/140/645344.

Mann:1993:POF

[Man93] K. H. Mann. Physical oceanography, food chains, and fish stocks: a review. ICES Journal of Marine Science, 50(2): 105–119, ???? 1993. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/50/2/105/774774.

Martin:1991:R

[Mar91] A. R. Martin. Reviews. ICES Journal of Marine Science, 48(2):247, August 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/48/2/247/644517.

Martin:1993:NSI

[Mar93] J. H. A. Martin. Norwegian Sea intermediate water in the Faroe–Shetland Channel. ICES Journal of Marine Science, 50 (2):195–201, ???? 1993. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/50/2/195/774817.

Mauchline:1995:BAL

[Mau95] John Mauchline. Bathymetric adaptations of life history patterns of congeneric species (Euchaeta: Calanoida) in a 2000 m water column. ICES Journal of Marine Science, 52(3– 4):511–516, June 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.oup.

com/icesjms/article/52/3-4/511/841634.

Maucorps:1996:OA

[Mau96] Alain Maucorps. Opening address. ICES Journal of Marine Science, 53(6):885–886, December 1996. CODEN ICESEC.

ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/53/6/885/592516.

McClatchie:1996:CSM

[MAYC96] Sam McClatchie, Jon Alsop, Zhen Ye, and Roger F. Coombs. Consequence of swimbladder model choice and fish orientation to target strength of three New Zealand fish species. ICES Journal of Marine Science, 53(5):847–862, October 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
53/5/847/704336.

Macdonald:1996:CAM

[MB96a] R. W. Macdonald and J. M. Bewers. Contaminants in the Arctic marine environment: priorities for protection. ICES Journal of Marine Science, 53(3):537–563, June 1996. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
53/3/537/625704.

Misund:1996:TSE

[MB96b] Ole Arve Misund and Arvid K. Beltestad. Target-strength estimates of schooling herring and mackerel using the comparison method. ICES Journal of Marine Science, 53(2):281–284, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/53/2/281/792525.

Morgan:1997:TGV

[MB97] M. J. Morgan and W. R. Bowering. Temporal and geographic variation in maturity at length and age of Greenland halibut (Reinhardtius hippoglossoides) from the Canadian north-west Atlantic with implications for fisheries management. ICES Journal of Marine Science, 54(5):875–885, October 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
54/5/875/751534.

Myers:1996:CCE

[MBHQ96] R. A. Myers, N. J. Barrowman, J. M. Hoenig, and Z. Qu. The collapse of cod in Eastern Canada: the evidence from tagging data. ICES Journal of Marine Science, 53(3):629–640, June 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/53/3/629/625715.

Milton:1995:FEP

[MBR95] D. A. Milton, S. J. M. Blaber, and N. J. F. Rawlinson. Fecundity and egg production of four species of short-lived clupeoid from Solomon Islands, tropical South Pacific. ICES Journal of Marine Science, 52(1):111–125, February 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
52/1/111/699738.

Mayo:1990:GMI

[MBSM90] R. K. Mayo, J. Burnett, T. D. Smith, and C. A. Muchant. Growth-maturation interactions of Acadian redfish (Sebastes fasciatus Storer) in the Gulf of Maine–Georges Bank region of the Northwest Atlantic. ICES Journal of Marine Science, 46 (3):287–305, ???? 1990. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/46/3/287/596408.

Myers:1995:SRA

[MBT95] R. A. Myers, N. J. Barrowman, and K. R. Thompson. Synchrony of recruitment across the North Atlantic: an update. (or, “now you see it, now you don’t!”). ICES Journal of Marine Science, 52(1):103–110, February 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/52/1/103/699725.

Martin:1990:CAL

[MC90] I. Martin and R. M. Cook. Combined analysis of length and age-at-length data. ICES Journal of Marine Science, 46(2): 178–186, ???? 1990. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/46/2/178/645374.

Mulligan:1998:SBE

[MC98] T. J. Mulligan and D. G. Chen. A split-beam echo counting model: development of statistical procedures. ICES Journal of Marine Science, 55(5):905–917, October 1998. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
55/5/905/690997.

Morgan:1999:VMA

[MC99] M. J. Morgan and E. B. Colbourne. Variation in maturity-at-age and size in three populations of American plaice. ICES Journal of Marine Science, 56(5):673–688, October 1999. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
56/5/673/691324.

McDermott:1998:WPB

[McD98] John J. McDermott. The western Pacific brachyuran (Hemigrapsus sanguineus: Grapsidae), in its new habitat along the Atlantic coast of the United States: geographic distribution and ecology. ICES Journal of Marine Science, 55(2):289–298, April 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/55/2/289/781756.

McDowall:1999:DKD

[McD99] R. M. McDowall. Different kinds of diadromy: Different kinds of conservation problems. ICES Journal of Marine Science, 56 (4):410–413, August 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/56/4/410/624619.

McGinnity:1997:GIE

[McG97] P. McGinnity. Genetic impact of escaped farmed Atlantic salmon (Salmo salar L.) on native populations: use of DNA profiling to assess freshwater performance of wild, farmed, and hybrid progeny in a natural river environment. ICES Journal of Marine Science, 54(6):998–1008, December 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
54/6/998/674090.

McVicar:1997:DPI

[McV97] A. H. McVicar. Disease and parasite implications of the coexistence of wild and cultured Atlantic salmon populations. ICES Journal of Marine Science, 54(6):1093–1103, December 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
54/6/1093/673944.

Mazzocchi:1995:RPZ

[Md95] M. G. Mazzocchi and M. Ribera d’Alcalà. Recurrent patterns in zooplankton structure and succession in a variable coastal environment. ICES Journal of Marine Science, 52(3– 4):679–691, June 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.oup.

com/icesjms/article/52/3-4/679/841881.

Martinez:1997:RTN

[MEH97] I. Martinez, E. O. Elvevoll, and T. Haug. RAPD typing of north-east Atlantic minke whale (Balaenoptera acutorostrata). ICES Journal of Marine Science, 54(3):478–484, June 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
54/3/478/652495.

Moksness:1991:DSA

[MF91] E. Moksness and P. Fossum. Distinguishing spring- and autumn-spawned herring larvae (Clupea harengus L.) by otolith microstructure. ICES Journal of Marine Science, 48(1):61–66, May 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/48/1/61/696122.

Moksness:1992:DGR

[MF92] Erlend Moksness and Petter Fossum. Daily growth rate and hatching-date distribution of Norwegian spring-spawning herring (Clupea harengus L.). ICES Journal of Marine Science, 49(2):217–221, May 1992. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/49/2/217/683917.

Modin:1996:COM

[MFGP96] Johan Modin, Björn Fagerholm, Björn Gunnarsson, and Leif Pihl. Changes in otolith microstructure at metamorphosis of plaice, Pleuronectes platessa L. ICES Journal of Marine Science, 53(4):745–748, August 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/53/4/745/799773.

Mousseau:1998:APF

[MFL98] Laure Mousseau, Louis Fortier, and Louis Legendre. Annual production of fish larvae and their prey in relation to sizefractionated primary production (Scotian Shelf, NW Atlantic). ICES Journal of Marine Science, 55(1):44–57, February 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
55/1/44/615945.

Misund:1998:THS

[MFPT98] Ole Arve Misund, Anders Fernö, Tony Pitcher, and Bjørn Totland. Tracking herring schools with a high resolution sonar. variations in horizontal area and relative echo intensity. ICES Journal of Marine Science, 55(1):58–66, February 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
55/1/58/615948.

McFarlane:1993:R

[MGF93] G. A. McFarlane, Trygve. Gjedrem, and Michael J. Fogarty. Reviews. ICES Journal of Marine Science, 50(4):501–504, ???? 1993. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/50/4/501/668267.

Michalsen:1996:DVC

[MGF96] Kathrine Michalsen, Olav Rune Godø, and Anders Fernö. Diel variation in the catchability of gadoids and its influence on the reliability of abundance indices. ICES Journal of Marine Science, 53(2):389–395, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/53/2/389/792726.

McClatchie:1994:STV

[MGMS94] Sam McClatchie, Charles H. Greene, Michael C. Macaulay, and Darrell R. M. Sturley. Spatial and temporal variability of Antarctic krill: implications for stock assessment. ICES Journal of Marine Science, 51(1):11–18, ???? 1994. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
51/1/11/760283.

Morrison:1991:MMH

[MGN91] J. A. Morrison, J. C. Gamble, and I. R. Napier. Mass mortality of herring eggs associated with a sedimenting diatom bloom. ICES Journal of Marine Science, 48(2):237–245, August 1991.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
48/2/237/644514.

Marchal:1995:MAT

[MH95] Paul Marchal and Joseph Horwood. Multi-annual TACs and minimum biological levels. ICES Journal of Marine Science, 52 (5):797–807, October 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/52/5/797/783856.

MacLennan:1996:FPA

[MH96] David N. MacLennan and D. V. Holliday. Fisheries and plankton acoustics: past, present, and future. ICES Journal of Marine Science, 53(2):513–516, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/53/2/513/792974.

Millar:1997:EGH

[MH97] Russell B. Millar and René Holst. Estimation of gillnet and hook selectivity using log-linear models. ICES Journal of Marine Science, 54(3):471–477, June 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/54/3/471/652492.

Marchal:1998:IFM

[MH98a] P. Marchal and J. Horwood. Increasing fisheries management options with a flexible cost function. ICES Journal of Marine Science, 55(2):213–227, April 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/55/2/213/781720.

Mo:1998:OLS

[MH98b] T. A. Mo and P. A. Heuch. Occurrence of Lepeophtheirus salmonis (Copepoda: Caligidae) on sea trout (Salmo trutta) in the inner Oslo Fjord, south-eastern Norway. ICES Journal of Marine Science, 55(2):176–180, April 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/55/2/176/781697.

Mukai:1996:DDT

[MI96] Tohru Mukai and Kohji Iida. Depth dependence of target strength of live Kokanee salmon in accordance with Boyle’s law.

ICES Journal of Marine Science, 53(2):245–248, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
53/2/245/792462.

Millar:1992:MEE

[Mil92] Russell B. Millar. Modelling environmental effects on growth of cod: fitting to growth increment data versus fitting to sizeat-age data. ICES Journal of Marine Science, 49(3):289–295, August 1992. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/49/3/289/819764.

Millar:1994:STG

[Mil94] Russell B. Millar. Sampling from trawl gears used in sized selectivity experiments. ICES Journal of Marine Science, 51 (3):293–298, ???? 1994. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/51/3/293/620201.

Mills:1995:MSC

[Mil95] Claudia E. Mills. Medusae, siphonophores, and ctenophores as planktivorous predators in changing global ecosystems. ICES Journal of Marine Science, 52(3–4):575–581, June 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/52/3-4/575/841728.

Millner:1996:CLT

[Mil96] R. Millner. Comparison of long-term trends in growth of sole and plaice populations. ICES Journal of Marine Science, 53(6): 1196–1198, December 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/53/6/1196/592502.

Miller:1999:R

[Mil99] P. Miller. Review. ICES Journal of Marine Science, 56(3):404, June 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/56/3/404/623531.

Misund:1993:DMM

[Mis93] Ole Arve Misund. Dynamics of moving masses: variability in packing density, shape, and size among herring, sprat, and saithe schools. ICES Journal of Marine Science, 50(2):145–160, ???? 1993. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/50/2/145/774791.

Misund:1996:IMS

[Mis96] O. Misund. Improved mapping of schooling fish near the surface: comparison of abundance estimates obtained by sonar and echo integration. ICES Journal of Marine Science, 53 (2):383–388, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/53/2/383/792713.

McGehee:1996:TDS

[MJ96] Duncan McGehee and Jules S. Jaffe. Three-dimensional swimming behavior of individual zooplankters: observations using the acoustical imaging system FishTV. ICES Journal of Marine Science, 53(2):363–369, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/53/2/363/792682.

Magnusson:1997:DAG

[MJ97a] J. V. Magnússon and G. Jóhannesson. Distribution and abundance of 0-group redfish in the Irminger Sea and off East Greenland: relationships with adult abundance indices. ICES Journal of Marine Science, 54(5):830–845, October 1997. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
54/5/830/751497.

Malmberg:1997:TDC

[MJ97b] Svend-Aage Malmberg and Steingrímur Jónsson. Timing of deep convection in the Greenland and Iceland Seas. ICES Journal of Marine Science, 54(3):300–309, June 1997. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
54/3/300/652451.

Madin:1995:DFF

[MK95] L. P. Madin and P. Kremer. Determination of the filter-feeding rates of salps (tunicata, thaliacea). ICES Journal of Marine Science, 52(3–4):583–595, June 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https:/

/academic.oup.com/icesjms/article/52/3-4/583/841740.

Mulligan:1996:SBE

[MK96] Timothy J. Mulligan and Robert Kieser. A split-beam echocounting model for riverine use. ICES Journal of Marine Science, 53(2):403–406, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/53/2/403/792753.

McAllister:1998:BSA

[MK98] M. K. McAllister and G. P. Kirkwood. Bayesian stock assessment: a review and example application using the logistic model. ICES Journal of Marine Science, 55(6):1031–1060, December 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/55/6/1031/638969.

Mcallister:1999:AMC

[MK99a] M. K. Mcallister and G. P. Kirkwood. Applying multivariate conjugate priors in fishery-management system evaluation: how much quicker is it and does it bias the ranking of management options? ICES Journal of Marine Science, 56(6): 884–899, December 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/56/6/884/658139.

McKinnell:1999:STC

[MK99b] S. M. McKinnell and Ö. Karlström. Spatial and temporal covariation in the recruitment and abundance of Atlantic salmon populations in the Baltic Sea. ICES Journal of Marine Science, 56(4):433–443, August 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/56/4/433/624627.

Masse:1996:SSD

[MKP96] Jacques Massé, Constantin Koutsikopoulos, and Wilhelmina Patty. The structure and spatial distribution of pelagic fish schools in multispecies clusters: an acoustic study. ICES Journal of Marine Science, 53(2):155–160, April 1996. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
53/2/155/791603.

McLaren:1995:AEG

[ML95] Ian A. McLaren and Angela Leonard. Assessing the equivalence of growth and egg production of copepods. ICES Journal of Marine Science, 52(3–4):397–408, June 1995. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/52/3-
4/397/841545.

Marchal:1996:AEU

[ML96] Emile Marchal and Anne Lebourges. Acoustic evidence for unusual diel behaviour of a mesopelagic fish (Vinciguerria nimbaria) exploited by tuna. ICES Journal of Marine Science, 53 (2):443–447, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/53/2/443/792834.

Mellergaard:1999:DPB

[ML99] S. Mellergaard and T. Lang. Diseases and parasites of Baltic cod (Gadus morhua) from the Mecklenburg Bight to the Estonian coast. ICES Journal of Marine Science, 56(2):164–168, April 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/56/2/164/622458.

MacKenzie:1998:SLC

[MLBM98] K. MacKenzie, M. Longshaw, G. S. Begg, and A. H. McVicar. Sea lice (Copepoda: Caligidae) on wild sea trout (Salmo trutta L.) in Scotland. ICES Journal of Marine Science, 55 (2):151–162, April 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/55/2/151/781688.

Misund:1999:TSS

[MLCB99] O. A. Misund, N. Luyeye, J. Coetzee, and D. Boyer. Trawl sampling of small pelagic fish off Angola: effects of avoidance, towing speed, tow duration, and time of day. ICES Journal of Marine Science, 56(3):275–283, June 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/56/3/275/623497.

MacLennan:1996:IST

[MM96] D. N. MacLennan and A. Menz. Interpretation of in situ targetstrength data. ICES Journal of Marine Science, 53(2):233–236, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/53/2/233/792443.

Montevecchi:1997:CDO

[MM97] W. A. Montevecchi and R. A. Myers. Centurial and decadal oceanographic influences on changes in northern gannet populations and diets in the north-west Atlantic: implications for climate change. ICES Journal of Marine Science, 54(4):608–614, August 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/54/4/608/607460.

McClatchie:1998:TSS

[MMHC98] Sam McClatchie, Gavin Macaulay, Stuart Hanchet, and Roger F. Coombs. Target strength of southern blue whiting (Micromesistius australis) using swimbladder modelling, split beam and deconvolution. ICES Journal of Marine Science, 55 (3):482–493, June 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/55/3/482/655075.

Mohn:1999:RPS

[Moh99] R. Mohn. The retrospective problem in sequential population analysis: an investigation using cod fishery and simulated data. ICES Journal of Marine Science, 56(4):473–488, August 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
56/4/473/624639.

Moksness:1992:DOM

[Mok92a] Erlend Moksness. Differences in otolith microstructure and body growth rate of North Sea herring (Clupea harengus L.) larvae in the period 1987–1989. ICES Journal of Marine Science, 49(2):223–230, May 1992. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/49/2/223/683922.

Moksness:1992:VDI

[Mok92b] Erlend Moksness. Validation of daily increments in the otolith microstructure of Norwegian spring-spawning herring (Clupea harengus L.). ICES Journal of Marine Science, 49(2):231–235, May 1992. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/49/2/231/683925.

Moller:1990:ABD

[Möl90] H. Möller. Association between diseases of flounder (Platichthys flesus) and environmental conditions in the Elbe estuary, FRG. ICES Journal of Marine Science, 46(2):187–199, ???? 1990. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
46/2/187/645383.

Michalsen:1998:GNE

[MON98] K. Michalsen, G. Ottersen, and O. Nakken. Growth of Northeast Arctic cod (Gadus morhua L.) in relation to ambient temperature. ICES Journal of Marine Science, 55(5):863–877, October 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/55/5/863/690938.

Methven:1991:SAV

[MP91] David A. Methven and John F. Piatt. Seasonal abundance and vertical distribution of capelin (Mallotus villosus) in relation to water temperature at a coastal site off eastern Newfoundland. ICES Journal of Marine Science, 48(2):187–193, August 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
48/2/187/644485.

Martinez:1999:RTC

[MP99] I. Martinez and L. A. Pastene. RAPD-typing of Central and Eastern North Atlantic and Western North Pacific minke whales, Balaenoptera acutorostrata. ICES Journal of Marine Science, 56(5):640–651, October 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/56/5/640/691317.

Meincke:1997:AON

[MRF97] J. Meincke, B. Rudels, and H. J. Friedrich. The Arctic Ocean– Nordic Seas thermohaline system. ICES Journal of Marine Science, 54(3):283–299, June 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/54/3/283/652449.

Markussen:1992:FCA

[MRL92] Nina Hedlund Markussen, Morten Ryg, and Christian Lydersen. Food consumption of the NE Atlantic minke whale (Balaenoptera acutorostrata) population estimated with a simulation model. ICES Journal of Marine Science, 49(3):317–323, August 1992. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/49/3/317/819789.

Myers:1994:STR

[MRM^{+}94] R. A. Myers, A. A. Rosenberg, P. M. Mace, N. Barrowman, and V. R. Restrepo. In search of thresholds for recruitment overfishing. ICES Journal of Marine Science, 51(2):191–205, ???? 1994. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/51/2/191/674482.

Murphy:1994:MSE

[MRN94] E. J. Murphy, P. G. Rodhouse, and C. P. Nolan. Modelling the selective effects of fishing on reproductive potential and population structure of squid. ICES Journal of Marine Science, 51(3):299–313, ???? 1994. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/51/3/299/620220.

McClatchie:1991:SGM

[MRSF91] Sam McClatchie, Stanislaw Rakusa-Suszczewski, and Krzysztof Filcek. Seasonal growth and mortality of Euphausia superba in Admiralty Bay, South Shetland Islands, Antarctica. ICES Journal of Marine Science, 48(3):335–342, November 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
48/3/335/737280.

McMahon:1995:SOF

[MRTB95] T. McMahon, R. Raine, O. Titov, and S. Boychuk. Some oceanographic features of northeastern Atlantic waters west of Ireland. ICES Journal of Marine Science, 52(2):221–232, April 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/52/2/221/593903.

Mesnil:1990:HAL

[MS90] B. Mesnil and J. G. Shepherd. A hybrid age- and lengthstructured model for assessing regulatory measures in multiplespecies, multiple-fleet fisheries. ICES Journal of Marine Science, 47(2):115–132, ???? 1990. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/47/2/115/611288.

Meisfjord:1996:GVB

[MS96] J. Meisfjord and R. C. Sundt. Genetic variation between populations of the harp seal, Phoca groenlandica. ICES Journal of Marine Science, 53(1):89–95, February 1996. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
53/1/89/731592.

Maynou:1998:ASS

[MSC98] Francesc X. Maynou, Francisco Sardà, and Gérard Y. Conan. Assessment of the spatial structure and biomass evaluation of Nephrops norvegicus (L.) populations in the northwestern Mediterranean by geostatistics. ICES Journal of Marine Science, 55(1):102–120, February 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/55/1/102/615932.

Mitson:1996:UTF

[MSG96] R. B. Mitson, Y. Simard, and C. Goss. Use of a two-frequency algorithm to determine size and abundance of plankton in three widely spaced locations. ICES Journal of Marine Science, 53 (2):209–215, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/53/2/209/792405.

MacLennan:1992:FMV

[MSPG92] D. N. MacLennan, J. G. Shepherd, J. G. Pope, and H. Gislason. Fishing mortality and the variation of catches: a time series approach. ICES Journal of Marine Science, 49(4):425– 430, November 1992. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/49/4/425/653333.

Mcallister:1999:FQM

[MSRK99] M. K. Mcallister, P. J. Starr, V. R. Restrepo, and G. P. Kirkwood. Formulating quantitative methods to evaluate fisherymanagement systems: what fishery processes should be modelled and what trade-offs should be made? ICES Journal of Marine Science, 56(6):900–916, December 1999. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
56/6/900/658141.

Martin:1996:ACZ

[MSWL96] Linda V. Martin, Timothy K. Stanton, Peter H. Wiebe, and James F. Lynch. Acoustic classification of zooplankton. ICES Journal of Marine Science, 53(2):217–224, April 1996. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
53/2/217/792418.

Morgan:1996:SSR

[MT96] M. J. Morgan and E. A. Trippel. Skewed sex ratios in spawning shoals of Atlantic cod (Gadus morhua). ICES Journal of Marine Science, 53(5):820–826, October 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/53/5/820/704321.

McKinnell:1997:REC

[MT97a] S. McKinnell and A. J. Thomson. Recent events concerning Atlantic salmon escapees in the Pacific. ICES Journal of Marine Science, 54(6):1221–1225, December 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/54/6/1221/674049.

Mergardt:1997:DPF

[MT97b] N. Mergardt and A. Temming. Diel pattern of food intake in whiting (Merlangius merlangus) investigated from the weight of partly digested food particles in the stomach and laboratory determined particle decay functions. ICES Journal of Marine Science, 54(2):226–242, April 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/54/2/226/620550.

Mozetic:1998:SIA

[MUCM98] P. Mozetiĉ, S. Fonda Umani, B. Cataletto, and A. Malej. Seasonal and inter-annual plankton variability in the Gulf of Trieste (northern Adriatic). ICES Journal of Marine Science, 55 (4):711–722, August 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/55/4/711/698165.

Munilla:1997:HSC

[Mun97] Ignacio Munilla. Henslow’s swimming crab (Polybius henslowii) as an important food for yellow-legged gulls (Larus cachinnans) in NW Spain. ICES Journal of Marine Science, 54(4): 631–634, August 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/54/4/631/607469.

Murray:1996:CGR

[Mur96] A. W. A. Murray. Comparison of geostatistical and random sample survey analyses of Antarctic krill acoustic data. ICES Journal of Marine Science, 53(2):415–421, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
53/2/415/792783.

Mutlu:1994:DNI

[Mut94] E. Mutlu. Distribution of the new invader Mnemiopsis sp. and the resident Aurelia aurita and Pleurobrachia pileus populations in the Black Sea in the years 1991–1993. ICES Journal of Marine Science, 51(4):407–421, August 1994. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
51/4/407/714248.

Mutlu:1996:TSC

[Mut96] Erhan Mutlu. Target strength of the common jellyfish (Aurelia aurita): a preliminary experimental study with a dualbeam acoustic system. ICES Journal of Marine Science, 53 (2):309–311, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/53/2/309/792569.

Misra:1990:WPA

[MUV90] R. K. Misra, J. F. Uthe, and W. Vyncke. A weighted procedure of analysing for time trends in contaminant levels in CMP data: application to cod and flounder data of the Belgian coast (1978–1985). ICES Journal of Marine Science, 47(1):65–75, ???? 1990. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/47/1/65/638533.

Millner:1996:LTC

[MW96] R. S. Millner and C. L. Whiting. Long-term changes in growth and population abundance of sole in the North Sea from 1940 to the present. ICES Journal of Marine Science, 53(6):1185– 1195, December 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/53/6/1185/592497.

MacKenzie:1997:TSI

[MW97] B. R. MacKenzie and F. E. Werner. Theme session on intermediate-scale physical processes and their influence on the transport and food environment of fish. ICES Journal of Marine Science, 54(2):159, April 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/54/2/159/620530.

Mills:1998:UCP

[MWRH98] D. K. Mills, A. Walne, P. C. Reid, and S. I. Heaney. Updating the continuous plankton recorder: an improved tool for integrated plankton monitoring. ICES Journal of Marine Science, 55(4):814–817, August 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/55/4/814/698257.

Nakashima:1990:EDI

[Nak90] B. S. Nakashima. Escapement from a diamond IX midwater trawl during acoustic surveys for capelin (Mallotus villosus) in the Northwest Atlantic. ICES Journal of Marine Science, 47(1):76–82, ???? 1990. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/47/1/76/638540.

Nejstgaard:1995:ACC

[NBBS95] J. C. Nejstgaard, U. Båmstedt, E. Bagøien, and P. T. Solberg. Algal constraints on copepod grazing. Growth state, toxicity, cell size, and season as regulating factors. ICES Journal of Marine Science, 52(3–4):347–357, June 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https:/

/academic.oup.com/icesjms/article/52/3-4/347/841517.

Nedreaas:1990:ADN

[Ned90] Kjell Nedreaas. Age determination of Northeast Atlantic sebastes species. ICES Journal of Marine Science, 47(2):208–230, ???? 1990. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/47/2/208/611347.

Nehring:1998:ETP

[Neh98] Stefan Nehring. Establishment of thermophilic phytoplankton species in the North Sea: biological indicators of climatic changes? ICES Journal of Marine Science, 55(4):818–823, August 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/55/4/818/698265.

Nicholson:1998:MTP

[NFM98] Mike Nicholson, Rob Fryer, and David Maxwell. Multivariate trends in phytoplankton species groups in the southern North Sea. ICES Journal of Marine Science, 55(4):581–586, August 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/55/4/581/698137.

Nilssen:1995:FHH

[NHP^{+}95] Kjell T. Nilssen, Tore Haug, Vladimir Potelov, Vladimir A. Stasenkov, and Yuri K. Timoshenko. Food habits of harp seals (Phoca groenlandica) during lactation and moult in March-May in the southern Barents Sea and White Sea. ICES Journal of Marine Science, 52(1):33–41, February 1995. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
52/1/33/699778.

Niermann:1994:DAE

[Nie94] U. Niermann. Distribution of anchovy eggs and larvae (Engraulis encrasicolus Cuv.) in the Black Sea in 1991–1992. ICES Journal of Marine Science, 51(4):395–406, August 1994. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
51/4/395/714246.

Nielsen:1999:EHS

[Nie99] J. L. Nielsen. The evolutionary history of steelhead (Oncorhynchus mykiss) along the US Pacific Coast: Developing a conservation strategy using genetic diversity. ICES Journal of Marine Science, 56(4):449–458, August 1999. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
56/4/449/624633.

Nilsson:1997:MMV

[Nil97] J. Nilsson. MtDNA and microsatellite variation in Baltic Atlantic salmon. ICES Journal of Marine Science, 54(6):1173– 1176, December 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/54/6/1173/674009.

Nedreaas:1994:GSR

[NJN94] Kjell Nedreaas, Torild Johansen, and Gunnar Nævdal. Genetic studies of redfish (Sebastes spp.) from Icelandic and Greenland waters. ICES Journal of Marine Science, 51(4):461–467, August 1994. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/51/4/461/714258.

Nordeide:1999:SSC

[NK99] J. T. Nordeide and E. Kjellsby. Sound from spawning cod at their spawning grounds. ICES Journal of Marine Science, 56 (3):326–332, June 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/56/3/326/623513.

Nielsen:1999:HES

[NL99] J. Rasmus Nielsen and Bo Lundgren. Hydroacoustic ex situ target strength measurements on juvenile cod (Gadus morhua L.). ICES Journal of Marine Science, 56(5):627–639, October 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/56/5/627/691314.

Njifonjou:1999:MFM

[NLCS99] O. Njifonjou, F. Laloë, C. Chaboud, and M. Simier. Marine fisheries management in Cameroon: overview and perspectives for improvement of the survey system and data collection. ICES Journal of Marine Science, 56(6):1033–1043, December 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
56/6/1033/658102.

Noack:1997:CSL

[NLP97] P. T. Noack, L. M. Laird, and I. G. Priede. Carotenoids of sea lice (Lepeophtheirus salmonis) as potential indicators of host Atlantic salmon (Salmo salar L.) origin. ICES Journal of Marine Science, 54(6):1140–1143, December 1997. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
54/6/1140/673982.

Nedreaas:1991:GSR

[NN91a] Kjell Nedreaas and Gunnar Nævdal. Genetic studies of redfish (Sebastes spp.) along the continental slopes from Norway to East Greenland. ICES Journal of Marine Science, 48(2): 173–186, August 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/48/2/173/644480.

Nedreaas:1991:IGR

[NN91b] Kjell Nedreaas and Gunnar Nævdal. Identification of 0- and 1-group redfish (genus Sebastes) using electrophoresis. ICES Journal of Marine Science, 48(1):91–99, May 1991. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
48/1/91/696138.

Nottestad:1998:EGB

[Nøt98] L. Nøttestad. Extensive gas bubble release in Norwegian springspawning herring (Clupea harengus) during predator avoidance.

ICES Journal of Marine Science, 55(6):1133–1140, December 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/55/6/1133/638994.

Nordeide:1998:HFV

[NP98] Jarle Tryti Nordeide and Ivar H. Pettersen. Haemoglobin frequencies and vertebral numbers of cod (Gadus morhua L.) off northern Norway — test of a population structure hypothesis. ICES Journal of Marine Science, 55(1):134–140, February 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/55/1/134/615935.

Nielsen:1996:CCF

[NR96] E. Nielsen and K. Richardson. Can changes in the fisheries yield in the Kattegat (1950–1992) be linked to changes in primary production? ICES Journal of Marine Science, 53(6): 988–994, December 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/53/6/988/592545.

Northridge:1995:E

[NTWW95a] S. P. Northridge, M. L. Tasker, A. Webb, and J. M. Williams. Erratum. ICES Journal of Marine Science, 52(6):1005– 1012, December 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/52/6/1005/649244.

Northridge:1995:DRA

[NTWW95b] Simon P. Northridge, Mark L. Tasker, Andrew Webb, and James M. Williams. Distribution and relative abundance of harbour porpoises (Phocoena phocoena L.), white-beaked dolphins (Lagenorhynchus albirostris Gray), and minke whales (Balaenoptera acutorostrata Lacepède) around the British Isles. ICES Journal of Marine Science, 52(1):55–66, February 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
52/1/55/699791.

Nichols:1993:FOD

[NW93] J. H. Nichols and S. Warnes. Field observations of the daily pattern of spawning in the western stock of mackerel (Scomber scombrus L.). ICES Journal of Marine Science, 50(2):219–225, ???? 1993. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/50/2/219/774837.

Orr:1997:MAF

[OB97] D. C. Orr and W. R. Bowering. A multivariate analysis of food and feeding trends among Greenland halibut (Reinhardtius hippoglossoides) sampled in Davis Strait, during 1986. ICES Journal of Marine Science, 54(5):819–829, October 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
54/5/819/751493.

OBrien:1999:ND

[O’B99] C. M. O’Brien. A note on the distribution of G_{loss}. ICES Journal of Marine Science, 56(2):180–183, April 1999. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
56/2/180/622469.

Ollason:1997:PSD

[OBD^{+}97] J. G. Ollason, A. D. Bryant, P. M. Davis, B. E. Scott, and M. L. Tasker. Predicted seabird distributions in the North Sea: the consequences of being hungry. ICES Journal of Marine Science, 54(4):507–517, August 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/54/4/507/607427.

OConnell:1992:EIM

[ODR92] M. F. O’Connell, J. B. Dempson, and D. G. Reddin. Evaluation of the impacts of major management changes in the Atlantic salmon (Salmo salar L.) fisheries of Newfoundland and Labrador, Canada, 1984–1988. ICES Journal of Marine Science, 49(1):69–87, February 1992. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/49/1/69/732788.

Oelofsen:1999:FMN

[Oel99] B. W. Oelofsen. Fisheries management: the Namibian approach. ICES Journal of Marine Science, 56(6):999–1004, December 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/56/6/999/658162.

OBrien:1995:IDA

[OF95] Barry O’Brien and Julie M. Fives. Ichthyoplankton distribution and abundance off the west coast of Ireland. ICES Journal of Marine Science, 52(2):233–245, April 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/52/2/233/593905.

Oh:1999:PDC

[OHN99] C.-W. Oh, R. G. Hartnoll, and R. D. M. Nash. Population dynamics of the common shrimp, Crangon crangon (L.), in Port Erin Bay, Isle of Man, Irish Sea. ICES Journal of Marine Science, 56(5):718–733, October 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/56/5/718/691333.

Omori:1995:LHS

[OIF95] Makoto Omori, Haruto Ishii, and Ai Fujinaga. Life history strategy of Aurelia aurita (Cnidaria, Scyphomedusae) and its impact on the zooplankton community of Tokyo Bay. ICES Journal of Marine Science, 52(3–4):597–603, June 1995. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/52/3-4/597/841751.

Ojaveer:1998:HZD

[OLO98] Evald Ojaveer, Alide Lumberg, and Henn Ojaveer. Highlights of zooplankton dynamics in Estonian waters (Baltic Sea). ICES Journal of Marine Science, 55(4):748–755, August 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
55/4/748/698199.

Ona:1996:ASS

[OM96] E. Ona and R. B. Mitson. Acoustic sampling and signal processing near the seabed: the deadzone revisited. ICES Journal of Marine Science, 53(4):677–690, August 1996. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
53/4/677/799719.

OFlynn:1997:CCT

[OMF^{+}97] F. M. O’Flynn, S. A. McGeachy, G. W. Friars, T. J. Benfey, and J. K. Bailey. Comparisons of cultured triploid and diploid Atlantic salmon(Salmo salar L.). ICES Journal of Marine Science, 54(6):1160–1165, December 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/54/6/1160/673999.

Ottersen:1998:ATDa

[OMN98a] Geir Ottersen, Kathrine Michalsen, and Odd Nakken. Ambient temperature and distribution of north-east Arctic cod. ICES Journal of Marine Science, 55(1):67–85, February 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
55/1/67/615950.

Ottersen:1998:ATDb

[OMN98b] Geir Ottersen, Kathrine Michalsen, and Odd Nakken. Ambient temperature and distribution of north-east Arctic cod. ICES Journal of Marine Science, 55(5):979, October 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
55/5/979/691022.

Oro:1997:ETD

[OR97] Daniel Oro and Xavier Ruiz. Exploitation of trawler discards by breeding seabirds in the north-western Mediterranean: differences between the Ebro Delta and the Balearic Islands areas. ICES Journal of Marine Science, 54(4):695–707, August 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
54/4/695/607492.

Orlowski:1999:ASS

[Orl99] A. Orlowski. Acoustic studies of spatial gradients in the Baltic: Implications for fish distribution. ICES Journal of Marine Science, 56(4):561–570, August 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/56/4/561/624679.

Ostvedt:1990:EES

[Øst90] Ole J. Østvedt. Emmy Egidius: 1 September 1929–3 February 1989. ICES Journal of Marine Science, 46(2):224–225, ???? 1990. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/46/2/224/645404.

Ostvedt:1999:GSA

[Øst99] Ole J. Østvedt. Gunnar Sætersdal. 20 April 1922–10 July 1997. ICES Journal of Marine Science, 56(1):126–127, February 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/56/1/126/652191.

Olaso:1998:IDB

[OVP98] I. Olaso, F. Velasco, and N. Pérez. Importance of discarded blue whiting (Micromesistius poutassou) in the diet of lesser spotted dogfish (Scyliorhinus canicula) in the Cantabrian Sea. ICES Journal of Marine Science, 55(3):331–341, June 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
55/3/331/655009.

Ohman:1995:IM

[OW95] Mark D. Ohman and Simon N. Wood. The inevitability of mortality. ICES Journal of Marine Science, 52(3–4):517– 522, June 1995. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/52/3-4/517/841640.

Parsons:1999:TEC

[Par99] E. C. M. Parsons. Trace element concentrations in whole fish from North lantau waters, Hong Kong. ICES Journal of Marine Science, 56(5):791–794, October 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/56/5/791/691350.

Patterson:1998:AFS

[Pat98] K. R. Patterson. Assessing fish stocks when catches are misreported: model, simulation tests, and application to cod, haddock, and whiting in the ICES area. ICES Journal of Marine Science, 55(5):878–891, October 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/55/5/878/690941.

Payne:1999:SVR

[Pay99] Andrew I. L. Payne. Some views of representatives of the South African fishing industry about local fisheries management. ICES Journal of Marine Science, 56(6):1067–1072, December 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/56/6/1067/658111.

Pillar:1997:DVB

[PB97] Stan C. Pillar and Manuel Barange. Diel variability in bottom trawl catches and feeding activity of the Cape hakes off the west coast of South Africa. ICES Journal of Marine Science, 54(3):485–499, June 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/54/3/485/652497.

Pierce:1998:DAF

[PBSN98] G. J. Pierce, N. Bailey, Y. Stratoudakis, and A. Newton. Distribution and abundance of the fished population of Loligo forbesi in Scottish waters: analysis of research cruise data. ICES Journal of Marine Science, 55(1):14–33, February 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
55/1/14/615938.

Paul:1991:IVC

[PCH91] A. J. Paul, K. O. Coyle, and L. Haldorson. Interannual variations in copepod nauplii prey of larval fish in an Alaskan Bay. ICES Journal of Marine Science, 48(2):157–165, August 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
48/2/157/644465.

Porteiro:1996:ASA

[PCM96] C. Porteiro, P. Carrera, and J. Miquel. Analysis of Spanish acoustic surveys for sardine, 1991–1993: abundance estimates and inter-annual variability. ICES Journal of Marine Science, 53(2):429–433, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/53/2/429/792810.

Pedersen:1995:FHS

[Ped95] Søren Anker Pedersen. Feeding habits of starry ray (Raja radiata) in West Greenland waters. ICES Journal of Marine Science, 52(1):43–53, February 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/52/1/43/699783.

Pedersen:1996:MCB

[Ped96a] Britta Pedersen. Metal concentration in biota in the North sea: changes and causes. ICES Journal of Marine Science, 53(6): 1008–1013, December 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/53/6/1008/592456.

Pedersen:1996:DFL

[Ped96b] Jens Pedersen. Discrimination of fish layers using the three-dimensional information obtained by a split-beam echosounder. ICES Journal of Marine Science, 53(2):371–376, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/53/2/371/792694.

Pepin:1990:BCR

[Pep90] P. Pepin. Biological correlates of recruitment variability in North Sea fish stocks. ICES Journal of Marine Science, 47 (1):89–98, ???? 1990. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/47/1/89/638550.

Pepin:1999:PRD

[PES99] P. Pepin, G. T. Evans, and T. H. Shears. Patterns of RNA/ DNA ratios in larval fish and their relationship to survival in the field. ICES Journal of Marine Science, 56(5):697–706, October 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/56/5/697/691328.

Petitgas:1993:GFS

[Pet93] Pierre Petitgas. Geostatistics for fish stock assessments: a review and an acoustic application. ICES Journal of Marine Science, 50(3):285–298, ???? 1993. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/50/3/285/685311.

Petitgas:1997:SED

[Pet97] Pierre Petitgas. Sole egg distributions in space and time characterised by a geostatistical model and its estimation variance. ICES Journal of Marine Science, 54(2):213–225, April 1997.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
54/2/213/620547.

Petitgas:1998:BDD

[Pet98] P. Petitgas. Biomass-dependent dynamics of fish spatial distributions characterized by geostatistical aggregation curves. ICES Journal of Marine Science, 55(3):443–453, June 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
55/3/443/655062.

Punt:1994:CFM

[PH94] André E. Punt and Ray Hilborn. A comparison of fishery models with and without cannibalism with implications for the management of the Cape hake resource off southern Africa. ICES Journal of Marine Science, 51(1):19–29, ???? 1994. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
51/1/19/760296.

Philippart:1998:LTI

[Phi98] Catharina J. M. Philippart. Long-term impact of bottom fisheries on several by-catch species of demersal fish and benthic invertebrates in the south-eastern North Sea. ICES Journal of Marine Science, 55(3):342–352, June 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/55/3/342/655012.

Prestidge:1995:MIC

[PHT95] M. C. Prestidge, R. P. Harris, and A. H. Taylor. A modelling investigation of copepod egg production in the Irish sea. ICES Journal of Marine Science, 52(3–4):693–703, June 1995. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/52/3-4/693/841889.

Poulet:1995:TMS

[PILB95] S. A. Poulet, A. Ianora, M. Laabir, and W. C. M. Klein Breteler. Towards the measurement of secondary production and recruitment in copepods. ICES Journal of Marine Science, 52(3–4):359–368, June 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/52/3-4/359/841526.

Patterson:1995:CPA

[PK95a] K. R. Patterson and G. P. Kirkwood. Comparative performance of ADAPT and laurec–shepherd methods for estimating fish population parameters and in stock management. ICES Journal of Marine Science, 52(2):183–196, April 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
52/2/183/593894.

Pedersen:1995:FWG

[PK95b] S. A. Pedersen and P. Kanneworff. Fish on the West Greenland shrimp grounds, 1988–1992. ICES Journal of Marine Science, 52(2):165–182, April 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/52/2/165/593892.

Polacheck:1999:IEM

[PKMP99] T. Polacheck, N. L. Klaer, C. Millar, and A. L. Preece. An initial evaluation of management strategies for the southern bluefin tuna fishery. ICES Journal of Marine Science, 56(6): 811–826, December 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/56/6/811/658122.

Pelletier:1992:MUU

[PL92] D. Pelletier and A. Laurec. Management under uncertainty: defining strategies for reducing overexploitation. ICES Journal of Marine Science, 49(4):389–401, November 1992. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
49/4/389/653316.

Petitgas:1996:SOP

[PL96] Pierre Petitgas and Jean Jacques Levenez. Spatial organization of pelagic fish: echogram structure, spatio-temporal condition, and biomass in Senegalese waters. ICES Journal of Marine Science, 53(2):147–153, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/53/2/147/791595.

Pech:1997:UPC

[PL97] Nicolas Pech and Francis Laloë. Use of Principal Component Analysis with Instrumental Variables (PCAIV) to analyse fisheries catch data. ICES Journal of Marine Science, 54 (1):32–47, February 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/54/1/32/782182.

Potter:1990:CMI

[PLPN90] D. C. Potter, R. G. Lough, R. I. Perry, and J. D. Neilson. Comparison of the MOCNESS and IYGPT pelagic samplers for the capture of 0-group cod (Gadus morhua) on Georges Bank. ICES Journal of Marine Science, 46(2):121–128, ???? 1990. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
46/2/121/645331.

Philippart:1996:LTF

[PLvdM^{+}96] Catharina J. M. Philippart, Han J. Lindeboom, Jaap van der Meer, Henk W. van der Veer, and Johannes Ij. Witte. Longterm fluctuations in fish recruit abundance in the western Wadden Sea in relation to variation in the marine environment. ICES Journal of Marine Science, 53(6):1120–1129, December 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/53/6/1120/592483.

Pavlov:1996:SSD

[PM96a] Dimitri A. Pavlov and Erlend Moksness. Sensitive stages during embryonic development of wolffish, Anarhichas lupus L. determining the final numbers of rays in unpaired fins and skeletal abnormalities. ICES Journal of Marine Science, 53(4):731–740, August 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/53/4/731/799756.

Pope:1996:ESS

[PM96b] J. G. Pope and C. T. Macer. An evaluation of the stock structure of North Sea cod, haddock, and whiting since 1920, together with a consideration of the impacts of fisheries and predation effects on their biomass and recruitment. ICES Journal of Marine Science, 53(6):1157, December 1996. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
53/6/1157/592491.

Pitcher:1996:ABH

[PMF^{+}96] Tony J. Pitcher, Ole Arve Misund, Anders Fernö, Bjørn Totland, and Vebjørn Melle. Adaptive behaviour of herring schools in the Norwegian Sea as revealed by high-resolution sonar. ICES Journal of Marine Science, 53(2):449–452, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
53/2/449/792847.

Pavlov:1993:LHP

[PN93] D. A. Pavlov and G. G. Novikov. Life history and peculiarities of common wolffish (Anarhichas lupus) in the White Sea. ICES Journal of Marine Science, 50(3):271–277, ???? 1993. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
50/3/271/685305.

Pollock:1993:ROR

[Pol93] David E. Pollock. Recruitment overfishing and resilience in spiny lobster populations. ICES Journal of Marine Science, 50(1):9–14, ???? 1993. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/50/1/9/652892.

Paul:1990:EIC

[PPS90a] A. J. Paul, J. M. Paul, and R. L. Smith. Energy ingestion and conversion rate in pollock (Theragra chalcogramma) fed different prey types. ICES Journal of Marine Science, 46 (3):232–234, ???? 1990. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/46/3/232/596392.

Paul:1990:ROC

[PPS90b] A. J. Paul, J. M. Paul, and R. L. Smith. Rates of oxygen consumption of yellowfin sole (Limanda aspera (Pallas)) relative to body size, food intake, and temperature. ICES Journal of Marine Science, 47(2):205–207, ???? 1990. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/47/2/205/611338.

Paul:1992:ERR

[PPS92] A. J. Paul, J. M. Paul, and R. L. Smith. Energy and ration requirements of fiathead sole (Hippoglossoides elassodon Jordan and Gilbert 1880) based on energy consumption and growth. ICES Journal of Marine Science, 49(4):413–416, November 1992. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/49/4/413/653326.

Pedersen:1993:FHR

[PR93] S. A. Pedersen and F. Riget. Feeding habits of redfish (Sebastes spp.) and Greenland halibut (Reinhardtius hippoglossoides) in West Greenland waters. ICES Journal of Marine Science, 50 (4):445–459, ???? 1993. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/50/4/445/668247.

Piet:1998:CDF

[PR98] G. J. Piet and A. D. Rijnsdorp. Changes in the demersal fish assemblage in the south-eastern North Sea following the establishment of a protected area (“plaice box”). ICES Journal of Marine Science, 55(3):420–429, June 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/55/3/420/655047.

Petrakis:1997:SSD

[PS97] George Petrakis and Konstantinos I. Stergiou. Size selectivity of diamond and square mesh codends for four commercial Mediterranean fish species. ICES Journal of Marine Science, 54(1):13–23, February 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/54/1/13/782151.

Punt:1999:HSE

[PS99] A. E. Punt and A. D. M. Smith. Harvest strategy evaluation for the eastern stock of gemfish (Rexea solandri). ICES Journal of Marine Science, 56(6):860–875, December 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
56/6/860/658134.

Planque:1998:LTC

[PT98] Benjamin Planque and Arnold H. Taylor. Long-term changes in zooplankton and the climate of the North Atlantic. ICES Journal of Marine Science, 55(4):644–654, August 1998. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
55/4/644/698148.

Pauly:1998:DCT

[PTCC98] D. Pauly, A. W. Trites, E. Capuli, and V. Christensen. Diet composition and trophic levels of marine mammals. ICES Journal of Marine Science, 55(3):467–481, June 1998. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
55/3/467/655070.

Pedersen:1995:WDC

[PTO95] G. Pedersen, K. Tande, and G. O. Ottesen. Why does a component of Calanus finmarchicus stay in the surface waters during the overwintering period in high latitudes? ICES Journal of Marine Science, 52(3–4):523–531, June 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/52/3-4/523/841649.

Pihl:1993:MPJ

[PU93] Leif Pihl and Mats Ulmestrand. Migration pattern of juvenile cod (Gadus morhua) on the Swedish west coast. ICES Journal of Marine Science, 50(1):63–70, ???? 1993. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/50/1/63/652885.

Purdom:1991:Ra

[Pur91a] C. E. Purdom. Reviews. ICES Journal of Marine Science, 48(2):248, August 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/48/2/248/644522.

Purdom:1991:Rb

[Pur91b] C. E. Purdom. Reviews. ICES Journal of Marine Science, 48 (2):248–249, August 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/48/2/248/644526.

Polacheck:1993:ASV

[PV93] Tom Polacheck and Jon Helge Vølstad. Analysis of spatial variability of Georges Bank haddock (Melanogrammus aeglefinus) from trawl survey data using a linear regression model with spatial interaction. ICES Journal of Marine Science, 50(1):1– 8, ???? 1993. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/50/1/1/652869.

Piontkovski:1995:MVT

[PW95] S. A. Piontkovski and R. Williams. Multiscale variability of tropical ocean zooplankton biomass. ICES Journal of Marine Science, 52(3–4):643–656, June 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https:/

/academic.oup.com/icesjms/article/52/3-4/643/841851.

Ribic:1997:RSA

[RA97] Christine A. Ribic and D. G. Ainley. The relationships of seabird assemblages to physical habitat features in Pacific equatorial waters during spring 1984–1991. ICES Journal of Marine Science, 54(4):593–599, August 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/54/4/593/607453.

Radach:1998:QLT

[Rad98] Günther Radach. Quantification of long-term changes in the German Bight using an ecological development index. ICES Journal of Marine Science, 55(4):587–599, August 1998. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
55/4/587/698138.

Reeves:1992:EMS

[RAFC92] S. A. Reeves, D. W. Armstrong, R. J. Fryer, and K. A. Coull. The effects of mesh size, cod-end extension length and codend diameter on the selectivity of Scottish trawls and seines. ICES Journal of Marine Science, 49(3):279–288, August 1992. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
49/3/279/819752.

Raine:1993:DBD

[Rai93] R. Raine. The development of a bloom of the dinoflagellate [Gyrodinium aureolum (Hulbert)] on the south-west Irish coast. ICES Journal of Marine Science, 50(4):461–469, ???? 1993. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
50/4/461/668249.

Rajasilta:1993:SHC

[Raj93] M. Rajasilta. Spawning of herring (Clupea harengus membras L.) in the Archipelago Sea. ICES Journal of Marine Science, 50(3):233–246, ???? 1993. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/50/3/233/685298.

Ramster:1991:WR

[Ram91] John Ramster. Working with RJHB. ICES Journal of Marine Science, 48(3):373–374, November 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/48/3/373/737316.

Rijnsdorp:1998:MSD

[RBSV98] A. D. Rijnsdorp, A. M. Buys, F. Storbeck, and E. G. Visser. Micro-scale distribution of beam trawl effort in the southern North Sea between 1993 and 1996 in relation to the trawling frequency of the sea bed and the impact on benthic organisms. ICES Journal of Marine Science, 55(3):403–419, June 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
55/3/403/655043.

Richardson:1993:ADG

[RCE^{+}93] C. A. Richardson, S. A. Collis, K. Ekaratne, P. Dare, and D. Key. The age determination and growth rate of the European flat oyster, Ostrea edulis, in British waters determined from acetate peels of umbo growth lines. ICES Journal of Marine Science, 50(4):493–500, ???? 1993. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/50/4/493/668262.

Rankine:1990:VHL

[RCM90] Paul W. Rankine, Lindsay H. Cargill, and John A. Morrison. Variation in the hatching length of spring-spawned herring larvae (Clupea harengus L.) on Ballantrae Bank in the Firth of Clyde. ICES Journal of Marine Science, 46(3):333–339, ???? 1990. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/46/3/333/596419.

Rose:1995:CGM

[RdC95] G. A. Rose, B. deYoung, and E. B. Colbourne. Cod (Gadus morhua L.) migration speeds and transport relative to currents on the north-east Newfoundland shelf. ICES Journal of Marine Science, 52(6):903–913, December 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/52/6/903/649259.

Ribic:1997:DSN

[RDHP97] Christine A. Ribic, Randall Davis, Nancy Hess, and Dwight Peake. Distribution of seabirds in the northern Gulf of Mexico in relation to mesoscale features: initial observations. ICES Journal of Marine Science, 54(4):545–551, August 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
54/4/545/607440.

Rijnsdorp:1991:RVN

[RDvBH91] A. D. Rijnsdorp, N. Daan, F. A. van Beek, and H. J. L. Heessen. Reproductive variability in North Sea plaice, sole, and cod. ICES Journal of Marine Science, 47(3):352–375, ???? 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
47/3/352/601178.

Reynisson:1996:ETI

[Rey96] Páll Reynisson. Evaluation of threshold-induced bias in the integration of single-fish echoes. ICES Journal of Marine Science, 53(2):345–350, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/53/2/345/792644.

Reynisson:1998:MEB

[Rey98] P. Reynisson. Monitoring of equivalent beam angles of hullmounted acoustic survey transducers in the period 1983–1995. ICES Journal of Marine Science, 55(6):1125–1132, December 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/55/6/1125/638991.

Richardson:1993:MDV

[RG93] Linda R. Richardson and John R. Gold. Mitochondrial DNA variation in red grouper (Epinephelus morio) and greater amberjack (Seriola dumerili) from the Gulf of Mexico. ICES Journal of Marine Science, 50(1):53–62, ???? 1993. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
50/1/53/652883.

Rice:1996:PCS

[RG96] Jake Rice and Henrik Gislason. Patterns of change in the size spectra of numbers and diversity of the North Sea fish assemblage, as reflected in surveys and models. ICES Journal of Marine Science, 53(6):1214–1225, December 1996. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
53/6/1214/592508.

Ross:1993:EMS

[RGH93] A. H. Ross, W. S. C. Gurney, and M. R. Heath. Ecosystem models of Scottish sea lochs for assessing the impact of nutrient enrichment. ICES Journal of Marine Science, 50(4):359–367, ???? 1993. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/50/4/359/668224.

Roe:1996:VBD

[RGHC96] Howard S. J. Roe, Gwyn Griffiths, Mark Hartman, and Nick Crisp. Variability in biological distributions and hydrography from concurrent Acoustic Doppler Current Profiler and Sea-Soar surveys. ICES Journal of Marine Science, 53(2):131–138, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/53/2/131/791566.

Richter:1995:SCV

[Ric95] Claudio Richter. Seasonal changes in the vertical distribution of mesozooplankton in the Greenland Sea Gyre (75^{o}n): distribution strategies of calanoid copepods. ICES Journal of Marine Science, 52(3–4):533–539, June 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https:/

/academic.oup.com/icesjms/article/52/3-4/533/841660.

Rijnsdorp:1991:CFF

[Rij91] A. D. Rijnsdorp. Changes in fecundity of female North Sea plaice (Pleuronectes platessa L.) between three periods since 1900. ICES Journal of Marine Science, 48(3):253–280, November 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/48/3/253/737240.

Ritter:1997:CAS

[Rit97] J. A. Ritter. The contribution of Atlantic salmon (Salmo salar L.) enhancement to a sustainable resource. ICES Journal of Marine Science, 54(6):1177–1187, December 1997. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
54/6/1177/674014.

Rijnsdorp:1990:SSM

[RJ90] A. D. Rijnsdorp and A. Jaworski. Size-selective mortality in plaice and cod eggs: a new method in the study of egg mortality. ICES Journal of Marine Science, 47(2):256–263, ???? 1990. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
47/2/256/611385.

Rosenberg:1992:CIC

[RKCM92] A. A. Rosenberg, G. P. Kirkwood, R. M. Cook, and R. A. Myers. Combining information from commercial catches and research surveys to estimate recruitment: a comparison of methods. ICES Journal of Marine Science, 49(4):379–387, November 1992. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/49/4/379/653314.

Rijnsdorp:1996:TPD

[RM96a] A. D. Rijnsdorp and R. S. Millner. Trends in population dynamics and exploitation of North Sea plaice (Pleuronectes platessa L.) since the late 1800s. ICES Journal of Marine Science, 53(6):1170–1184, December 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/53/6/1170/592494.

Rogers:1996:FAA

[RM96b] S. I. Rogers and R. S. Millner. Factors affecting the annual abundance and regional distribution of English inshore demersal fish populations: 1973 to 1995. ICES Journal of Marine Science, 53(6):1094–1112, December 1996. CODEN ICESEC.

ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/53/6/1094/592477.

Robb:1992:CGB

[Rob92] A. P. Robb. Changes in the gall bladder of whiting (Merlangius merlangus) in relation to recent feeding history. ICES Journal of Marine Science, 49(4):431–436, November 1992. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
49/4/431/653335.

Rochet:1998:STE

[Roc98] Marie-Joëlle Rochet. Short-term effects of fishing on life history traits of fishes. ICES Journal of Marine Science, 55(3):371–391, June 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/55/3/371/655033.

Rantajarvi:1998:ESF

[ROH^{+}98] Eija Rantajärvi, Riitta Olsonen, Seija Hällfors, Juha-Markku Leppänen, and Mika Raateoja. Effect of sampling frequency on detection of natural variability in phytoplankton: unattended high-frequency measurements on board ferries in the Baltic Sea. ICES Journal of Marine Science, 55(4):697–704, August 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
55/4/697/698160.

Rose:1998:ATS

[Ros98] G. A. Rose. Acoustic target strength of capelin in Newfoundland waters. ICES Journal of Marine Science, 55(5): 918–923, October 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/55/5/918/690999.

Rijnsdorp:1995:MSD

[RP95] A. D. Rijnsdorp and M. A. Pastoors. Modelling the spatial dynamics and fisheries of North Sea plaice (Pleuronectes platessa L.) based on tagging data. ICES Journal of Marine Science, 52(6):963–980, December 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/52/6/963/649277.

Rose:1996:TSS

[RP96] G. A. Rose and D. R. Porter. Target-strength studies on Atlantic cod (Gadus morhua) in Newfoundland waters. ICES Journal of Marine Science, 53(2):259–265, April 1996. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
53/2/259/792495.

Richardson:1998:ENP

[RP98] Katherine Richardson and Flemming Bo Pedersen. Estimation of new production in the North Sea: consequences for temporal and spatial variability of phytoplankton. ICES Journal of Marine Science, 55(4):574–580, August 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/55/4/574/698134.

Restrepo:1999:PCR

[RP99] V. R. Restrepo and J. E. Powers. Precautionary control rules in US fisheries management: specification and performance. ICES Journal of Marine Science, 56(6):846–852, December 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
56/6/846/658130.

Rees:1999:CBB

[RPW^{+}99] H. L. Rees, M. A. Pendle, R. Waldock, D. S. Limpenny, and S. E. Boyd. A comparison of benthic biodiversity in the North Sea, English Channel, and Celtic Seas. ICES Journal of Marine Science, 56(2):228–246, April 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/56/2/228/622497.

Robinson:1998:DIE

[RR98] R. F. Robinson and C. A. Richardson. The direct and indirect effects of suction dredging on a razor clam (Ensis arcuatus) population. ICES Journal of Marine Science, 55(5): 970–977, October 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/55/5/970/691012.

Reisenbichler:1999:GCA

[RR99] R. R. Reisenbichler and S. P. Rubin. Genetic changes from artificial propagation of Pacific salmon affect the productivity and viability of supplemented populations. ICES Journal of Marine Science, 56(4):459–466, August 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/56/4/459/624634.

Rees:1992:STT

[RRL^{+}92] H. L. Rees, S. M. Rowlatt, M. A. Lambert, R. G. Lees, and D. S. Limpenny. Spatial and temporal trends in the benthos and sediments in relation to sewage sludge disposal off the northeast coast of England. ICES Journal of Marine Science, 49 (1):55–64, February 1992. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/49/1/55/732784.

Rasmussen:1999:SVC

[RSB99] E. B. Rasmussen, M.-B. Salhauge, and J. Boje. The suitability of vertebral counts in stock delineation studies of Greenland halibut, Reinhardtius hippoglossoides (Walbaum), in West Greenland. ICES Journal of Marine Science, 56(1): 75–83, February 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/56/1/75/652209.

Reeve:1995:I

[RSH95] Michael R. Reeve, Hein Rune Skjoldal, and Roger Harris. Introduction. ICES Journal of Marine Science, 52(3–4):261– 262, June 1995. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/52/3-4/261/841447.

Ransom:1996:CHN

[RSN96] Bruce H. Ransom, Tracey W. Steig, and Patrick A. Nealson. Comparison of hydroacoustic and net catch estimates of Pacific salmon smolt (Oncorhynchus spp.) passage at hydropower dams in the Columbia River Basin, USA. ICES Journal of Marine Science, 53(2):477–481, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/53/2/477/792912.

Rudjakov:1995:SVM

[RTK95] J. A. Rudjakov, V. B. Tseitlin, and V. J. Kitain. Seasonal variations of mesoplankton biomass in the upper layer of the Bering sea; understanding biomass oscillations in the ocean. ICES Journal of Marine Science, 52(3–4):747–753, June 1995. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/52/3-4/747/841934.

Reid:1997:CSP

[RTWC97] David G. Reid, William R. Turrell, Martin Walsh, and Ad Corten. Cross-shelf processes north of Scotland in relation to the southerly migration of western mackerel. ICES Journal of Marine Science, 54(2):168–178, April 1997. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
54/2/168/620535.

Reddin:1990:IPG

[RVD90] D. G. Reddin, E. Verspoor, and P. R. Downton. An integrated phenotypic and genotypic approach to the stock discrimination of Atlantic salmon. ICES Journal of Marine Science, 47(1):83–88, ???? 1990. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/47/1/83/638544.

Rijnsdorp:1996:CGN

[RvL96] A. D. Rijnsdorp and P. I. van Leeuwen. Changes in growth of North Sea plaice since 1950 in relation to density, eutrophication, beam-trawl effort, and temperature. ICES Journal of Marine Science, 53(6):1199–1213, December 1996. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
53/6/1199/592504.

Rijnsdorp:1996:CAD

[RvLDH96] Adriaan D. Rijnsdorp, Piet I. van Leeuwen, Niels Daan, and Henk J. L. Heessen. Changes in abundance of demersal fish species in the North Sea between 1906–1909 and 1990–1995. ICES Journal of Marine Science, 53(6):1054–1062, December 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/53/6/1054/592469.

Richardson:1991:ASP

[RW91] C. A. Richardson and P. Walker. The age structure of a population of the hard-shell clam, Mercenaria mercenaria from Southampton Water, England, derived from acetate peel replicas of shell sections. ICES Journal of Marine Science, 48(2): 229–236, August 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/48/2/229/644507.

Revie:1990:IFE

[RWJF90] J. Revie, D. E. Weston, F. R. Harden Jones, and G. P. Fox. Identification of fish echoes located at 65 km range by shorebased sonar. ICES Journal of Marine Science, 46(3):313–324, ???? 1990. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/46/3/313/596415.

Ryman:1997:MAE

[Rym97] N. Ryman. Minimizing adverse effects of fish culture: understanding the genetics of populations with overlapping generations. ICES Journal of Marine Science, 54(6):1149–1159, December 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/54/6/1149/673994.

Simmonds:1996:SIU

[SAC96] John E. Simmonds, F. Armstrong, and Philip J. Copland. Species identification using wideband backscatter with neural network and discriminant analysis. ICES Journal of Marine Science, 53(2):189–195, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/53/2/189/792373.

Sakshaug:1997:BPD

[Sak97] Egil Sakshaug. Biomass and productivity distributions and their variability in the Barents Sea. ICES Journal of Marine Science, 54(3):341–350, June 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/54/3/341/652465.

Sampson:1990:LSP

[Sam90] David B. Sampson. A length-structured population model for southern fin whales and a test for density dependence. ICES Journal of Marine Science, 46(3):249–268, ???? 1990. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
46/3/249/596396.

Sampson:1991:FTF

[Sam91] David B. Sampson. Fishing tactics and fish abundance, and their influence on catch rates. ICES Journal of Marine Science, 48(3):291–301, November 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/48/3/291/737249.

Schneider:1994:PDT

[SB94] G. Schneider and G. Behrends. Population dynamics and the trophic role of Aurelia aurita medusae in the Kiel Bight and western Baltic. ICES Journal of Marine Science, 51(4): 359–367, August 1994. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/51/4/359/714240.

Soule:1995:EBE

[SBH95] Michael Soule, Manuel Barange, and Ian Hampton. Evidence of bias in estimates of target strength obtained with a splitbeam echo-sounder. ICES Journal of Marine Science, 52(1): 139–144, February 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/52/1/139/699753.

Soule:1997:PNP

[SBSH97] Michael Soule, Manuel Barange, Haakon Solli, and Ian Hampton. Performance of a new phase algorithm for discriminating between single and overlapping echoes in a split-beam echosounder. ICES Journal of Marine Science, 54(5):934– 938, October 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/54/5/934/751569.

Stokes:1999:CUE

[SBSP99] T. K. Stokes, D. S. Butterworth, R. L. Stephenson, and A. I. L. Payne. Confronting uncertainty in the evaluation and implementation of fisheries-management systems: Introduction. ICES Journal of Marine Science, 56(6):795–796, December 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/56/6/795/658118.

Spencer:1996:SPP

[SC96] Paul D. Spencer and Jeremy S. Collie. A simple predator–prey model of exploited marine fish populations incorporating alternative prey. ICES Journal of Marine Science, 53(3):615–628, June 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/53/3/615/625714.

Schwalme:1999:SDF

[SC99] K. Schwalme and G. A. Chouinard. Seasonal dynamics in feeding, organ weights, and reproductive maturation of Atlantic cod (Gadus morhua) in the southern Gulf of St Lawrence. ICES Journal of Marine Science, 56(3):303–319, June 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
56/3/303/623510.

Shackell:1994:MPS

[SCM94] N. L. Shackell, J. E. Carscadden, and D. S. Miller. Migration of pre-spawning capelin (Mallotus villosus) as related to temperature on the northern Grand Bank, Newfoundland. ICES Journal of Marine Science, 51(1):107–114, ???? 1994. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
51/1/107/760280.

Stanton:1996:ASC

[SCW96] Timothy K. Stanton, Dezhang Chu, and Peter H. Wiebe. Acoustic scattering characteristics of several zooplankton groups. ICES Journal of Marine Science, 53(2):289–295, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/53/2/289/792541.

Scalabrin:1996:NAI

[SDW^{+}96] Carla Scalabrin, Nöel Diner, Alain Weill, Alain Hillion, and Marie-Catherine Mouchot. Narrowband acoustic identification of monospecific fish shoals. ICES Journal of Marine Science, 53(2):181–188, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/53/2/181/792331.

Selivanovsky:1996:SSH

[SE96] Dmitry A. Selivanovsky and Alexander B. Ezersky. Sound scattering by hydrodynamic wakes of sea animals. ICES Journal of Marine Science, 53(2):377–381, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/53/2/377/792705.

Soldal:1997:SYG

[SE97] Aud Vold Soldal and Arill Engås. Survival of young gadoids excluded from a shrimp trawl by a rigid deflecting grid. ICES Journal of Marine Science, 54(1):117–124, February 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
54/1/117/782137.

Salminen:1998:CCR

[SE98] M. Salminen and E. Erkamo. Comparison of coastal and river releases of Atlantic salmon smolts in the river Kokemäenjoki, Baltic Sea. ICES Journal of Marine Science, 55(6):1071– 1081, December 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/55/6/1071/638975.

Serebryakov:1990:PFR

[Ser90] V. P. Serebryakov. Population fecundity and reproductive capacity of some food fishes in relation to year-class-strength fluctuations. ICES Journal of Marine Science, 47(2):267–272, ???? 1990. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/47/2/267/611394.

Simmonds:1996:WBR

[SF96] E. J. Simmonds and R. J. Fryer. Which are better, random or systematic acoustic surveys? A simulation using North Sea herring as an example. ICES Journal of Marine Science, 53 (1):39–50, February 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/53/1/39/731551.

Stratoudakis:1999:FDS

[SFCP99] Y. Stratoudakis, R. J. Fryer, R. M. Cook, and G. J. Pierce. Fish discarded from Scottish demersal vessels: Estimators of total discards and annual estimates for targeted gadoids. ICES Journal of Marine Science, 56(5):592–605, October 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
56/5/592/691309.

Soria:1996:AVI

[SFG96] Marc Soria, Pierre Fréon, and François Gerlotto. Analysis of vessel influence on spatial behaviour of fish schools using a multi-beam sonar and consequences for biomass estimates by echo-sounder. ICES Journal of Marine Science, 53(2):453–458, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/53/2/453/792859.

Sorokin:1995:TCM

[SG95] Yu. I. Sorokin and O. Giovanardi. Trophic characteristics of the Manila clam (Tapes philippinarum Adams and Reeve). ICES Journal of Marine Science, 52(5):853–862, October 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
52/5/853/783881.

Straile:1998:RDC

[SG98] Dietmar Straile and Walter Geller. The response of Daphnia to changes in trophic status and weather patterns: a case study from Lake Constance. ICES Journal of Marine Science, 55 (4):775–782, August 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/55/4/775/698229.

Steele:1995:PCP

[SH95] John H. Steele and Eric W. Henderson. Predation control of plankton demography: III. Food chain control. ICES Journal of Marine Science, 52(3–4):565–573, June 1995. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/52/3-
4/565/841714.

Secor:1998:ULS

[SH98] D. H. Secor and E. D. Houde. Use of larval stocking in restoration of Chesapeake Bay striped bass. ICES Journal of Marine Science, 55(2):228–239, April 1998. CODEN ICESEC.

ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/55/2/228/781728.

Soule:1996:PIC

[SHB96] Michael Soule, Ian Hampton, and Manuel Barange. Potential improvements to current methods of recognizing single targets with a split-beam echo-sounder. ICES Journal of Marine Science, 53(2):237–243, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/53/2/237/792451.

Shepherd:1991:SMS

[She91] J. G. Shepherd. Simple methods for short-term forecasting of catch and biomass. ICES Journal of Marine Science, 48(1):67– 78, May 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/48/1/67/696127.

Shepherd:1997:PYC

[She97] J. G. Shepherd. Prediction of year-class strength by calibration regression analysis of multiple recruit index series. ICES Journal of Marine Science, 54(5):741–752, October 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
54/5/741/751464.

Shepherd:1999:ESA

[She99] J. G. Shepherd. Extended survivors analysis: an improved method for the analysis of catch-at-age data and abundance indices. ICES Journal of Marine Science, 56(5):584–591, October 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/56/5/584/691307.

Saegrov:1997:EFA

[SHKL97] H. Sægrov, K. Hindar, S. Kålås, and H. Lura. Escaped farmed Atlantic salmon replace the original salmon stock in the River Vosso, western Norway. ICES Journal of Marine Science, 54 (6):1166–1172, December 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/54/6/1166/674003.

Secor:1995:MRE

[SHM95] D. H. Secor, E. D. Houde, and D. M. Monteleone. A markrelease experiment on larval striped bass Morone saxatilis in a Chesapeake Bay tributary. ICES Journal of Marine Science, 52 (1):87–101, February 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/52/1/87/699813.

Simmonds:1996:FPA

[Sim96] John E. Simmonds. Fisheries and plankton acoustics. ICES Journal of Marine Science, 53(2):129, April 1996. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
53/2/129/791553.

Sindermann:1991:CHE

[Sin91] Carl J. Sindermann. Case histories of effects of transfers and introductions on marine resources. ICES Journal of Marine Science, 47(3):377–378, ???? 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/47/3/377/601184.

Steig:1996:MFM

[SJ96] Tracey W. Steig and Samuel V. Johnston. Monitoring fish movement patterns in a reservoir using horizontally scanning split-beam techniques. ICES Journal of Marine Science, 53 (2):435–441, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/53/2/435/792823.

Smith:1990:ESS

[SJB90] P. J. Smith, A. Jamieson, and A. J. Birley. Electrophoretic studies and the stock concept in marine teleosts. ICES Journal of Marine Science, 47(2):231–245, ???? 1990. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
47/2/231/611359.

Svaasand:1990:ESCa

[SJK90] T. Svåsand, K. E. Jørstad, and T. S. Kristiansen. Enhancement studies of coastal cod in western Norway. Part I. Recruitment of wild and reared cod to a local spawning stock. ICES Journal of Marine Science, 47(1):5–12, ???? 1990. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/47/1/5/638523.

Svaasand:1990:ESCb

[SK90a] T. Svåsand and T. S. Kristiansen. Enhancement studies of coastal cod in western Norway. part II. Migration of reared coastal cod. ICES Journal of Marine Science, 47(1):13–22, ???? 1990. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/47/1/13/638503.

Svaasand:1990:ESCc

[SK90b] T. Svåsand and T. S. Kristiansen. Enhancement studies of coastal cod in western Norway. Part IV. Mortality of reared cod after release. ICES Journal of Marine Science, 47(1):30–39, ???? 1990. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/47/1/30/638509.

Squires:1995:RRS

[SK95a] Dale Squires and James Kirkley. Resource rents from single and multispecies individual transferable quota programs. ICES Journal of Marine Science, 52(2):153–164, April 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
52/2/153/593889.

Strachan:1995:PMD

[SK95b] N. J. C. Strachan and L. Kell. A potential method for the differentiation between haddock fish stocks by computer vision using canonical discriminant analysis. ICES Journal of Marine Science, 52(1):145–149, February 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/52/1/145/699758.

Stange:1997:OCF

[SK97] Kari Stange and Jarle Klungsøyr. Organochlorine contaminants in fish and polycyclic aromatic hydrocarbons in sediments from the Barents Sea. ICES Journal of Marine Science, 54(3):318–332, June 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/54/3/318/652459.

Serchuk:1996:STM

[SKD96] Fredric M. Serchuk, Eskild Kirkegaard, and Niels Daan. Status and trends of the major roundfish, flatfish, and pelagic fish stocks in the North Sea: thirty-year overview. ICES Journal of Marine Science, 53(6):1130–1145, December 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
53/6/1130/592486.

Schram:1998:SOL

[SKHM98] T. A. Schram, J. A. Knutsen, P. A. Heuch, and T. A. Mo. Seasonal occurrence of Lepeophtheirus salmonis and Caligus elongatus (Copepoda: Caligidae) on sea trout (Salmo trutta), off southern Norway. ICES Journal of Marine Science, 55 (2):163–175, April 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/55/2/163/781694.

Scholten:1998:TVC

[SKL98] M. C. Th. Scholten, K. J. M. Kramer, and R. W. P. M. Laane. Trends and variation in concentration of dissolved metals (Cd, Cu, Pb, and Zn) in the North Sea (1980–1989). ICES Journal of Marine Science, 55(5):825–834, October 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
55/5/825/690927.

Sakuramoto:1997:RBT

[SKS97] Kazumi Sakuramoto, Takeru Kitahara, and Hideki Sugiyama. Relationship between temperature and fluctuations in sandfish catch (Arctoscopus japonicus) in the coastal waters off akita prefecture. ICES Journal of Marine Science, 54(1): 1–12, February 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/54/1/1/782126.

Stokesbury:1997:HIH

[SL97] M. J. Stokesbury and G. L. Lacroix. High incidence of hatchery origin Atlantic salmon in the smolt output of a Canadian river. ICES Journal of Marine Science, 54(6):1074–1081, December 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/54/6/1074/673929.

Smayda:1998:PVC

[Sma98] Theodore J. Smayda. Patterns of variability characterizing marine phytoplankton, with examples from Narragansett Bay. ICES Journal of Marine Science, 55(4):562–573, August 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
55/4/562/698132.

Solari:1997:SRB

[SMGB97] A. P. Solari, J. M. Martín-González, and C. Bas. Stock and recruitment in Baltic cod (Gadus morhua): a new, non-linear approach. ICES Journal of Marine Science, 54(3):427–443, June 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/54/3/427/652482.

Smith:1995:ASM

[Smi95] Sharon L. Smith. The Arabian Sea: Mesozooplankton response to seasonal climate in a tropical ocean: II. Regional interactions of physics and population dynamics. ICES Journal of Marine Science, 52(3–4):427–438, June 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https:/

/academic.oup.com/icesjms/article/52/3-4/427/841574.

Shepherd:1991:MMC

[SN91] J. G. Shepherd and M. D. Nicholson. Multiplicative modelling of catch-at-age data, and its application to catch forecasts. ICES Journal of Marine Science, 47(3):284–294, ???? 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
47/3/284/601167.

Seiderer:1999:ARB

[SN99] L. J. Seiderer and R. C. Newell. Analysis of the relationship between sediment composition and benthic community structure in coastal deposits: Implications for marine aggregate dredging. ICES Journal of Marine Science, 56(5):757–765, October 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
56/5/757/691342.

Swain:1994:CCA

[SNSC94] D. P. Swain, G. A. Nielsen, A. F. Sinclair, and G. A. Chouinard. Changes in catchability of Atlantic cod (Gadus morhua) to an otter-trawl fishery and research survey in the southern Gulf of St Lawrence. ICES Journal of Marine Science, 51(4):493–504, August 1994. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/51/4/493/714263.

Sinclair:1999:GMT

[SOBP99] M. Sinclair, R. N. O’Boyle, D. L. Burke, and F. G. Peacock. Groundfish management in transition within the Scotia–Fundy area of Canada. ICES Journal of Marine Science, 56(6):1014– 1023, December 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/56/6/1014/658099.

Somerton:1999:IED

[Som99] D. Somerton. Incorporating experimentally derived estimates of survey trawl efficiency into the stock assessment process: a discussion. ICES Journal of Marine Science, 56(3):299–302, June 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/56/3/299/623506.

St-Pierre:1992:VDS

[SP92] Gilbert St-Pierre. Visual determination of sex in live Pacific halibut. ICES Journal of Marine Science, 49(3):373–376, August 1992. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/49/3/373/819865.

Smith:1996:ABA

[SP96] Stephen J. Smith and Fred H. Page. Associations between Atlantic cod (Gadus morhua) and hydrographic variables: implications for the management of the 4VsW cod stock. ICES Journal of Marine Science, 53(3):597–614, June 1996. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
53/3/597/625712.

Sparholt:1990:ETB

[Spa90a] Henrik Sparholt. An estimate of the total biomass of fish in the North Sea. ICES Journal of Marine Science, 46(2):200–210, ???? 1990. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/46/2/200/645388.

Sparholt:1990:IEN

[Spa90b] Henrik Sparholt. Improved estimates of the natural mortality rates of nine commercially important fish species included in the North Sea multispecies VPA model. ICES Journal of Marine Science, 46(2):211–223, ???? 1990. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/46/2/211/645398.

Sparholt:1990:SIV

[Spa90c] Henrik Sparholt. A stochastic integrated VPA for herring in the Baltic Sea using acoustic estimates as auxiliary information for estimating natural mortality. ICES Journal of Marine Science, 46(3):325–332, ???? 1990. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/46/3/325/596418.

Sparholt:1995:UMM

[Spa95] Henrik Sparholt. Using the MSVPA/MSFOR model to estimate the right-hand side of the Ricker curve for Baltic cod. ICES Journal of Marine Science, 52(5):819–826, October 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
52/5/819/783866.

Sparholt:1996:CCB

[Spa96] H. Sparholt. Causal correlation between recruitment and spawning stock size of central Baltic cod? ICES Journal of Marine Science, 53(5):771–779, October 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/53/5/771/704298.

Suuronen:1996:SRM

[SPCLT96] Petri Suuronen, Jose A. Perez-Comas, Esa Lehtonen, and Vesa Tschernij. Size-related mortality of herring (Clupea harengus L.) escaping through a rigid sorting grid and trawl codend meshes. ICES Journal of Marine Science, 53(4):691–700, August 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/53/4/691/799724.

Stergiou:1997:SEM

[SPCP97] K. I. Stergiou, C.-Y. Politou, E. D. Christou, and G. Petrakis. Selectivity experiments in the NE Mediterranean: the effect of trawl codend mesh size on species diversity and discards. ICES Journal of Marine Science, 54(5):774–786, October 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
54/5/774/751473.

Sanchez:1998:DAM

[SPL98] F. Sánchez, N. Pérez, and J. Landa. Distribution and abundance of megrim (Lepidorhombus boscii and Lepidorhombus whiffiagonis) on the northern Spanish shelf. ICES Journal of Marine Science, 55(3):494–514, June 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/55/3/494/655080.

Stephenson:1999:SAM

[SRAL99] R. L. Stephenson, K. Rodman, D. G. Aldous, and D. E. Lane. An in-season approach to management under uncertainty: the case of the SW Nova Scotia herring fishery. ICES Journal of Marine Science, 56(6):1005–1013, December 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
56/6/1005/658096.

Saborido-Rey:1998:HAV

[SRJ98] Fran Saborido-Rey and Susana Junquera. Histological assessment of variations in sexual maturity of cod (Gadus morhua L.) at the Flemish Cap (north-west Atlantic). ICES Journal of Marine Science, 55(3):515–521, June 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/55/3/515/655084.

Schobben:1993:PMM

[SS93a] H. P. M. Schobben and M. C. Th. Scholten. Probabilistic methods for marine ecological risk assessment. ICES Journal of Marine Science, 50(4):349–358, ???? 1993. CODEN ICESEC.

ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/50/4/349/668222.

Suthers:1993:DGP

[SS93b] Iain M. Suthers and Svein Sundby. Dispersal and growth of pelagic juvenile arcto — Norwegian cod (Gadus morhua), inferred from otolith microstructure and water temperature. ICES Journal of Marine Science, 50(3):261–270, ???? 1993. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
50/3/261/685303.

Sprules:1995:SBB

[SS95a] W. G. Sprules and J. D. Stockwell. Size-based biomass and production models in the st Lawrence great lakes. ICES Journal of Marine Science, 52(3–4):705–710, June 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/52/3-4/705/841896.

Stockwell:1995:STP

[SS95b] J. D. Stockwell and W. G. Sprules. Spatial and temporal patterns of zooplankton biomass in lake erie. ICES Journal of Marine Science, 52(3–4):557–564, June 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https:/

/academic.oup.com/icesjms/article/52/3-4/557/841699.

Suthers:1996:RMS

[SS96] Iain M. Suthers and Svein Sundby. Role of the midnight sun: comparative growth of pelagic juvenile cod (Gadus morhua) from the Arcto–Norwegian and a Nova Scotian stock. ICES Journal of Marine Science, 53(5):827–836, October 1996. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
53/5/827/704325.

Stensholt:1999:FMV

[SS99] E. Stensholt and B. K. Stensholt. Fish movement vectors and the temperature gradient: a geometric analysis method for the depth–temperature time series from data storage tags. ICES Journal of Marine Science, 56(4):537–544, August 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
56/4/537/624666.

Selivanovsky:1996:AIP

[SSD96] D. A. Selivanovsky, P. A. Stunzhas, and I. N. Didenkulov. Acoustical investigation of phytoplankton. ICES Journal of Marine Science, 53(2):313–316, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/53/2/313/792576.

Shackell:1997:GCG

[SSFB97] N. L. Shackell, W. T. Stobo, K. T. Frank, and D. Brickman. Growth of cod (Gadus morhua) estimated from markrecapture programs on the Scotian Shelf and adjacent areas. ICES Journal of Marine Science, 54(3):383–398, June 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
54/3/383/652476.

Schnack-Schiel:1995:LCS

[SSH95] Sigrid B. Schnack-Schiel and Wilhelm Hagen. Life-cycle strategies of Calanoides acutus, Calanus propinquus, and Metridia gerlachei (Copepoda: Calanoida) in the eastern Weddell Sea, Antarctica. ICES Journal of Marine Science, 52(3–4):541– 548, June 1995. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/52/3-4/541/841670.

Sherman:1998:BAZ

[SSJK98] K. Sherman, A. Solow, J. Jossi, and J. Kane. Biodiversity and abundance of the zooplankton of the northeast shelf ecosystem. ICES Journal of Marine Science, 55(4):730–738, August 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
55/4/730/698179.

Swartzman:1994:RDP

[SSKP94] G. Swartzman, W. Stuetzle, K. Kulman, and M. Powojowski. Relating the distribution of pollock schools in the Bering Sea to environmental factors. ICES Journal of Marine Science, 51 (4):481–492, August 1994. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/51/4/481/714261.

Stefansson:1998:AEB

[SSS98] G. Stefánsson, U. Skúladóttir, and B. Æ. Steinarsson. Aspects of the ecology of a boreal system. ICES Journal of Marine Science, 55(5):859–862, October 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/55/5/859/690936.

Smith:1999:IEF

[SSS99] A. D. M. Smith, K. J. Sainsbury, and R. A. Stevens. Implementing effective fisheries-management systems — management strategy evaluation and the Australian partnership approach. ICES Journal of Marine Science, 56(6):967–979, December 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/56/6/967/658156.

Stanton:1994:AEZ

[Sta94] T. Stanton. On acoustic estimates of zooplankton biomass. ICES Journal of Marine Science, 51(4):505–512, August 1994. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
51/4/505/714266.

Stefansson:1996:AGS

[Ste96] Gunnar Stefánsson. Analysis of groundfish survey abundance data: combining the GLM and delta approaches. ICES Journal of Marine Science, 53(3):577–588, June 1996. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
53/3/577/625707.

Stergiou:1999:ECS

[Ste99] K. I. Stergiou. Effects of changes in the size and shape of codend on catch of Aegean Sea fishes. ICES Journal of Marine Science, 56(1):96–102, February 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/56/1/96/652215.

Sutinen:1999:WWW

[Sut99] J. G. Sutinen. What works well and why: evidence from fishery-management experiences in OECD countries. ICES Journal of Marine Science, 56(6):1051–1058, December 1999.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
56/6/1051/658107.

Sparholt:1991:BSR

[SV91] Henrik Sparholt and Morten Vinther. The biomass of starry ray (Raja radiata) in the North Sea. ICES Journal of Marine Science, 47(3):295–302, ???? 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/47/3/295/601168.

Sanchez-Velasco:1998:DCF

[SV98] L. Sánchez-Velasco. Diet composition and feeding habits of fish larvae of two co-occurring species (Pisces: Callionymidae and Bothidae) in the North-western Mediterranean. ICES Journal of Marine Science, 55(2):299–308, April 1998. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
55/2/299/781762.

Suthers:1999:GHD

[SVJ99] I. M. Suthers, T. Van Der Meeren, and K. E. Jørstad. Growth histories derived from otolith microstructure of three Norwegian cod stocks co-reared in mesocosms; effect of initial size and prey size changes. ICES Journal of Marine Science, 56 (5):658–672, October 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/56/5/658/691321.

Sanchez-Velasco:1999:SDS

[SVS99] L. Sánchez-Velasco and B. Shirasago. Spatial distribution of some groups of microzooplankton in relation to oceanographic processes in the vicinity of a submarine canyon in the northwestern Mediterranean Sea. ICES Journal of Marine Science, 56(1):1–14, February 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/56/1/1/652183.

Stanley:1996:AFA

[SW96] David R. Stanley and Charles A. Wilson. Abundance of fishes associated with a petroleum platform as measured with dualbeam hydroacoustics. ICES Journal of Marine Science, 53 (2):473–475, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/53/2/473/792898.

Swartzman:1997:ASD

[Swa97] G. Swartzman. Analysis of the summer distribution of fish schools in the Pacific Eastern Boundary Current. ICES Journal of Marine Science, 54(1):105–116, February 1997. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
54/1/105/782132.

Swartzman:1999:SPA

[Swa99] G. Swartzman. Spatial proximity of age-0 walleye pollock (Theragra chalcogramma) to zooplankton near the Pribilof Islands, Bering Sea, Alaska. ICES Journal of Marine Science, 56 (4):545–560, August 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/56/4/545/624674.

Socha:1996:VBP

[SWB96] David G. Socha, Jonathan L. Watkins, and Andrew S. Brierley. A visualization-based post-processing system for analysis of acoustic data. ICES Journal of Marine Science, 53(2):335–338, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/53/2/335/792617.

Stanton:1994:ACD

[SWCG94] Timothy K. Stanton, Peter H. Wiebe, Dezhang Chu, and Luois Goodman. Acoustic characterization and discrimination of marine zooplankton and turbulence. ICES Journal of Marine Science, 51(4):469–479, August 1994. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/51/4/469/714259.

Temming:1994:MGE

[TA94] Axel Temming and Niels G. Andersen. Modelling gastric evacuation without meal size as a variable. a model applicable for the estimation of daily ration of cod (Gadus morhua L.) in the field. ICES Journal of Marine Science, 51(4):429–438, August 1994. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
51/4/429/714253.

Taylor:1995:NSS

[Tay95] Arnold H. Taylor. North–south shifts of the Gulf stream and their climatic connection with the abundance of zooplankton in the UK and its surrounding seas. ICES Journal of Marine Science, 52(3–4):711–721, June 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https:/

/academic.oup.com/icesjms/article/52/3-4/711/841904.

Tregenza:1997:HPP

[TBHL97] N. J. C. Tregenza, S. D. Berrow, P. S. Hammond, and R. Leaper. Harbour porpoise (Phocoena phocoena L.) bycatch in set gillnets in the Celtic Sea. ICES Journal of Marine Science, 54(5):896–904, October 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/54/5/896/751546.

Tuck:1997:PBN

[TCA97] I. D. Tuck, C. J. Chapman, and R. J. A. Atkinson. Population biology of the Norway lobster, Nephrops norvegicus (L.) in the Firth of Clyde, Scotland — I: Growth and density. ICES Journal of Marine Science, 54(1):125–135, February 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/54/1/125/782143.

Takao:1996:DBE

[TF96] Yoshimi Takao and Masahiko Furusawa. Dual-beam echo integration method for precise acoustic surveys. ICES Journal of Marine Science, 53(2):351–358, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/53/2/351/792656.

Tiselius:1995:CWU

[THJ^{+}95] Peter Tiselius, Benni Hansen, Per Jonsson, Thomas Kiørboe, Torkel Gissel Nielsen, Sergey Piontkovski, and Enric Saiz. Can we use laboratory-reared copepods for experiments? A comparison of feeding behaviour and reproduction between a field and a laboratory population of Acartia tonsa. ICES Journal of Marine Science, 52(3–4):369–376, June 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https:/

/academic.oup.com/icesjms/article/52/3-4/369/841530.

Thurow:1997:ETF

[Thu97] Fritz Thurow. Estimation of the total fish biomass in the Baltic Sea during the 20th century. ICES Journal of Marine Science, 54(3):444–461, June 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/54/3/444/652485.

Tingley:1997:OLS

[TIR97] G. A. Tingley, M. J. Ives, and I. C. Russell. The occurrence of lice on sea trout (Salmo trutta L.) captured in the sea off the East Anglian coast of England. ICES Journal of Marine Science, 54(6):1120–1128, December 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/54/6/1120/673968.

Thorarinsson:1997:CVC

[TJ97] Kristján Thórarinsson and Gardar Jóhannesson. Correcting for variation in catchability: maturity-related catchability variation in Icelandic cod. ICES Journal of Marine Science, 54(5): 787–796, October 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/54/5/787/751478.

Thompson:1996:DFS

[TL96] Charles H. Thompson and Richard H. Love. Determination of fish size distributions and areal densities using broadband lowfrequency measurements. ICES Journal of Marine Science, 53 (2):197–201, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/53/2/197/792385.

Toresen:1990:LTC

[Tor90] R. Toresen. Long-term changes in growth of Norwegian springspawning herring. ICES Journal of Marine Science, 47(1): 48–56, ???? 1990. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/47/1/48/638520.

Toresen:1991:PEN

[Tor91] Reidar Toresen. Predation on the eggs of Norwegian springspawning herring (Clupea harengus L.) on a spawning ground on the west coast of Norway. ICES Journal of Marine Science, 48(1):15–21, May 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/48/1/15/696105.

Tasker:1997:SME

[TR97] Mark L. Tasker and James B. Reid. Seabirds in the marine environment. ICES Journal of Marine Science, 54(4):505–506, August 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/54/4/505/607421.

Traynor:1996:TSM

[Tra96] Jimmie J. Traynor. Target-strength measurements of walleye pollock (Theragra chalcogramma) and Pacific whiting (Merluccius productus). ICES Journal of Marine Science, 53(2): 253–258, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/53/2/253/792486.

Tameishi:1996:UJS

[TSAS96] H. Tameishi, H. Shinomiya, I. Aoki, and T. Sugimoto. Understanding Japanese sardine migrations using acoustic and other aids. ICES Journal of Marine Science, 53(2):167–171, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/53/2/167/791622.

Turrell:1996:HES

[TSP^{+}96] W. R. Turrell, G. Slesser, R. Payne, R. D. Adams, and P. A. Gillibrand. Hydrography of the East Shetland Basin in relation to decadal North Sea variability. ICES Journal of Marine Science, 53(6):899–916, December 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/53/6/899/592521.

Tarbox:1996:AAS

[TT96] K. E. Tarbox and R. E. Thorne. Assessment of adult salmon in near-surface waters of Cook Inlet, Alaska. ICES Journal of Marine Science, 53(2):397–401, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/53/2/397/792738.

Turrell:1992:NHC

[Tur92] W. R. Turrell. New hypotheses concerning the circulation of the northern North Sea and its relation to North Sea fish stock recruitment. ICES Journal of Marine Science, 49(1): 107–123, February 1992. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/49/1/107/732755.

Tackx:1995:MSF

[TZD^{+}95] M. L. M. Tackx, L. Zhu, W. De Coster, R. Billones, and M. H. Daro. Measuring selectivity of feeding by estuarine copepods using image analysis combined with microscopic and Coulter counting. ICES Journal of Marine Science, 52(3– 4):419–425, June 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.oup.

com/icesjms/article/52/3-4/419/841566.

Ueberschar:1992:CNC

[UC92] B. Ueberschär and C. Clemmesen. A comparison of the nutritional condition of herring larvae as determined by two biochemical methods — tryptic enzyme activity and RNA/DNA ratio measurements. ICES Journal of Marine Science, 49 (2):245–249, May 1992. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/49/2/245/683933.

Ugland:1993:FCI

[UJA^{+}93] K. I. Ugland, K. A. Jødestøl, P. E. Aspholm, A. B. Krøyer, and T. Jakobsen. Fish consumption by invading harp seals off the Norwegian coast in 1987 and 1988. ICES Journal of Marine Science, 50(1):27–38, ???? 1993. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/50/1/27/652878.

Ulltang:1996:SAB

[Ull96] Øyvind Ulltang. Stock assessment and biological knowledge: can prediction uncertainty be reduced? ICES Journal of Marine Science, 53(4):659–675, August 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/53/4/659/799712.

vanderMeer:1995:APS

[vdML95] Jaap van der Meer and Mardik F. Leopold. Assessing the population size of the European storm petrel (Hydrobates pelagicus) using spatial autocorrelation between counts from segments of criss-cross ship transects. ICES Journal of Marine Science, 52 (5):809–818, October 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/52/5/809/783860.

vanderWateren:1990:CES

[vdWRvdLF90] P. van der Wateren, A. D. Rijnsdorp, M. A. van der Land, and M. Fonds. Changes in egg size of plaice (Pleuronectes platessa L.) during incubation, and the effect of fixation. ICES Journal of Marine Science, 47(2):264–266, ???? 1990. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
47/2/264/611389.

Verspoor:1997:GDA

[Ver97] E. Verspoor. Genetic diversity among Atlantic salmon (Salmo salar L.) populations. ICES Journal of Marine Science, 54(6): 965–973, December 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/54/6/965/674075.

Viette:1997:RBS

[VGF97] Michèle Viette, Piero Giulio Giulianini, and Enrico Antonio Ferrero. Reproductive biology of scad, Trachurus mediterraneus (Teleostei, Carangidae), from the Gulf of Trieste. ICES Journal of Marine Science, 54(2):267–272, April 1997. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
54/2/267/620555.

VanDamme:1997:SDL

[VGHO97] P. A. Van Damme, A. Geets, O. Hamerlynck, and F. Ollevier. The suprapopulation dynamics of Lernaeocera branchialis and L. lusci in the oosterschelde: seasonal abundance on three definitive host species. ICES Journal of Marine Science, 54 (1):24–31, February 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/54/1/24/782177.

Vuorinen:1998:PCB

[VHV^{+}98] I. Vuorinen, J. Hänninen, M. Viitasalo, U. Helminen, and H. Kuosa. Proportion of copepod biomass declines with decreasing salinity in the Baltic Sea. ICES Journal of Marine Science, 55(4):767–774, August 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/55/4/767/698220.

Virtala:1998:SVP

[VKA98] M. Virtala, S. Kuikka, and E. Arjas. Stochastic virtual population analysis. ICES Journal of Marine Science, 55(5): 892–904, October 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/55/5/892/690959.

vanLeussen:1996:NWE

[vLRvR^{+}96] Wim van Leussen, Günther Radach, Wim van Raaphorst, Franciscus Colijn, and Remi Laane. The North-West European Shelf Programme (NOWESP): integrated analysis of shelf processes based on existing data sets and models. ICES Journal of Marine Science, 53(6):926–932, December 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
53/6/926/592526.

Valdes:1998:TSA

[VM98] Luis Valdés and Mercedes Moral. Time-series analysis of copepod diversity and species richness in the southern Bay of Biscay off Santander, Spain, in relation to environmental conditions. ICES Journal of Marine Science, 55(4):783–792, August 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
55/4/783/698232.

Vollestad:1990:OME

[Vøl90] Leif Asbjørn Vøllestad. Optimal management of European eel in the Imsa River, Norway. ICES Journal of Marine Science, 46(2):115–120, ???? 1990. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/46/2/115/645327.

Verdoit:1999:GMI

[VPT99] M. Verdoit, D. Pelletier, and C. Talidec. A growth model that incorporates individual variability for the Norway lobster population (Nephrops norvegicus, L. 1758) from the Bay of Biscay. ICES Journal of Marine Science, 56(5):734–745, October 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
56/5/734/691337.

Verheye:1998:LTI

[VR98] Hans M. Verheye and Anthony J. Richardson. Long-term increase in crustacean zooplankton abundance in the southern Benguela upwelling region (1951–1996): bottom-up or topdown control? ICES Journal of Marine Science, 55(4):803–807, August 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/55/4/803/698239.

Vranken:1991:EPL

[VVH91] G. Vranken, R. Vanderhaeghen, and C. Heip. Effects of pollutants on life-history parameters of the marine nematode Monhystera disjuncta. ICES Journal of Marine Science, 48(3): 325–334, November 1991. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/48/3/325/737264.

Volovik:1995:MMS

[VVM95] Yegor S. Volovik, Stanislav P. Volovik, and Zinaida A. Myrzoyan. Modelling of the Mnemiopsis sp. population in the Azov Sea. ICES Journal of Marine Science, 52(3–4):735– 746, June 1995. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/52/3-4/735/841922.

Weston:1990:MEF

[WA90] D. E. Weston and H. W. Andrews. Monthly estimates of fish numbers using a long-range sonar. ICES Journal of Marine Science, 47(1):104–111, ???? 1990. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/47/1/104/638497.

Waldron:1994:VAS

[Wal94] Miranda E. Waldron. Validation of annuli of the South African anchovy, Engraulis capensis, using daily otolith growth increments. ICES Journal of Marine Science, 51(2):233–234, ???? 1994. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/51/2/233/674485.

Walsh:1999:I

[Wal99] S. J. Walsh. Introduction. ICES Journal of Marine Science, 56(3):251, June 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/56/3/251/623485.

Wanless:1997:MEE

[Wan97] S. Wanless. Modelling environmental and energetic effects on feeding performance and distribution of shags (Phalacrocorax aristotelis): integrating telemetry, geographical information systems, and modelling techniques. ICES Journal of Marine Science, 54(4):524–544, August 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/54/4/524/607437.

Warren:1994:PSP

[War94] William G. Warren. The potential of sampling with partial replacement for fisheries surveys. ICES Journal of Marine Science, 51(3):315–324, ???? 1994. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/51/3/315/620223.

Woods:1995:MOZ

[WB95] J. D. Woods and W. Barkmann. Modelling oligotrophic zooplankton production: seasonal oligotrophy off the Azores. ICES Journal of Marine Science, 52(3–4):723–734, June 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/52/3-4/723/841913.

Watkins:1996:PPT

[WB96] Jonathan L. Watkins and Andrew S. Brierley. A postprocessing technique to remove background noise from echo integration data. ICES Journal of Marine Science, 53(2):

339–344, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/53/2/339/792631.

Walter:1997:OCS

[WB97a] U. Walter and P. H. Becker. Occurrence and consumption of seabirds scavenging on shrimp trawler discards in the Wadden Sea. ICES Journal of Marine Science, 54(4):684–694, August 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/54/4/684/607487.

Wright:1997:SCC

[WB97b] P. J. Wright and G. S. Begg. A spatial comparison of common guillemots and sandeels in Scottish waters. ICES Journal of Marine Science, 54(4):578–592, August 1997. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
54/4/578/607450.

Walker:1996:LTC

[WH96] P. A. Walker and H. J. L. Heessen. Long-term changes in ray populations in the North Sea. ICES Journal of Marine Science, 53(6):1085–1093, December 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/53/6/1085/592475.

Wlodarczyk:1997:SSV

[WH97] Ewa Wlodarczyk and Katarzyna Horbowa. Size-specific vertical distribution of Baltic cod (Gadus morhua L.) eggs in the Bornholm Basin in 1993 and 1994. ICES Journal of Marine Science, 54(2):206–212, April 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/54/2/206/620542.

Walker:1998:SSR

[WH98] P. A. Walker and J. R. G. Hislop. Sensitive skates or resilient rays? Spatial and temporal shifts in ray species composition in the central and north-western North Sea between 1930 and the present day. ICES Journal of Marine Science, 55(3):392–402, June 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/55/3/392/655038.

Wanless:1998:SSC

[WHG98] S. Wanless, M. P. Harris, and S. P. R. Greenstreet. Summer sandeel consumption by seabirds breeding in the Firth of Forth, south-east Scotland. ICES Journal of Marine Science, 55(6): 1141–1151, December 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/55/6/1141/638996.

Walker:1997:DMS

[WHM97] Paddy Walker, Gary Howlett, and Richard Millner. Distribution, movement and stock structure of three ray species in the North Sea and eastern English Channel. ICES Journal of Marine Science, 54(5):797–808, October 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/54/5/797/751483.

Winger:1999:SEA

[WHW99] P. D. Winger, P. He, and S. J. Walsh. Swimming endurance of American plaice (Hippoglossoides platessoides) and its role in fish capture. ICES Journal of Marine Science, 56(3):252–265, June 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/56/3/252/623489.

Witbaard:1996:GVA

[Wit96] Rob Witbaard. Growth variations in Arctica islandica L. (Mollusca): a reflection of hydrography-related food supply. ICES Journal of Marine Science, 53(6):981–987, December 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
53/6/981/592543.

Witting:1999:OMP

[Wit99] L. Witting. Optimization of management procedures with control on uncertainty risk. ICES Journal of Marine Science, 56 (6):876–883, December 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/56/6/876/658136.

Witbaard:1994:LTT

[WK94] R. Witbaard and R. Klein. Long-term trends on the effects of the southern North Sea beamtrawl fishery on the bivalve mollusc Arctica islandica L. (Mollusca, bivalvia). ICES Journal of Marine Science, 51(1):99–105, ???? 1994. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/51/1/99/760314.

Wainwright:1999:MER

[WK99] T. C. Wainwright and R. G. Kope. Methods of extinction risk assessment developed for US West Coast salmon. ICES Journal of Marine Science, 56(4):444–448, August 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
56/4/444/624630.

Weinberg:1999:EAL

[WM99] K. L. Weinberg and P. T. Munro. The effect of artificial light on escapement beneath a survey trawl. ICES Journal of Marine Science, 56(3):266–274, June 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/56/3/266/623494.

White:1998:NDA

[WMO98] M. White, C. Mohn, and M. J. Orren. Nutrient distributions across the Porcupine Bank. ICES Journal of Marine Science, 55(6):1082–1094, December 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/55/6/1082/638980.

Walin:1993:PBK

[WO93] Gøsta Walin and Ingemar Olsson. Professor Börje Kullenberg 1906–1991. ICES Journal of Marine Science, 50(1):101–102, ???? 1993. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/50/1/101/652872.

Walsh:1995:UMM

[WRT95] Martin Walsh, David G. Reid, and William R. Turrell. Understanding mackerel migration off Scotland: Tracking with echosounders and commercial data, and including environmental correlates and behaviour. ICES Journal of Marine Science, 52(6):925–939, December 1995. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/52/6/925/649267.

Williamson:1996:AOD

[WT96] Neal J. Williamson and Jimmie J. Traynor. Application of a one-dimensional geostatistical procedure to fisheries acoustic surveys of Alaskan pollock. ICES Journal of Marine Science, 53(2):423–428, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/53/2/423/792797.

Wiklund:1999:RAA

[WTB99] T. Wiklund, I. Tabolina, and T. V. Bezgachina. Recovery of atypical Aeromonas salmonicida from ulcerated fish from the Baltic Sea. ICES Journal of Marine Science, 56(2):175–179, April 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/56/2/175/622463.

Winters:1996:EPF

[WW96] G. H. Winters and J. P. Wheeler. Environmental and phenotypic factors affecting the reproductive cycle of Atlantic herring. ICES Journal of Marine Science, 53(1):73–88, February 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/53/1/73/731588.

Winters:1993:VRO

[WWS93] G. H. Winters, J. P. Wheeler, and D. Stansbury. Variability in the reproductive output of spring-spawning herring in the north-west Atlantic. ICES Journal of Marine Science, 50(1):15–25, ???? 1993. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/50/1/15/652875.

Ye:1996:FDA

[YCL96] Zhen Ye, Terry Curran, and David Lemon. Fish detection by the acoustic scintillation technique. ICES Journal of Marine Science, 53(2):317–321, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/53/2/317/792589.

Ye:1998:AEC

[Ye98] Y. Ye. Assessing effects of closed seasons in tropical and subtropical penaeid shrimp fisheries using a length-based yieldper-recruit model. ICES Journal of Marine Science, 55(6):

1112–1124, December 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/55/6/1112/638986.

Ye:1996:ASF

[YF96] Zhen Ye and David M. Farmer. Acoustic scattering by fish in the forward direction. ICES Journal of Marine Science, 53 (2):249–252, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/53/2/249/792474.

Yndestad:1999:ENIb

[Ynd99a] H. Yndestad. Earth nutation influence on system dynamics of Northeast Arctic cod. ICES Journal of Marine Science, 56(5): 652–657, October 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/56/5/652/691320.

Yndestad:1999:ENIa

[Ynd99b] H. Yndestad. Earth nutation influence on the temperature regime of the Barents Sea. ICES Journal of Marine Science, 56(3):381–387, June 1999. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/56/3/381/623528.

Youngson:1997:FOR

[YWMW97] A. F. Youngson, J. H. Webb, J. C. MacLean, and B. M. Whyte. Frequency of occurrence of reared Atlantic salmon in Scottish salmon fisheries. ICES Journal of Marine Science, 54(6):1216– 1220, December 1997. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/54/6/1216/674027.

Zeldis:1998:DEP

[ZF98] J. R. Zeldis and R. I. C. C. Francis. A daily egg production method estimate of snapper biomass in Hauraki Gulf, New Zealand. ICES Journal of Marine Science, 55(3):522–534, June 1998. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/55/3/522/655086.

Zhang:1993:CWT

[ZM93] Z. Zhang and E. Moksness. A chemical way of thinning otoliths of adult Atlantic herring (Clupea harengus) to expose the microstructure in the nucleus region. ICES Journal of Marine Science, 50(2):213–217, ???? 1993. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/50/2/213/774830.

Zakharia:1996:WSF

[ZMHD96] Manell E. Zakharia, François Magand, François Hetroit, and Noël Diner. Wideband sounder for fish species identification at sea. ICES Journal of Marine Science, 53(2):203–208, April 1996. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/53/2/203/792397.